### v1

# Two-Timescale Optimization Framework for Decentralized Linear-Quadratic Optimal Control

Abstract
This study investigates a decentralized linear-quadratic optimal control problem, and several approximate separable constrained optimization problems are formulated for the first time based on the selection of sparsity promoting functions. First, for the optimization problem with weighted $\ell_{1}$ sparsity promoting function, a two-timescale algorithm is adopted that is based on the BSUM (Block Successive Upper-bound Minimization) framework and a differential equation solver. Second, a piecewise quadratic sparsity promoting function is introduced, and the induced optimization problem demonstrates an accelerated convergence rate by performing the same two-timescale algorithm. Finally, the optimization problem with $\ell_{0}$ sparsity promoting function is considered that is nonconvex and discontinuous, and can be approximated by successive coordinatewise convex optimization problems.
Keywords: decentralized linear-quadratic optimal control, separable optimization problem, two-timescale algorithm, $\ell_{0}$ sparsity promoting function.
1 Introduction
During the last few decades, the concept of decentralized control has gained significant attention due to its appealing characteristics, such as the potential to lower communication costs and parallelizable processing. In particular, decentralized linear-quadratic (LQ) optimal control is also a hot topic, which has been extensively investigated [25, 26, 7, 29]. Nevertheless, decentralized LQ problems present a significant challenge, as obtaining a closed solution is exceedingly difficult, with the exception of highly specific cases [46]. Optimization, an “end-to-end” method allowing us to optimize the performance criterion directly, has significantly contributed to the centralized LQ problems [11, 35, 16, 9, 8, 38, 10, 53] and decentralized LQ problems [52, 13, 28, 49], respectively. However, the question of how to design a controller with a general sparsity structure has yet to be resolved. Here, the term “sparsity structure” in this context pertains to the presence of zero elements within the stabilizing linear feedback gain matrix. Motivated by recent achievements in optimization-based LQ controls, the objective of this study is to establish a convex optimization framework for decentralized control incorporating a sparsity promoting penalty, which allows us to obtain an feedback gain with arbitrary sparsity structure.
1.1 Related Works
1) Optimization-based LQ control. LQ optimal control is a landmark problem in the field of optimal control theory, and direct policy gradient is a hot topic of centralized LQ problems in recent years. If the full state is measured, the LQ performance criterion satisfies the Polyak-Lojasiewicz (PL) condition [10]. Hence, gradient-based methods can be ensured the convergence to the global optimal linear feedback gain exponentially in both model-based and model-free settings [11, 35, 38]. The challenge of LQ problems is significantly amplified when the state cannot be fully measured, mostly due to the absence of the PL condition and the resulting disconnection of feasible set [16, 9, 8, 53, 14].
On the other hands, a canonical decentralized LQ optimal control can be cast as a centralized problem where the stabilizing feedback gain matrix is restricted to lie in a particular subspace $\mathcal{K}$. If $\mathcal{K}$ denotes the block-diagonal matrix space, classic convex parametrization methods can be utilized to obtain an approximate convex problem [9], and ADMM (Alternating Direction Method of Multipliers) based method is shown to converge to the global optimizer of the abovementioned approximate convex problem [52, 49, 24]. If $\mathcal{K}$ possesses general sparsity structure, the SDP (Semidefinite Programming) relaxation perspective is of concern. If the SDP relaxation has a rank-$1$ solution, a globally linear optimal distributed controller can be recovered from the solution of SDP [13]. Additionally, zero-order policy optimization algorithm is utilized to optimize the decentralized LQ cost directly, and it can converge to a stationary point with a PAC (Probably Approximately Correct) guarantee [28].
2) Sparse optimization. Sparse optimization, a task that is intrinsically NP-hard, is a fundamental challenge in the fields of optimization, machine learning, image processing, and control theory. One popular method is the relaxed approximation method, for example, convex relaxed method [4, 51, 44], nonconvex relaxed method [48, 15] and DC (difference of convex functions) relaxed method [21, 27]. Due to an enhanced comprehension of the variational characteristics associated with the $\ell_{0}$ norm, numerous studies have addressed sparse optimization by directly optimizing the $\ell_{0}$ norm [30, 2]. Additionally, successive coordinatewise convex problem can approximate $\ell_{0}$ norm penalty problem [42].
3) Separable constrained optimization. It aims to solve the optimization problem (12) below. The most classic method to solve this problem is the ADMM method [17], which is based on the primal-dual framework. Additionally, there exists a wide range of variations of ADMM; see, for example, the work [20, 19, 37]. Currently, some dynamical system based methods are proposed to solve optimization problem (12), for instance, first-order dynamical system [3, 5, 6], second-order dynamical system [50, 23]. In particular, [31] introduces a new primal-dual splitting algorithms with provable nonergodic convergence rates based on differential equation solver approach, which is utilized in this paper.
1.2 Contributions
This paper examines decentralized LQ optimal control with general sparsity structure. If compared with existing results, the main contributions of the presented paper are as follows.
  1. 1.
    Via the well-established concept of classic parameterization, we formulate a constrained optimization problem with a sparsity promoting penalty for the first time. Under our framework, the optimal linear feedback gain of a general decentralized LQ problem can be represented by the global minimizers of this issue in a certain sense.
  2. 2.
    Convex optimization framework introduced in [31] is feasible for the formulated optimization problem, provided that the sparsity promoting function is weighted $\ell_{1}$ norm. However, the framework of [31] cannot be utilized directly due to the presence of two distinct subproblems. The closed form solution of one subproblem is easily derived, while the remaining subproblems is complicated, necessitating the development of an iterative algorithm based on BSUM framework (see Algorithm 4). Observe that the process of BSUM entails the sequential resolution of various optimization issues, which have predefined structures yet, necessitating the need for individual design. Via the special and ingenious construction, each issue in BSUM possesses a closed form solution and BSUM exhibits $\mathcal{O}\left(\frac{1}{k}\right)$ convergence rate indeed.
  3. 3.
    Next, it is established a direct relationship between the acceleration of solving decentralized LQ problems and the design of a strongly convex sparsity promoting function. To our best knowledge, this phenomenon has not been reported in the filed of sparse optimization.
  4. 4.
    The variational properties of $\ell_{0}$-penalty is firstly studied for decentralized LQ problem. Motivated by the convex relaxation framework of [42], we show that the $\ell_{1}$-penalty induced problems can be seen as approximations of $\ell_{0}$-penalty induced problem in the context of decentralized LQ optimal control. Noting that [42] studies an unconstrained smooth optimization problem, the considered problem of this paper is constrained and nonsmooth. Hence, the conclusion and techniques of proof presented in our analysis do not exhibit parallel generalizations of those of [42].
The interconnection between the proposed optimization problems is illustrated in the diagram below.
[…]
Figure 1: Interconnection between the proposed optimization problems
[…]
For any proper closed convex function $f$ on $\mathcal{X}$, we say $f\in\mathcal{S}_{\mu}^{0}(\mathcal{X})$ with $\mu\geq 0$ if
$f(x_{1})-f(x_{2})-\langle p,x_{1}-x_{2}\rangle\geq\frac{\mu}{2}\|x_{1}-x_{2}\|^{2},\leavevmode\nobreak\ \forall(x_{1},x_{2})\in\mathcal{X}\times\mathcal{X},$
where $p\in\partial f(x_{2})$ with $\partial f(x_{2})$ the subdifferential of $f$ at $x_{2}\in\mathcal{X}$. We write $f\in\mathcal{S}_{\mu,L}^{1,1}(\mathcal{X})$ if $f\in\mathcal{S}_{\mu}^{0}(\mathcal{X})$ has $L$-Lipschitz continuous gradient:
$f(x_{1})-f(x_{2})-\langle\nabla f(x_{2}),x_{1}-x_{2}\rangle\leq\frac{L}{2}\|x_{1}-x_{2}\|^{2},\leavevmode\nobreak\ \forall(x_{1},x_{2})\in\mathcal{X}\times\mathcal{X}.$
Given
$v=\begin{pmatrix}v_{11}&\cdots&v_{1n}\\ \vdots&\ddots&\vdots\\ v_{m1}&\cdots&v_{mn}\end{pmatrix}\in\mathbb{R}^{m\times n},$
the function ${\rm vec}(v):\mathbb{R}^{m\times n}\longmapsto\mathbb{R}^{mn}$ is introduced as
${\rm vec}(v)=(v_{11},\dots,v_{1n},\dots,v_{m1},\dots,v_{mn})^{\top}.$
For any $\alpha\in\mathbb{R}$, the $\alpha$-level set of a function $f:\mathbb{R}^{n}\to[-\infty,\infty]$ is the set
${\rm lev}_{\leq\alpha}f=\{x\in\mathbb{R}^{n}\colon f(x)\leq\alpha\}.$
We introduce the notation $M\lesssim N$, meaning $M\leq CN$ with some generic bounded constant $C>0$ that is independent of $A,B,\mu_{f},\mu_{g},\gamma_{0},\beta_{0}$ (see Section 5) but can be different in each occurrence.
2 Formulation
Consider a linear time-invariant system
$\dot{x}(t)$ $=Ax(t)+B_{2}u(t)+B_{1}w(t),$ (1)
$z(t)$ $=Cx(t)+Du(t)$
with state $x(t)\in\mathbb{R}^{n}$, input $u(t)\in\mathbb{R}^{m}$, exogenous disturbance input $w(t)\in\mathbb{R}^{l}$, controlled output $z(t)\in\mathbb{R}^{q}$ and matrices $A\in\mathbb{R}^{n\times n},B_{2}\in\mathbb{R}^{n\times m},B_{1}\in\mathbb{R}^{n\times l},C\in\mathbb{R}^{q\times n},D\in\mathbb{R}^{q\times m}$. The infinite-horizon LQ problem of this paper is to find a linear static state feedback gain $K\in\mathbb{R}^{m\times n}$ such that
$u(t)=-Kx(t)$
minimizes the following performance criterion
$J=\int_{0}^{\infty}z(t)^{\top}z(t){\rm d}t.$ (2)
By [43], to optimize (2), it is equivalent to minimize the $\mathcal{H}_{2}$-norm of the transfer function
$H(s)=(C-DK)(sI_{n}-A+B_{2}K)^{-1}B_{1}$
from $w$ to $z$, and the performance criterion can be reformulated as
$J(K)=\|H(s)\|_{\mathcal{H}_{2}}^{2}={\rm Tr}\left((C-DK)W_{c}(C-DK)^{\top}\right),$
where $W_{c}$ is the controllability Gramian associated with the closed-loop system. In this paper, the stabilizing controller below refers to stabilizing system (1) internally; see Section 14 of [54] for details. The following assumption is introduced in [32, 28, 18].
Assumption 1.
Assume that $C^{\top}D=0,D^{\top}D\succ 0,B_{1}B_{1}^{\top}\succ 0,(A,B_{2})$ is stabilizable and $(A,C)$ has no unobservable modes on the imaginary axis.
We introduce the following notation
$F=\begin{bmatrix}A&B_{2}\\ 0&0\end{bmatrix}\in\mathbb{R}^{p\times p},\leavevmode\nobreak\ G=\begin{bmatrix}0\\ I_{m}\end{bmatrix}\in\mathbb{R}^{p\times m},$
$Q=\begin{bmatrix}B_{1}B_{1}^{\top}&0\\ 0&0\end{bmatrix}\in\mathbb{S}^{p},\leavevmode\nobreak\ R=\begin{bmatrix}C^{\top}C&0\\ 0&D^{\top}D\end{bmatrix}\in\mathbb{S}^{p}.$
Assumption 2.
The parameter $F$ is unknown but convex-bounded, i.e., $F$ belongs to a polyhedral domain, which is expressed as a convex combination of the extreme matrices, where $F=\sum_{i=1}^{M}\xi_{i}F_{i},\xi_{i}\geq 0,\sum_{i=1}^{M}\xi_{i}=1$, and $F_{i}=\begin{bmatrix}A_{i}&B_{2,i}\\ 0&0\end{bmatrix}\in\mathbb{R}^{p\times p}$ denotes the extreme vertex of the uncertain domain.
Let block matrix
$W=\begin{bmatrix}W_{1}&W_{2}\\ W_{2}^{\top}&W_{3}\end{bmatrix}\in\mathbb{S}^{p},$
where $W_{1}\in\mathbb{S}_{++}^{n},W_{2}\in\mathbb{R}^{n\times m},W_{3}\in\mathbb{S}^{m}$. Regard $\Theta_{i}(W)=F_{i}W+WF_{i}^{\top}+Q$ as a block matrix, i.e.,
$\Theta_{i}(W)=\begin{bmatrix}\Theta_{i,1}(W)&\Theta_{i,2}(W)\\ \Theta_{i,2}^{\top}(W)&\Theta_{i,3}(W)\end{bmatrix}\in\mathbb{S}^{p}.$
The following theorem introduces a subset of stabilizing controller gains, and find a upper bound of performance criterion (2).
Theorem 1.
One can define the set
$\mathscr{C}$ =$=\{W\in\mathbb{S}^{p}:W\succeq 0,\Theta_{i,1}(W)\preceq 0,\forall i=1,2,\ldots,M\},$
and let
$\mathscr{K}=\{K=W_{2}^{\top}W_{1}^{-1}\colon W\in\mathscr{C}\}.$
Then,
  1. 1.
    $K\in\mathscr{K}$ stabilizes the closed-loop system in the presence model uncertainties
  2. 2.
    $K\in\mathscr{K}$ gives
    $\langle R,W\rangle\geq\|H_{i}(s)\|_{\mathcal{H}_{2}}^{2},$
    where $\|H_{i}(s)\|_{\mathcal{H}_{2}}$ represents the $\mathcal{H}_{2}$-norm with respect to the $i$-th extreme system.
Proof.
The proof is familiar with that of Theorem 1 of [32], but for the sake of completeness, we will include short proof here. For statement 1, $\Theta_{i,1}\preceq 0$ implies that
$(A_{i}-B_{2,i}W_{2}^{\top}W_{1}^{-1})W_{1}+W_{1}(A_{i}-B_{2,i}W_{2}^{\top}W_{1}^{-1})^{\top}+B_{1}B_{1}^{\top}\preceq 0.$ (3)
It is obvious that $K=W_{2}^{\top}W_{1}^{-1}$ stabilizes the extreme system. To ensure the stability over the entire uncertain domain that is convex, it suffices to check the stability at the vertices of the convex polyhedron. Since, for all $i=1,\dots,M$, stability of extreme system $i$ is guaranteed, the stability of (1) holds.
For statement 2, we have
$W_{c,i}=\int_{0}^{+\infty}e^{(A-B_{2,i}K)t}B_{1}B_{1}^{\top}e^{(A-B_{2,i}K)^{\top}t}{\rm d}t.$
Hence, by (3), $W_{1}\succeq W_{c,i}$ holds. Then, based on Schur complement lemma, $W\succeq 0$ implies that
$W_{3}\succeq W_{2}^{\top}W_{1}^{-1}W_{2}=KW_{1}K^{\top}\succeq KW_{c,i}K^{\top}.$
Thus, combining with the assumption $C^{\top}D=0$, it follows that
$\langle R,W\rangle={\rm Tr}(CW_{1}C^{\top}+DW_{3}D^{\top})\geq{\rm Tr}(CW_{c,i}C^{\top}+DKW_{c,i}K^{\top}D^{\top})={\rm Tr}((C-DK)W_{c,i}(C-DK)^{\top}).$
The proof is complete. ∎
Remark 1.
The above theorem introduce a subset $\mathscr{K}$ of stabilizing feedback gain matrices and the upper bounded of performance criterion (2) in the presence of parameter uncertainties. Let us denote all the stabilizing feedback gain matrices as $\mathcal{S}$, and the we intend to solve the following optimization problem originally
$\min\limits_{K\in\mathcal{S}}J(K)={\rm Tr}\left((C-DK)W_{c}(C-DK)^{\top}\right),$ (4)
which is a constrained nonconvex problem. However, by the above theorem, we can solve the following optimization problem to approximate a “good” solution of (4)
$\min\limits_{W\in\mathscr{C}}\langle R,W\rangle,$ (5)
which is a convex problem since $\mathscr{C}$ is a convex set and objective function is linear. Thus, it seems that we obtain a “good” solution of a nonconvex problem by solving a convex optimization problem, and we have achieved substantial advancements. However, it is known that $J(K)$ is $L$-smooth and satisfies PL condition, so the gradient-based method can be utilized to obtain global minimum of (4) [10]. But if there are some sparsity constraints on $K$, the transformation of optimization problem from (4) to (5) is revolutionary, since the current state of scholarly knowledge remains uncertain regarding whether the gradient-based method could converge to the global minimum for the PL function with equality constraints.
In this study, we focus on a broad decentralized control problem, assuming that the feedback gain exhibits sparsity rather than a block diagonal structure [32]. The fundamental concept underlying our proposal is that the sparse stabilizing feedback gain, regardless of its structure, can be achieved by the design of a diagonal matrix $W_{1}$ and a sparse matrix $W_{2}$ with some sparse structure. Then, feedback gain $K=W_{2}^{\top}W_{1}^{-1}$ can preserve the sparsity structure of $W_{2}$ and stabilizes the closed-loop system.
Definition 1.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be proper and lower semi-continuous function. Then $f$ is said to be a sparsity promoting function provided that:
  1. 1.
    $f(0)=0$ and f achieves its global minimum at origin;
  2. 2.
    the set $\partial f(0)$ contains at least one nonzero element.
In this paper, we will use the weighted-$\ell_{1}$ norm, a piecewise quadratic function, and the $\ell_{0}$ norm as sparsity promoting functions, which correspond to distinct optimization problems. First, for the weighted-$\ell_{1}$ norm of a matrix $F$
$g(F)=\sum_{i,j}w_{ij}|F_{ij}|,$ (6)
introduce the following optimization problem
$\min_{W,P}\leavevmode\nobreak\ \leavevmode\nobreak$ $\langle R,W\rangle+\gamma g(P)$ (7)
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $W\in\mathbb{S}^{p}_{+},$
$-V_{2}(F_{i}W+WF_{i}^{\top}+Q)V_{2}^{\top}\in\mathbb{S}_{+}^{n},\forall i=1,\dots,M,$
$-V_{j1}WV_{j2}=0,\forall j=1,\dots,\frac{n(n-1)}{2},$
$V_{1}WV_{2}^{\top}-P=0$
with $\gamma>0$, $V_{1}=[0,I_{m}],V_{2}=[I_{n},0]$. Denote
$\Psi_{i}=-V_{2}(F_{i}W+WF_{i}^{\top}+Q)V_{2}^{\top},\forall i=1,\dots,M.$
Then, optimization problem (7) can be equivalently expressed in a compact form
$\min_{W,P}\leavevmode\nobreak\ \leavevmode\nobreak$ $\langle R,W\rangle+\gamma g(P)$ (8)
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $\mathcal{G}(W)\in\mathcal{K},$
$V_{1}WV_{2}^{\top}-P=0,$
where
$\mathcal{G}(W)=(W,\Psi_{1},\dots,\Psi_{M},V_{11}WV_{12},\dots,V_{N1}WV_{N2}),$
and convex cone $\mathcal{K}$ is defined as
$\mathcal{K}=\mathbb{S}_{+}^{p}\times\mathbb{S}_{+}^{n}\times\cdots\times\mathbb{S}_{+}^{n}\times 0\times\cdots\times 0.$
Since the positive semi-definite cone is self-dual, it is straightforward to express the dual cone $\mathcal{K}^{*}$ as
$\mathcal{K}^{*}=\mathbb{S}_{+}^{p}\times\mathbb{S}_{+}^{n}\times\cdots\times\mathbb{S}_{+}^{n}\times\mathbb{R}\times\cdots\times\mathbb{R}.$
Assumption 3.
Problem (8) is strictly feasible, i.e., Slater’s condition is satisfied and strong duality holds.
Remark 2.
This assumption is popular in solving decentralized LQ problems via optimization landscape [32, 49, 52]. The question of when the aforementioned assumption is valid remains unresolved, and [52] offers some insights from the standpoint of block-diagonal strategy when $(A,B_{2})$ is known. Specifically, let $D$ be the feasible set of optimization problem (8), and denote
$\widetilde{D}=\{W_{2}\in\mathbb{R}^{n\times m}\colon W_{2}={\rm blocdiag}\{W_{2,1},\dots,W_{2,n}\}\}$
with $W_{2,j}\in\mathbb{R}^{1\times D_{j}}$ for all $j=1,\dots,n$. It is demonstrate that, if system (1) is fully actuated or weakly coupled in terms of topological connections or weakly coupled in terms of dynamical interactions (see Definition 4, Definition 5 and Definition 8 of [52]), then $D\cap\widetilde{D}$ (a subset of $D$) is nonempty (see Proposition 1-3 of [52]). Hence, if system (1) is fully actuated or weakly coupled in terms of topological connections or weakly coupled in terms of dynamical interactions, problem (8) is feasible. Combining with the fact that $B_{1}B_{1}^{\top}\succ 0$ and $W_{3}$ is a free variable to be chosen such that $W\succ 0$ (utilizing Schur complement lemma), problem (8) is strictly feasible, i.e., Assumption 3 holds. It is important to note that the feasible set of (8) is larger than the ones in [32, 49, 18, 52], since there is no structural constraints on $W_{2}$ in this paper.
Remark 3.
In fact, let
$D_{k}=\begin{bmatrix}0\\ I_{k}\end{bmatrix}\in\mathbb{R}^{n\times k},$
then the $\frac{n(n-1)}{2}$ equality constraints
$V_{j1}WV_{j2}=0,j=1,\dots,\frac{n(n-1)}{2}$ (9)
can be converted into $n-1$ equality constraint:
$e_{k}V_{2}WV_{2}^{\top}D_{n-k}=0,k=1,\dots,n-1$
with $e_{k}=(0,\dots,1,\dots,0)$. For the convenience and simplicity of formulas’ deducing, we still use $\frac{n(n-1)}{2}$ equality constraints (i.e., (9)) to describe the diagonal property of matrix $W_{1}$.
In Section 3, we will introduce a two-timescale algorithm framework based on differential equation solver approach to address optimization problem (8), which encompasses two distinct subproblems with one possessing closed form solution. The other subproblem is investigated in Section 4, and Section 5 provides the convergence analysis of the proposed algorithm and studies an acceleration optimization problem. In Section 6, we will study an $\ell_{0}$ sparsity promoting optimization problem that differs from (8), and propose some theorems to characterize its theoretical features.
3 Two-Timescale Optimization Algorithm Framework
Review the problem (8)
$\min_{W,P}\leavevmode\nobreak\ \leavevmode\nobreak$ $\langle R,W\rangle+\gamma g(P)$
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $\mathcal{G}(W)\in\mathcal{K},$
$V_{1}WV_{2}^{\top}-P=0.$
Let $\widetilde{W}={\rm vec}(W),\widetilde{P}={\rm vec}(P)$, then (8) is equivalent to the following optimization problem
$\min_{\widetilde{W},\widetilde{P}}\leavevmode\nobreak\ \leavevmode\nobreak$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})+\gamma g(P)$ (10)
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $\left(V_{i2}^{\top}\otimes V_{i1}\right)\widetilde{W}=0,\leavevmode\nobreak\ \leavevmode\nobreak\ i=1,\dots,M,$
$(V_{2}\otimes V_{1})\widetilde{W}-\widetilde{P}=0,$
where $\widetilde{\Psi}_{i}={\rm vec}(\Psi_{i})$ and
$\Gamma_{+}^{p}=\left\{{\rm vec}(W)\colon W\in\mathbb{S}_{+}^{p}\right\}.$
If denote $\mathcal{A}\in\mathbb{R}^{(M+mn)\times p^{2}},\mathcal{B}\in\mathbb{R}^{(M+mn)\times mn}$ with
$\mathcal{A}=\begin{pmatrix}V_{12}^{\top}\otimes V_{11}\\ \vdots\\ V_{M2}^{\top}\otimes V_{M1}\\ V_{2}\otimes V_{1}\end{pmatrix},\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}=\begin{pmatrix}0\\ \vdots\\ 0\\ -I_{mn}\end{pmatrix},$
then optimization problem (10) is summarized in the following form
$\min_{\widetilde{W},\widetilde{P}}\leavevmode\nobreak\ \leavevmode\nobreak$ $f(\widetilde{W})+h(\widetilde{P})$ (11)
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $\mathcal{A}\widetilde{W}+\mathcal{B}\widetilde{P}=0,$
where
$f(\widetilde{W})$ $=\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M}),$
$h(\widetilde{P})$ $=\gamma g(P).$
To solve optimization problem (11), we briefly review the basic setup of primal-dual splitting method based on differential equation approach; see [31] for a comprehensive review. The following separable convex optimization problem is considered in [31]:
$\min_{x\in\mathcal{X},y\in\mathcal{Y}}F(x,y)=f(x)+g(y)\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm s.t.}\leavevmode\nobreak\ Ax+By=b,$ (12)
where $\mathcal{X}\subset\mathbb{R}^{m}$ and $\mathcal{Y}\subset\mathbb{R}^{n}$ are two closed convex sets, and $f$ and $g$ are two properly closed convex functions (not necessary differential). Assume that $g\in\mathcal{S}_{\mu_{g}}^{0}(\mathcal{Y})$ with $\mu_{g}>0$ and $f=f_{1}+f_{2}$ with $f_{2}\in\mathcal{S}_{0}^{0}(\mathcal{X})$, $f_{1}\in\mathcal{S}_{\mu_{f},L_{f}}^{1,1}(\mathcal{X})$ and $0\leq\mu_{f}\leq L_{f}<\infty$. Then, the following scheme is proposed to solve the above equality constrained optimization problem
[…]
where $(\tilde{v}_{k},\tilde{y}_{k},\tilde{\eta}_{f,k},\eta_{g,k})$ is given by
$\eta_{g,k}$ $=(\alpha_{k}+1)\beta_{k}+\mu_{g}\alpha_{k},$ (14)
$\tilde{y}_{k}$ $=y_{k}+\frac{\alpha_{k}\beta_{k}}{\eta_{g,k}}(w_{k}-y_{k}),$ (15)
$\tilde{\eta}_{f,k}$ $=\gamma_{k}+\mu_{f}\alpha_{k},$ (16)
$\tilde{v}_{k}$ $=\frac{1}{\tilde{\eta}_{f,k}}(\gamma_{k}v_{k}+\mu_{f}\alpha_{k}u_{k}).$ (17)
To emphasis, $\alpha_{k}>0$ is the step size and the parameter system is discretized implicitly by
$\frac{\theta_{k+1}-\theta_{k}}{\alpha_{k}}=-\theta_{k+1},\quad\frac{\gamma_{k+1}-\gamma_{k}}{\alpha_{k}}=\mu_{f}-\gamma_{k+1},\quad\frac{\beta_{k+1}-\beta_{k}}{\alpha_{k}}=\mu_{g}-\beta_{k+1}$ (18)
with initial conditions $\theta_{0}=1,\gamma_{0}>0$ and $\beta_{0}>0$.
Remark 4.
In fact, the following differential inclusion
[…]
is the continuous-time version of algorithm (13), where $\mathcal{L}$ is defined in (39) below (see Section 5). It can be observed that differential inclusion (19) encompasses two Nesterov-type differential inclusions (studied in [47]) pertaining to primal variables $x$ and $y$, as well as a gradient flow about the dual variable $\lambda$. Besides, an alternative presentation of (19) reads as
[…]
Therefore, (13) represents the semi-implicit Euler discretization of differential inclusion (20), which justifies its classification as a differential equation solver approach.
Remark 5.
[…]
Back to our topic (optimization problem (11)), denote
$f_{1}(\widetilde{W})$ $=\langle{\rm vec}(R),\widetilde{W}\rangle,$
$f_{2}(\widetilde{W})$ $=\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M}).$
We first consider the iteration of variable $\widetilde{P}\in\mathbb{R}^{mn\times 1}$. By (13),
$\widetilde{P}_{k+1}$ $=\mathbf{prox}_{\tau_{k}h}^{\mathbb{R}^{mn\times 1}}\left(\widetilde{P}_{k}+\frac{\alpha_{k}\beta_{k}}{\eta_{g,k}}(w_{k}-\widetilde{P}_{k})-\tau_{k}\mathcal{B}^{\top}\bar{\lambda}_{k+1}\right)$ (21)
$=\left(\underset{P_{ij}\in\mathbb{R}}{\operatorname*{argmin}}\left\{\sum_{i,j}\left(\gamma w_{ij}|P_{ij}|+\frac{\rho_{k}}{2}\left(P_{ij}-\Omega^{k}_{ij}\right)^{2}\right)\right\}\right)_{mn\times 1},$ (22)
where $\rho_{k}=\frac{1}{\tau_{k}}$ and
$\Omega^{k}={\operatorname*{matrix}}\left(\widetilde{P}_{k}+\frac{\alpha_{k}\beta_{k}}{\eta_{g,k}}(w_{k}-\widetilde{P}_{k})-\tau_{k}\mathcal{B}^{\top}\bar{\lambda}_{k+1}\right).$ (23)
Letting $\theta(\widetilde{P})=\sum_{i,j}\left(\gamma w_{ij}|P_{ij}|+\frac{\rho_{k}}{2}\left(P_{ij}-\Omega^{k}_{ij}\right)^{2}\right)$, it is obvious that
$\partial_{P_{ij}}\theta(\widetilde{P})=\left\{\begin{aligned} &\gamma w_{ij}[-1,1]+\rho_{k}\left(P_{ij}-\Omega^{k}_{ij}\right),\leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ P_{ij}=0,\\ &\gamma w_{ij}{\operatorname*{sign}}(P_{ij})+\rho_{k}\left(P_{ij}-\Omega^{k}_{ij}\right),\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm otherwise}.\end{aligned}\right.$
Denoting $a_{k}=\frac{\gamma}{\rho_{k}}w_{ij}$, we have
$P_{ij}^{*}=\left\{\begin{aligned} &\left(1-\frac{a_{k}}{|\Omega_{ij}^{k}|}\right)\Omega_{ij}^{k},\leavevmode\nobreak\ \leavevmode\nobreak\ &|\Omega_{ij}^{k}|>a_{k},\\ &0,\leavevmode\nobreak\ \leavevmode\nobreak\ &|\Omega_{ij}^{k}|\leq a_{k}.\end{aligned}\right.$ (24)
Hence, $\widetilde{P}$ has naturally sparsity structure.
4 Solving Inner Layer Optimization Problem
Algorithm (13) consists of two subproblems: updating $v_{k+1}$ and $y_{k+1}$. The solution to the latter can be obtained explicitly using equation (24), while the solution to the former will be discussed in this section. We first deduce the Lagrange dual problem of the above subproblem due to the appealing structure of dual problem. Then, the BSUM framework is utilized to address the dual problem. Observe that the process of BSUM entails the sequential resolution of various optimization issues. However, each problem has a predefined structure, necessitating the need for individual design. Via the special and ingenious construction, each issue possesses a closed form solution and BSUM exhibits $\mathcal{O}\left(\frac{1}{k}\right)$ convergence rate.
The update of $v_{k+1}$ is given by
$v_{k+1}=\underset{v\in\mathcal{X}}{\operatorname*{argmin}}\left\{f_{2}(v)+\langle d_{k},v\rangle+\frac{\alpha_{k}}{2\theta_{k}}\left\|Av+Bw_{k}-b\right\|^{2}+\frac{\widetilde{\eta}_{f,k}}{2\alpha_{k}}\left\|v-\widetilde{v}_{k}\right\|^{2}\right\},$ (25)
and when employing the symbol system of this paper, optimization problem (25) is transformed to
$\min_{{\rm vec}(W)}\leavevmode\nobreak$ $\langle d_{k},{\rm vec}(W)\rangle+\frac{\alpha_{k}}{2\theta_{k}}\left\|\mathcal{A}{\rm vec}(W)+\mathcal{B}w_{k}\right\|^{2}+\frac{\widetilde{\eta}_{f,k}}{2\alpha_{k}}\left\|{\rm vec}(W)-\widetilde{v}_{k}\right\|^{2}$ (26)
${\rm s.t.}\leavevmode\nobreak$ ${\rm vec}(W)\in\Gamma_{+}^{p},$
${\rm vec}(\Psi_{i})\in\Gamma_{+}^{n},\leavevmode\nobreak\ i=1,\dots,M,$
where
${\rm vec}(\Psi_{i})$ $={\rm vec}(-VF_{i}WV^{\top}-VWF_{i}^{\top}V^{\top}-VQV^{\top})$
$=-\left((V\otimes VF_{i})+(VF_{i}\otimes V)\right){\rm vec}(W)-(V\otimes V){\rm vec}(Q).$
By Assumption 3, for all $k$, problem (26) satisfies Slater’s condition and strong duality holds. Since $W,\Psi_{i}$ are symmetric matrices, there exist matrices $T_{1}\in\mathbb{R}^{\frac{p(p+1)}{2}\times p^{2}}$ and $T_{2}\in\mathbb{R}^{\frac{n(n+1)}{2}\times n^{2}}$ such that
${\rm svec}(W)$ $=T_{1}{\rm vec}(W)\leavevmode\nobreak\ {\rm with}\leavevmode\nobreak\ T_{1}T_{1}^{\top}=I_{\frac{p(p+1)}{2}},$
${\rm svec}(\Psi_{i})$ $=T_{2}{\rm vec}(\Psi_{i})\leavevmode\nobreak\ {\rm with}\leavevmode\nobreak\ T_{2}T_{2}^{\top}=I_{\frac{n(n+1)}{2}},$
where svec is the vectorization operator on $\mathbb{S}^{n}$ defined by
${\rm svec}(S)=\left[s_{11},s_{12},\cdots,s_{n1},\cdots,s_{nn}\right]^{\top}\in\mathbb{R}^{\frac{n(n+1)}{2}}.$
Remark 6.
We partition the identity matrix $I_{\frac{n(n+1)}{2}}$ as follows
$I_{\frac{n(n+1)}{2}}=(u_{11},u_{21},\dots,u_{n1},u_{22},\dots,u_{n2},\dots,u_{nn}).$
Then, for every matrix $A\in\mathbb{S}^{n}$, the elimination matrix $L$ (which performs the transformation ${\rm svec}(A)=L{\rm vec}(A)$) can be explicitly expressed as
$L=\sum_{i\geq j}u_{ij}({\rm vec}E_{ij})^{\top}=\sum_{i\geq j}(u_{ij}\otimes e_{j}^{\top}\otimes e_{i}^{\top})$
with $E_{ij}=e_{i}e_{j}^{\top}$. By Lemma 3.2 of [33], $L^{\top}=L^{\dagger}$, and naturally $LL^{\top}=I_{\frac{n(n+1)}{2}}$.
Remark 7.
Inversely, there exist matrix $D_{1}\in\mathbb{R}^{p^{2}\times\frac{p(p+1)}{2}}$ and $D_{2}\in\mathbb{R}^{n^{2}\times\frac{n(n+1)}{2}}$ such that
${\rm vec}(W)=D_{1}{\rm svec}(W),\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm vec}(\Psi_{i})=D_{2}{\rm svec}(\Psi_{i}).$
In addition, $D_{1}$ can be explicitly expressed as
$D_{1}=\sum_{i\geq j}u_{ij}{\rm vec}(T_{ij})^{\top},$
where $T_{ij}$ is a $p\times p$ matrix with $1$ in position $(i,j),(j,i)$ and zero elsewhere.
Then, by utilizing svec operator and denoting ($\widetilde{\Gamma}_{+}^{n}$ is defined analogously)
$\widetilde{\Gamma}_{+}^{p}=\left\{{\rm svec}(W)\colon W\in\mathbb{S}_{+}^{p}\right\},$
optimization (26) can be readily converted into the subsequent problem
$\min_{{\rm svec}(W)}\leavevmode\nobreak$ $\langle d_{k},D_{1}{\rm svec}(W)\rangle+\frac{\alpha_{k}}{2\theta_{k}}\left\|\mathcal{A}D_{1}{\rm svec}(W)+\mathcal{B}w_{k}\right\|^{2}+\frac{\widetilde{\eta}_{f,k}}{2\alpha_{k}}\left\|D_{1}{\rm svec}(W)-\widetilde{v}_{k}\right\|^{2}$ (27)
${\rm s.t.}\leavevmode\nobreak$ ${\rm svec}(W)\in\widetilde{\Gamma}_{+}^{p}$
${\rm svec}(\Psi_{i})\in\widetilde{\Gamma}_{+}^{n},\leavevmode\nobreak\ i=1,\dots,M,$
where ${\rm svec}(\Psi_{i})=T_{2}{\rm vec}(\Psi_{i}).$ The prime problem of (27) presents a level of difficulty that warrants exploration of its dual problem. The Lagrangian of (27) is defined as
$\mathcal{L}({\rm svec}(W);X)=$ $\langle d_{k},D_{1}{\rm svec}(W)\rangle+\frac{\alpha_{k}}{2\theta_{k}}\left\|\mathcal{A}D_{1}{\rm svec}(W)+\mathcal{B}w_{k}\right\|^{2}+\frac{\widetilde{\eta}_{f,k}}{2\alpha_{k}}\left\|D_{1}{\rm svec}(W)-\widetilde{v}_{k}\right\|^{2}$ (28)
$-\left\langle X_{0},{\rm svec}(W)\right\rangle+\left\langle(X_{1}+\dots+X_{M}),T_{2}(V\otimes V){\rm vec}(Q)\right\rangle$
$+\sum_{i=1}^{M}\left\langle T_{2}\left((V\otimes VF_{i})+(VF_{i}\otimes V)\right)D_{1}{\rm svec}(W),X_{i}\right\rangle,$
where $X=(X_{0},X_{1},\dots,X_{M})\in\widetilde{\Gamma}_{+}^{p}\times\widetilde{\Gamma}_{+}^{n}\times\dots\times\widetilde{\Gamma}_{+}^{n}$ is the Lagrange multiplier. Define
$\tilde{d}_{k}=D_{1}^{\top}d_{k}-X_{0}+D_{1}^{\top}\sum_{i=1}^{M}\left[(V\otimes VF_{i})^{\top}+(VF_{i}\otimes V)^{\top}\right]T_{2}^{\top}X_{i}\doteq\mathcal{L}_{k}(X_{0},X_{1},\dots,X_{M}),$
$\sigma_{1}=\frac{\alpha_{k}}{2\theta_{k}},\leavevmode\nobreak\ \sigma_{2}=\frac{\tilde{\eta}_{f,k}}{2\alpha_{k}},\leavevmode\nobreak\ \tilde{b}_{k}=\mathcal{B}w_{k},$
and by rearranging (28) we can derive
$\mathcal{L}({\rm svec}(W);X)=$ $\langle\tilde{d}_{k},{\rm svec}(W)\rangle+\sigma_{1}\left\|\mathcal{A}D_{1}{\rm svec}(W)+\tilde{b}_{k}\right\|^{2}+\sigma_{2}\left\|D_{1}{\rm svec}(W)-\widetilde{v}_{k}\right\|^{2}$
$+\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle.$
It may be inferred that the Lagrange dual function $\theta(X)$ is derived by
$\theta(X)=\min_{{\rm svec}(W)}\mathcal{L}({\rm svec}(W);X)=\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle+\min_{{\rm svec}(W)}\hat{\theta}({\rm svec}(W))$ (29)
with
$\hat{\theta}({\rm svec}(W))=\langle\tilde{d}_{k},{\rm svec}(W)\rangle+\sigma_{1}\left\|\mathcal{A}D_{1}{\rm svec}(W)+\tilde{b}_{k}\right\|^{2}+\sigma_{2}\left\|D_{1}{\rm svec}(W)-\widetilde{v}_{k}\right\|^{2}.$
Let
$\nabla\hat{\theta}({\rm svec}(W))=\tilde{d}_{k}+2\sigma_{1}(\mathcal{A}D_{1})^{\top}(\mathcal{A}D_{1}{\rm svec}(W)+\tilde{b}_{k})+2\sigma_{2}D_{1}^{\top}(D_{1}{\rm svec}(W)-\tilde{v}_{k})=0.$
Alternatively, it may be derived that
$\left(2\sigma_{1}(\mathcal{A}D_{1})^{\top}(\mathcal{A}D_{1})+2\sigma_{2}D_{1}^{\top}D_{1}\right){\rm svec}(W)=-2\sigma_{1}(\mathcal{A}D_{1})^{\top}\tilde{b}_{k}+2\sigma_{2}D_{1}^{\top}\tilde{v}_{k}-\tilde{d}_{k}.$
Denoting
$M=2\sigma_{1}(\mathcal{A}D_{1})^{\top}(\mathcal{A}D_{1})+2\sigma_{2}D_{1}^{\top}D_{1},\leavevmode\nobreak\ q_{k}=-2\sigma_{1}(\mathcal{A}D_{1})^{\top}\tilde{b}_{k}+2\sigma_{2}D_{1}^{\top}\tilde{v}_{k},$
we can obtain
${\rm svec}(W)=M^{-1}(q_{k}-\tilde{d}_{k}).$ (30)
Substituting (30) into (29), it could potentially inferred that
$\theta(X)$ $=\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle+\left\langle\tilde{d}_{k},M^{-1}(q_{k}-\tilde{d}_{k})\right\rangle$
$+\sigma_{1}\left\|\mathcal{A}D_{1}M^{-1}(q_{k}-\tilde{d}_{k})+\tilde{b_{k}}\right\|^{2}+\sigma_{2}\left\|D_{1}M^{-1}(q_{k}-\tilde{d}_{k})-\tilde{v}_{k}\right\|^{2}$
$=\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle+\tilde{d}_{k}^{\top}\left(\sigma_{1}(\mathcal{A}D_{1}M^{-1})^{\top}(\mathcal{A}D_{1}M^{-1})+\sigma_{2}M^{-2}-M^{-1}\right)\tilde{d}_{k}$
$+\left\langle\tilde{d}_{k},M^{-1}q_{k}+2\sigma_{1}(\mathcal{A}D_{1}M^{-1})^{\top}\xi_{k}^{1}+2\sigma_{2}(D_{1}M^{-1})^{\top}\xi_{k}^{2}\right\rangle+\|\xi_{k}^{1}\|^{2}+\|\xi_{k}^{2}\|^{2},$
where
$\xi_{k}^{1}=-\mathcal{A}D_{1}M^{-1}q_{k}-\tilde{b}_{k},\leavevmode\nobreak\ \xi_{k}^{2}=\tilde{v}_{k}-D_{1}M^{-1}q_{k}.$
Denoting
$\Omega_{k}$ $=M^{-1}q_{k}+2\sigma_{1}(\mathcal{A}D_{1}M^{-1})^{\top}\xi_{k}^{1}+2\sigma_{2}(D_{1}M^{-1})^{\top}\xi_{k}^{2},$
$\Psi$ $=\sigma_{1}(\mathcal{A}D_{1}M^{-1})^{\top}(\mathcal{A}D_{1}M^{-1})+\sigma_{2}M^{-2}-M^{-1},$
the formulation of the dual problem is presented in the subsequent manner:
$\max_{X=(X_{0},\dots,X_{M})}\leavevmode\nobreak$ $\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle+\left\langle\tilde{d}_{k},\Omega_{k}\right\rangle+\tilde{d}_{k}^{\top}\Psi\tilde{d}_{k}$ (31)
${\rm s.t.}\leavevmode\nobreak$ $X_{0}\in\widetilde{\Gamma}_{+}^{p},$
$X_{i}\in\widetilde{\Gamma}_{+}^{n},\leavevmode\nobreak\ i=1,\dots,M.$
In the following text, the sGS (symmetric Gauss-Seidel) semi-proximal ALM (Augmented Lagrangian Method) is presented to solve dual problem (31). Define
$\Theta(X_{0},X_{1},\dots,X_{M})$ $=-\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle-\left\langle\mathcal{L}_{k}(X_{0},\dots,X_{M}),\Omega_{k}\right\rangle$
$+\left\|\sqrt{-\Psi}\mathcal{L}_{k}(X_{0},\dots,X_{M})\right\|^{2}+\delta_{\widetilde{\Gamma}_{+}^{p}}(X_{0})+\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{1})+\dots+\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{M}),$
and dual problem (31) can be equivalently expressed as
$\min_{X=(X_{0}\dots,X_{M})}\leavevmode\nobreak\ \Theta(X_{0},X_{1},\dots,X_{M}).$ (32)
For $i=1,\dots,M$, we can compute that
$\partial_{X_{i}}\Theta(X_{0},X_{1},\dots,X_{M})$ $=-T_{2}(V\otimes V){\rm vec}(Q)-T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}\Omega_{k}$
$+2T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}(-\Psi)\mathcal{L}_{k}(X_{0},\dots,X_{i-1},X_{i+1},\dots,X_{M})$
$+2T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}(-\Psi)D_{1}^{\top}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]^{\top}T_{2}^{\top}X_{i}$
$+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{i}).$
Define
$H_{i}(X_{0},\dots,X_{i-1},X_{i+1},\dots,X_{M})=T_{2}(V\otimes V){\rm vec}(Q)+T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}\Omega_{k}$
$-2T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}(-\Psi)\mathcal{L}_{k}(X_{0},\dots,X_{i-1},X_{i+1},\dots,X_{M}),$
$L_{i}=2T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}(-\Psi)D_{1}^{\top}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]^{\top}T_{2}^{\top},$
then we have
$\partial_{X_{i}}\Theta(X_{0},X_{1},\dots,X_{M})=-H_{i}(X_{0},\dots,X_{i-1},X_{i+1},\dots,X_{M})+L_{i}X_{i}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{i}).$
The subproblem in terms of the variable $X_{i}$ in the backward sGS sweep is given by
$\bar{X}_{i}^{k+1}=\underset{X_{i}}{\operatorname*{argmin}}\leavevmode\nobreak\ \Theta(X_{0}^{k},X_{1}^{k},\dots,X_{i-1}^{k},X_{i},\bar{X}_{i+1}^{k+1},\dots,\bar{X}_{M}^{k+1})+\frac{1}{2}\left\|X_{i}-X_{i}^{k}\right\|_{\mathcal{S}_{i}}^{2},$
where $\mathcal{S}_{i}$ is a positive linear operator given by
$\mathcal{S}_{i}(X)=\rho_{i}\mathcal{I}(X)-L_{i}X,\leavevmode\nobreak\ \rho_{i}=\max\{{\rm eig}(L_{i})\}$
with an identity operator $\mathcal{I}$.
The optimality condition implies that
00 0 $\in-H_{i}(X_{0}^{k},X_{1}^{k},\dots,X_{i-1}^{k},\bar{X}_{i+1}^{k+1},\dots,\bar{X}_{M}^{k+1})+L_{i}X_{i}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{i})+\mathcal{S}_{i}(X_{i}-X_{i}^{k})$
$=\rho_{i}X_{i}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{i})-H_{i}(X_{0}^{k},X_{1}^{k},\dots,X_{i-1}^{k},\bar{X}_{i+1}^{k+1},\dots,\bar{X}_{M}^{k+1})-\rho_{i}X_{i}^{k}+L_{i}X_{i}^{k}.$
Denoting
$\Delta_{i}^{k}=\rho_{i}^{-1}(H_{i}(X_{0}^{k},X_{1}^{k},\dots,X_{i-1}^{k},\bar{X}_{i+1}^{k+1},\dots,\bar{X}_{M}^{k+1})+\rho_{i}X_{i}^{k}-L_{i}X_{i}^{k}),$
upon careful observation, it may be noted that
$\bar{X}_{i}^{k+1}=(I+\rho_{i}^{-1}\partial\delta_{\widetilde{\Gamma}_{+}^{n}})^{-1}(\Delta_{i}^{k})=\Pi_{\widetilde{\Gamma}_{+}^{n}}(\Delta_{i}^{k}),$ (33)
where the second equality is justified by the lemma provided.
Lemma 1 ([32]).
The projection operator $\Pi_{\mathcal{C}}(\cdot)$ with respect to the convex cone $\mathcal{C}$ can be expressed as
$\Pi_{\mathcal{C}}=(I+\alpha\partial\delta_{\mathcal{C}})^{-1},$
where ($\partial$($\cdot$)∂ ( ⋅ ) denotes the subdifferential operator, and $\alpha\in\mathbb{R}$ can be an arbitrary number.
For $i=0$, we can compute that
$\partial_{X_{0}}\Theta(X_{0},\dots,X_{M})=L_{0}X_{0}+\partial\delta_{\widetilde{\Gamma}_{+}^{p}}(X_{0})-H_{0}(X_{1},\dots,X_{M}),$
with
$L_{0}=-2\Psi,\leavevmode\nobreak\ H_{0}(X_{1},\dots,X_{M})=-2\Psi\mathcal{L}_{k}(X_{1},\dots,X_{M})-\Omega_{k}.$
Similarly, the subproblem in terms of the variable $X_{0}$ in the backward sGS sweep is given by
$X_{0}^{k+1}=\underset{X_{0}}{\operatorname*{argmin}}\leavevmode\nobreak\ \Theta(X_{0},\bar{X}_{1}^{k+1},\dots,\bar{X}_{M}^{k+1})+\frac{1}{2}\left\|X_{0}-X_{0}^{k}\right\|_{\mathcal{S}_{0}}^{2},$
where the positive linear operator $\mathcal{S}_{0}$ is given by
$\mathcal{S}_{0}(X)=\rho_{0}\mathcal{I}(X)-L_{0}X,\leavevmode\nobreak\ \rho_{0}=\max\{{\rm eig}(L_{0})\}.$
Then, by the same token, we have
$X_{0}^{k+1}=(I+\rho_{0}^{-1}\partial\delta_{\widetilde{\Gamma}_{+}^{p}})^{-1}(\Delta_{0}^{k})=\Pi_{\widetilde{\Gamma}_{+}^{p}}(\Delta_{0}^{k})$ (34)
with
$\Delta_{0}^{k}=\rho_{0}^{-1}(H_{0}(\bar{X}_{1}^{k+1},\dots,\bar{X}_{M}^{k+1})+\rho_{0}X_{0}^{k}-L_{0}X_{0}^{k}).$
For forward sGS sweep, for all $i=1,\dots,M$, $X_{i}^{k+1}$ is given by
$X_{i}^{k+1}=\Pi_{\widetilde{\Gamma}_{+}^{n}}(\tilde{\Delta}_{i}^{k}),$ (35)
where $\tilde{\Delta}_{i}^{k}$ is defined as
$\tilde{\Delta}_{i}^{k}=\alpha_{i}^{-1}(H_{i}(X_{0}^{k+1},X_{1}^{k+1},\dots,X_{i-1}^{k+1},\bar{X}_{i+1}^{k+1},\dots,\bar{X}_{M}^{k+1})+\rho_{i}\bar{X}_{i}^{k+1}-L_{i}\bar{X}_{i}^{k+1}).$
Remark 8.
The update of $X_{i}^{k+1}$ requires the projection onto the convex cone $\widetilde{\Gamma}_{+}^{n}$ for $i=1,\dots,M$. Obviously,
$\bar{X}_{i}^{k+1}=\Pi_{\widetilde{\Gamma}_{+}^{n}}(\Delta_{i}^{k})=\Pi_{\mathbb{S}_{+}^{n}}({\rm svec}^{-1}(\Delta_{i}^{k})),$
and obtaining an explicit solution appears to be challenging. However, the subsequent lemma provides a computationally efficient method for calculating the aforementioned projection.
Lemma 2.
Let $X=\sum_{i=1}^{n}\lambda_{i}v_{i}v_{i}^{\top}\in\mathbb{S}^{n}$ be the eigenvalue decomposition of matrix $X$ with the eigenvalues satisfying $\lambda_{1}\geq\dots\geq\lambda_{n}$, where $v_{i}$ denotes the eigenvector corresponding to the $i$-th eigenvalue. Then, the projection onto the positive semi-definite cone of matrix X can be expressed by
$\Pi_{\mathbb{S}_{+}^{n}}(X)=\sum_{i=1}^{n}\max\{\lambda_{i},0\}v_{i}v_{i}^{\top}.$
Introducing the proximal term is first proposed in [12], and the convergence analysis has been considered in [22]. By utilizing the proximal term $\|X_{i}-X_{i}^{k}\|_{\mathcal{S}_{i}}^{2}$, it is possible to derive the closed form expression for the optimization variable $\bar{X}_{i}^{k+1}$. However, a fuller understanding of the proximal term can be achieved by examining it from the perspective of auxiliary functions [22].
Definition 2.
Consider the following optimization problem
$\min f(x)=g(x_{1},\dots,x_{K})+\sum_{i=1}^{K}h_{i}(x_{i})\leavevmode\nobreak\ {\rm s.t.}\leavevmode\nobreak\ x_{i}\in\mathcal{X}_{i},\leavevmode\nobreak\ i=1,\dots,K,$
where $g(\cdot)$is a smooth convex function and $h_{i}$ is a closed convex function. We call that $\{u_{i}(\cdot,\cdot)\}$ is strict auxiliary function, if
  1. 1.
    $u_{i}(y_{i},y)=g(y)\leavevmode\nobreak\ \leavevmode\nobreak\ \forall y\in\mathcal{X}=\mathcal{X}_{1}\times\dots\times\mathcal{X}_{K},\forall i$,
  2. 2.
    $u_{i}(x_{i},y)\geq g(y_{1},\dots,y_{i-1},x_{i},y_{i+1},\dots,y_{n})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall x_{i}\in\mathcal{X}_{i},\forall y\in\mathcal{X},\forall i,$
  3. 3.
    $\nabla u_{i}(y_{i},y)=\nabla_{i}g(y),\forall y\in\mathcal{X},\forall i$,
  4. 4.
    $u_{i}(x_{i},y)$ is continuous in $(x_{i},y),\forall i$. Further, for any given $y\in\mathcal{X}$, it is proper, closed and strongly convex function of $x_{i}$, satisfying
    $u_{i}(x_{i},y)\geq u_{i}(\hat{x}_{i},y)+\left\langle\nabla u_{i}(\hat{x}_{i},y),x_{i}-\hat{x}_{i}\right\rangle+\frac{\gamma_{i}}{2}\|x_{i}-\hat{x}_{i}\|^{2},\forall x_{i},\hat{x}_{i}\in\mathcal{X}_{i},$
    where $\gamma_{i}>0$ is independent of the choice of $y$.
  5. 5.
    For any given $y\in\mathcal{X}$, $u_{k}(x_{i},y)$ has Lipschitz continuous gradient, i.e.,
    $\|\nabla u_{i}(x_{i},y)-\nabla u_{i}(\hat{x}_{i},y)\|\leq L_{i}\|x_{k}-\hat{x}_{i}\|,\leavevmode\nobreak\ \forall\hat{x}_{i},x_{i}\in\mathcal{X}_{i},\forall i,$
    where $L_{i}>0$ is some constant. Further, we have
    $\|\nabla u_{i}(x_{i},y)-\nabla u_{i}(x_{i},z)\|\leq G_{i}\|y-z\|,\leavevmode\nobreak\ \forall x_{i}\in\mathcal{X}_{i},\forall i,\forall y,z\in\mathcal{X}.$
    Define $L_{{\rm max}}=\max_{i}L_{i},G_{{\rm max}}=\max_{i}G_{i}$.
Denote
$g(X_{0},\dots,X_{M})=\left\|\sqrt{-\Psi}\mathcal{L}_{k}(X_{0},\dots,X_{M})\right\|^{2},$
$h_{0}(X_{0})=\langle X_{0},\Omega_{k}\rangle,$
$h_{i}(X_{i})=-\left\langle X_{i},T_{2}(V^{\top}\otimes V){\rm vec}(Q)+T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]T_{1}^{\top}\Omega_{k}\right\rangle,\leavevmode\nobreak\ i=1,\dots,M,$
then optimization problem (32) is equivalent to
$\min\limits_{X_{0}\dots,X_{M}}$ $\widehat{\Theta}(X_{0},\dots,X_{M})=g(X_{0},\dots,X_{M})+\sum_{i=0}^{M}h_{i}(X_{i})$
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak$ $X_{0}\times X_{1}\dots\times X_{M}\in\mathcal{X}=\widetilde{\Gamma}_{+}^{p}\times\widetilde{\Gamma}_{+}^{n}\times\dots\times\widetilde{\Gamma}_{+}^{n}.$
Hence, ignoring the forward sGS sweep (i.e., regard algorithm 1 as BSUM algorithm with G-S (Gauss-Seidel) rule), let
$u_{i}(X_{i},X^{k})=g(X_{1}^{k},\dots,X_{i},\dots,X_{M}^{k})+\frac{1}{2}\|X_{i}-X_{i}^{k}\|_{\mathcal{S}_{i}}^{2},\leavevmode\nobreak\ i=0,\dots,M$
with $X^{k}=(X_{0}^{k},\dots,X_{M}^{k})$. It is evident that function $u_{i}$ serves as a strictly auxiliary function, leading us to derive the subsequent theorem.
Theorem 2.
Let $\{X^{k}\}$ be the sequence generated by the BSUM algorithm with G-S rule. Then it holds
$\widehat{\Theta}(X^{k})-\widehat{\Theta}^{*}\leq\frac{c_{1}}{\sigma_{1}}\frac{1}{k},\leavevmode\nobreak\ \forall k\geq 1,$
where $\gamma=\frac{1}{2}\min_{k}\gamma_{k}$,
$R=\max\limits_{X\in\mathcal{X}}\max\limits_{X^{*}\in\mathcal{X}^{*}}\left\{\|X-X^{*}\|\colon\widehat{\Theta}(X)\leq\widehat{\Theta}(X^{0})\right\},$
$c_{1}=\max\{4\sigma_{1}-2,\widehat{\Theta}(X^{0})-\widehat{\Theta}^{*},2\},$
$\sigma_{1}=\frac{\gamma}{(M+1)G_{\rm max}^{2}R^{2}}.$
Proof.
Based on the above discussion, this theorem is a direct corollary of Theorem 1 of [24]. ∎
In algorithm 1 below, the symmetric G-S rule is employed instead of the G-S rule in order to enhance the algorithm’s performance using empirical means. Review the optimality condition of (32):
00 0 $\in\partial_{X_{0}}\Theta(X_{0},\dots,X_{M})=-H_{0}+L_{0}X_{0}+\partial\delta_{\widetilde{\Gamma}_{+}^{p}}(X_{0}),$
00 0 $\in\partial_{X_{1}}\Theta(X_{0},X_{1},\dots,X_{M})=-H_{1}+L_{1}X_{1}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{1}),$
$\vdots$
00 0 $\in\partial_{X_{M}}\Theta(X_{0},X_{1},\dots,X_{M})=-H_{M}+L_{M}X_{M}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{M}),$
which are equivalent to
$X_{0}-\Pi_{\widetilde{\Gamma}_{+}^{p}}\left(X_{0}-L_{0}X_{0}+H_{0}\right)=0,$
$X_{1}-\Pi_{\widetilde{\Gamma}_{+}^{n}}\left(X_{1}-L_{1}X_{1}+H_{1}\right)=0,$
[…]
$X_{M}-\Pi_{\widetilde{\Gamma}_{+}^{n}}\left(X_{M}-L_{M}X_{M}+H_{M}\right)=0.$
Then, the relative residual error can be defined as
${\rm err}_{X_{0}}^{k}$ $=\frac{\left\|X_{0}^{k}-\Pi_{\widetilde{\Gamma}_{+}^{p}}\left(X_{0}^{k}-L_{0}X_{0}^{k}+H_{0}\right)\right\|}{1+\|X_{0}^{k}\|+\|L_{0}X_{0}^{k}-H_{0}\|},$ (36)
${\rm err}_{X_{i}}^{k}$ $=\frac{\left\|X_{i}^{k}-\Pi_{\widetilde{\Gamma}_{+}^{n}}\left(X_{i}^{k}-L_{i}X_{i}^{k}+H_{i}\right)\right\|}{1+\|X_{i}^{k}\|+\|L_{i}X_{i}^{k}-H_{i}\|},\leavevmode\nobreak\ i=1,\dots,M.$
Let
${\rm err}^{k}=\max\{{\rm err}_{X_{0}}^{k},{\rm err}_{X_{1}}^{k},\dots,{\rm err}_{X_{M}}^{k}\},$ (37)
and the inner layer optimization process will terminate if ${\rm err}^{k}<\epsilon$. By (30),
${\rm svec}(W)=M^{-1}(q_{k}-\mathcal{L}_{k}(X_{0}^{k},X_{1}^{k},\dots,X_{M}^{k}))$ (38)
is the optimal solution to (26).
  Algorithm 1 Coordinate Descent with Proximal Term (Inner Layer Solver)
1:Given : Initial point $(X_{0}^{0},X_{1}^{0},\dots,X_{M}^{0})$, stopping criterion parameter $\epsilon$
2:Result : ${\rm vec}(W)$
3:for $k=0,1,2,...$ do
4:     Update $\bar{X}_{M}^{k+1},\dots,\bar{X}_{1}^{k+1},X_{0}^{k+1}$ by (33), (34)
5:     Update $X_{1}^{k+1},\dots,X_{M}^{k+1}$ by (35)
6:     Compute ${\rm err}_{X_{0}}^{k+1},{\rm err}_{X_{1}}^{k+1},\dots,{\rm err}_{X_{M}}^{k+1}$ by (36)
7:     Determine ${\rm err}^{k+1}$ by (37)
8:     if ${\rm err}^{k+1}<\epsilon$ then
9:         Compute ${\rm svec}(W)$ by (38)
10:         return ${\rm vec}(W)=D_{1}{\rm svec}(W)$
11:     end if
12:end for
5 Discussion of Convergence Rate and Acceleration
  Algorithm 2 ALM based Decentralized LQ (Outer Layer Solver)
1:Given : Initial point $\widetilde{W}_{0}\in\Gamma^{p},v_{0}\in\Gamma^{p},\widetilde{P}_{0}\in\mathbb{R}^{mn},w_{0}\in\mathbb{R}^{mn},\lambda_{0}\in\mathbb{R}^{mn},\theta_{0}=1,\beta_{0}>0,\gamma_{0}>0$, smoothness parameter $L_{f}$
2:Result : $\widetilde{W}_{k}$, $\widetilde{P}_{k}$
3:for $k=0,1,2,...$ do
4:     Compute $\alpha_{k}$ by (40)
5:     Compute $\eta_{g,k},\tilde{y}_{k},\tilde{\eta}_{f,k},\tilde{v}_{k}$ by (14), (15), (16), (17)
6:     Determine $u_{k}=(1+\alpha_{k})^{-1}(\widetilde{W}_{k}+\alpha_{k}v_{k}),\leavevmode\nobreak\ d_{k}=\nabla f_{1}(u_{k})+\mathcal{A}^{\top}\lambda_{k}$
7:     Update $v_{k+1}\leftarrow$Inner Layer Solver
8:     Update $\widetilde{W}_{k+1}=(1+\alpha_{k})^{-1}(\widetilde{W}_{k}+\alpha_{k}v_{k+1})$
9:     Update $\bar{\lambda}_{k+1}=\lambda_{k}+\alpha_{k}/\theta_{k}(\mathcal{A}v_{k+1}+\mathcal{B}w_{k})$
10:     Update $\widetilde{P}_{k+1}$ by (24)
11:     Update $w_{k+1}=\widetilde{P}_{k+1}+\alpha_{k}^{-1}(\widetilde{P}_{k+1}-\widetilde{P}_{k})$
12:     Update $\lambda_{k+1}=\lambda_{k}+\alpha_{k}/\theta_{k}(\mathcal{A}v_{k+1}+\mathcal{B}w_{k+1})$
13:     Update $\theta_{k+1},\gamma_{k+1},\beta_{k+1}$ by (18)
14:     if Stopping Criterion==absent$=$ =True then
15:         return $\widetilde{W}_{k+1},\widetilde{P}_{k+1}$
16:     end if
17:end for
Algorithm 2 is the outer layer solver of optimization problem (11). Let the Lagrange function associated to problem (11) be
$\mathcal{L}(\widetilde{W},\widetilde{P},\lambda)=f(\widetilde{W})+h(\widetilde{P})+\left\langle\lambda,\mathcal{A}\widetilde{W}+\mathcal{B}\widetilde{P}\right\rangle,$ (39)
and let the discrete Lyapunov function be
$\mathcal{E}_{k}=\mathcal{L}(\widetilde{W}_{k},\widetilde{P}_{k},\lambda^{*})-\mathcal{L}(\widetilde{W}^{*},\widetilde{P}^{*},\lambda_{k})+\frac{\gamma_{k}}{2}\|v_{k}-\widetilde{W}^{*}\|^{2}+\frac{\beta_{k}}{2}\|w_{k}-\widetilde{P}^{*}\|^{2}+\frac{\theta_{k}}{2}\|\lambda_{k}-\lambda^{*}\|^{2},$
where $(\widetilde{W}^{*},\widetilde{P}^{*},\lambda^{*})$ is a saddle point of $\mathcal{L}$, i.e.,
$\mathcal{L}(\widetilde{W}^{*},\widetilde{P}^{*},\lambda)\leq\mathcal{L}(\widetilde{W}^{*},\widetilde{P}^{*},\lambda^{*})\leq\mathcal{L}(\widetilde{W},\widetilde{P},\lambda^{*}),\leavevmode\nobreak\ \forall(\widetilde{W},\widetilde{P},\lambda).$
The convergence analysis of Algorithm 2 can be derived, as stated in the following theorem.
Theorem 3.
Under the initial condition
$\gamma_{0}>0,\leavevmode\nobreak\ \leavevmode\nobreak\ \text{\rm and }\beta_{0}>0$
and the condition
$\|\mathcal{B}\|^{2}\alpha_{k}^{2}=\beta_{k}\theta_{k},$ (40)
it holds that $\{\widetilde{W}_{k},\widetilde{P}_{k}\}\subset\mathbb{R}^{p^{2}}\times\mathbb{R}^{mn}$ and
$\left\|\mathcal{A}\widetilde{W}_{k}+\mathcal{B}\widetilde{P}_{k}\right\|\leq\theta_{k}\mathcal{R}_{0},\quad\left|F(\widetilde{W}_{k},\widetilde{P}_{k})-F^{*}\right|\leq\theta_{k}\left(\mathcal{E}_{0}+\left\|\lambda^{*}\right\|\mathcal{R}_{0}\right),$
with $F(\widetilde{W},\widetilde{P})=f(\widetilde{W})+h(\widetilde{P})$. Above, $\mathcal{R}_{0}=\sqrt{2\mathcal{E}_{0}}+\|\lambda_{0}-\lambda^{*}\|+\|\mathcal{A}\widetilde{W}_{0}+\mathcal{B}\widetilde{P}_{0}\|$ and $\theta_{k}$ satisfies
$\theta_{k}\lesssim\frac{\|\mathcal{B}\|}{\sqrt{\beta_{0}}k},$
provided that $\beta_{0}\leq\left\|\mathcal{B}\right\|^{2}$.
Proof.
Noticing that $f_{1}(\widetilde{W})$ is a linear function, $\nabla f_{1}$ is Lipschitz continuous with Lipschitz parameter $L_{f_{1}}=0$. Then, this theorem is a direct corollary of Theorem 4.1 of [31]. ∎
The subsequent proposition elucidates the asymptotic behavior of optimization problem (8) when the parameter $\gamma$ approaches zero. Therefore, it is possible to achieve a near-optimal centralized controller by selecting a sufficiently small value for $\gamma$ and utilizing Algorithm 2.
Proposition 4.
Assume that $C^{\top}C\succ 0$, and denote
$F^{\gamma_{n}}(\widetilde{W})$ $=f(\widetilde{W})+\gamma_{n}g((V_{2}\otimes V_{1})\widetilde{W})+\delta_{\widehat{\mathcal{A}}\widetilde{W}=0}(\widetilde{W}),$
$F(\widetilde{W})$ $=f(\widetilde{W})+\delta_{\widehat{\mathcal{A}}\widetilde{W}=0}(\widetilde{W}),$
where $\widehat{\mathcal{A}}$ is defined in (46). Then it holds
$\inf F^{\gamma_{n}}\to\inf F$
Proof.
Obviously, $F^{\gamma_{n}}(\widetilde{W})\stackrel{{\scriptstyle\rm e}}{{\longrightarrow}}F(\widetilde{W})$ (the epigraph convergence given below in Section 6). Based on the fact
$\langle{\rm vec}(R),\widetilde{W}\rangle=\langle R,W\rangle\geq\lambda_{{\rm min}}(R){\rm tr}(W)\geq\lambda_{{\rm min}}(R)\|W\|_{2},$
$F(\widetilde{W})$ is level-bounded. Additionally, all $F^{\gamma_{n}}$ are closed convex function, and thus all the sets ${\rm lev}_{\leq\alpha}F^{\gamma_{n}}$ is connected. Then, $\{F^{\gamma_{n}}\}_{n\in\mathbb{N}}$ is eventually level-bounded by 7.32 (c) of [41]. Hence
$\inf F^{\gamma_{n}}\to\inf F$
by Theorem 7.33 of [41]. ∎
Above, our choice of sparsity promoting function is the weighted-$l_{1}$ norm, i.e., $g(P)=\sum_{i,j}w_{ij}|P_{ij}|$, which is a convex but not a strongly convex function. Hence, regretfully, $\mu_{g}=0$, and algorithm (13) exhibits $\mathcal{O}\left(\frac{1}{k}\right)$ convergence rate. Alternatively, designing a strongly convex sparsity promoting function is one approach to resolving the issue described above. For instance, let $g_{2}(x):\mathbb{R}\to\mathbb{R}$ be
$g_{2}(x)=\left\{\begin{aligned} &\frac{1}{2}a_{1}x^{2}+b_{1}x,\leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ x\leq 0,\\ &\frac{1}{2}a_{2}x^{2}+b_{2}x,\leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ x>0,\end{aligned}\right.$ (41)
and $g_{\mathcal{Q}}(P)=\sum_{ij}w_{ij}g_{2}(p_{ij})$ called a weighted piecewise quadratic function. The proof of the following lemma is straightforward.
Lemma 3.
$g_{\mathcal{Q}}(P)$ is sparsity promoting if and only if
$a_{1}\geq 0,\leavevmode\nobreak\ a_{2}\geq 0,\leavevmode\nobreak\ b_{1}\leq 0\leq b_{2},\leavevmode\nobreak\ {\rm and}\leavevmode\nobreak\ b_{2}-b_{1}>0.$
Obviously, by choosing $a_{1},a_{2}>0,b_{1}<0<b_{2}$, $g_{\mathcal{Q}}(P)$ becomes a strongly convex sparsity promoting function. Hence, there comes a surprising result that designing an algorithm with an accelerated convergence rate is equivalent to designing a strongly convex sparsity promoting function. If we consider the piecewise quadratic function $g_{\mathcal{Q}}(P)$, the primal optimization problem becomes
$\min_{\widetilde{W},\widetilde{P}}\leavevmode\nobreak\ \leavevmode\nobreak$ $f(\widetilde{W})+\gamma g_{\mathcal{Q}}(\widetilde{P})$
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $\mathcal{A}\widetilde{W}+\mathcal{B}\widetilde{P}=0.$
Similarly, by (13), we have
$\widetilde{P}_{k+1}$ $=\mathbf{prox}_{\tau_{k}\gamma g_{\mathcal{Q}}}^{\mathbb{R}^{mn\times 1}}(\widetilde{P}_{k}-\tau_{k}\mathcal{B}^{\top}\bar{\lambda}_{k+1})$
$=\left(\underset{P_{ij}\in\mathbb{R}}{\operatorname*{argmin}}\left\{\sum_{i,j}\left(\gamma w_{ij}g_{2}(P_{ij})+\frac{\rho_{k}}{2}\left(P_{ij}-\Omega^{k}_{ij}\right)^{2}\right)\right\}\right)_{mn\times 1},$
where $\Omega^{k}$ is given by (23). Focus on the following subproblem
$\min_{P_{ij}\in\mathbb{R}}\gamma w_{ij}g_{2}(P_{ij})+\frac{\rho_{k}}{2}\left(x-\Omega^{k}_{ij}\right)^{2}$ (42)
with
$g_{2}(P_{ij})=\left\{\begin{aligned} &\frac{1}{2}a_{1}P_{ij}^{2}+b_{1}P_{ij},\leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ P_{ij}\leq 0\\ &\frac{1}{2}a_{2}P_{ij}^{2}+b_{2}P_{ij},\leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ P_{ij}>0\end{aligned}\right.$
given by (41), and let $P_{ij}^{*}$ be the global minimizer of optimization problem (42).
Theorem 5.
Provided that $a_{1}>0,a_{2}>0,b_{1}<0<b_{2}$, then the existence and uniqueness of $P_{ij}^{*}$ holds. Moreover,
$P_{ij}^{*}=\left\{\begin{aligned} &0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ \Omega_{ij}^{k}<\frac{\gamma w_{ij}b_{2}}{\rho_{k}},\Omega_{ij}^{k}\geq 0,\\ &0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ \Omega_{ij}^{k}>\frac{\gamma w_{ij}b_{1}}{\rho_{k}},\Omega_{ij}^{k}<0,\\ &\frac{\rho_{k}\Omega_{ij}^{k}-\gamma w_{ij}b_{2}}{\gamma w_{ij}a_{2}+\rho_{k}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ \Omega_{ij}^{k}\geq\frac{\gamma w_{ij}b_{2}}{\rho_{k}},\Omega_{ij}^{k}\geq 0,\\ &\frac{\rho_{k}\Omega_{ij}^{k}-\gamma w_{ij}b_{1}}{\gamma w_{ij}a_{1}+\rho_{k}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ \Omega_{ij}^{k}\leq\frac{\gamma w_{ij}b_{1}}{\rho_{k}},\Omega_{ij}^{k}<0.\end{aligned}\right.$ (43)
Proof.
Noticing the fact that for all $x,y\in\mathbb{R}$ and $\alpha\in[0,1]$
$|\alpha x+(1-\alpha)y|^{2}=\alpha|x|^{2}+(1-\alpha)|y|^{2}-\alpha(1-\alpha)|x-y|^{2},$
it is obvious that optimization problem (42) is strongly convex. Thus, $P_{ij}^{*}$ exists and is unique. Denote $\varphi(P_{ij})=\gamma w_{ij}g_{2}(P_{ij})+\frac{\rho_{k}}{2}\left(P_{ij}-\Omega^{k}_{ij}\right)^{2}$. When $\Omega_{ij}^{k}\geq 0$, we can obtain
$\varphi(P_{ij})=\left\{\begin{aligned} &\left(\frac{\gamma w_{ij}a_{1}+\rho_{k}}{2}\right)P_{ij}^{2}+\left(\gamma w_{ij}b_{1}-\rho_{k}\Omega_{ij}^{k}\right)P_{ij}+\frac{\rho_{k}}{2}{\Omega_{ij}^{k}}^{2},\leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ P_{ij}\leq 0,\\ &\left(\frac{\gamma w_{ij}a_{2}+\rho_{k}}{2}\right)P_{ij}^{2}+\left(\gamma w_{ij}b_{2}-\rho_{k}\Omega_{ij}^{k}\right)P_{ij}+\frac{\rho_{k}}{2}{\Omega_{ij}^{k}}^{2},\leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ P_{ij}>0.\end{aligned}\right.$
Hence, $\varphi(P_{ij})$ monotonously decreases on $(-\infty,0]$. Moreover, if $\Omega_{ij}^{k}<\frac{\gamma w_{ij}b_{2}}{\rho_{k}}$, $\varphi(P_{ij})$ will monotonously increase on $[0,+\infty)$, and if $\Omega_{ij}^{k}\geq\frac{\gamma w_{ij}b_{2}}{\rho_{k}}$, $\varphi(P_{ij})$ will monotonously decrease on $\left[0,\frac{\rho_{k}\Omega_{ij}^{k}-\gamma w_{ij}b_{2}}{\gamma w_{ij}a_{2}+\rho_{k}}\right]$, while increase on $\left[\frac{\rho_{k}\Omega_{ij}^{k}-\gamma w_{ij}b_{2}}{\gamma w_{ij}a_{2}+\rho_{k}},+\infty\right)$. Consequently, when $\Omega_{ij}^{k}\geq 0$, we have
$P_{ij}^{*}=\left\{\begin{aligned} &0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ \Omega_{ij}^{k}<\frac{\gamma w_{ij}b_{2}}{\rho_{k}},\\ &\frac{\rho_{k}\Omega_{ij}^{k}-\gamma w_{ij}b_{2}}{\gamma w_{ij}a_{2}+\rho_{k}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ \Omega_{ij}^{k}\geq\frac{\gamma w_{ij}b_{2}}{\rho_{k}}.\end{aligned}\right.$
When $\Omega_{ij}^{k}<0$, by the same token, we have
$P_{ij}^{*}=\left\{\begin{aligned} &0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ \Omega_{ij}^{k}>\frac{\gamma w_{ij}b_{1}}{\rho_{k}},\\ &\frac{\rho_{k}\Omega_{ij}^{k}-\gamma w_{ij}b_{1}}{\gamma w_{ij}a_{1}+\rho_{k}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ \Omega_{ij}^{k}\leq\frac{\gamma w_{ij}b_{1}}{\rho_{k}}.\end{aligned}\right.$
Here, we finish the proof. ∎
Remark 9.
Under abovementioned discussion, $g_{2}(x)$ is strongly convex with parameter $\mu_{g_{2}}=\min\{a_{1},a_{2}\}$. It is not hard to show that $g_{\mathcal{Q}}(P)=\sum_{i,j}w_{ij}g_{2}(P_{ij})$ is strongly convex with parameter $\mu_{g_{\mathcal{Q}}}=\min_{i,j}w_{ij}\mu_{g_{2}}$, if for all $i,j$, $w_{ij}\neq 0$.
  Algorithm 3 Accelerated ALM based Decentralized LQ (Outer Layer Solver)
1:Given : Initial point $\widetilde{W}_{0}\in\Gamma^{p},v_{0}\in\Gamma^{p},y_{0}\in\mathbb{R}^{mn},w_{0}\in\mathbb{R}^{mn},\lambda_{k}\in\mathbb{R}^{mn},\theta_{0}=1,\beta_{0}>0,\gamma_{0}>0$, smoothness parameter $L_{f}$
2:Result : $\widetilde{W}_{k}$, $\widetilde{P}_{k}$
3:for $k=0,1,2,...$ do
4:     Compute $\alpha_{k}$ by (44)
5:     Compute $\eta_{g,k},\tilde{y}_{k},\tilde{\eta}_{f,k},\tilde{v}_{k}$ by (14), (15), (16), (17)
6:     Determine $u_{k}=(1+\alpha_{k})^{-1}(\widetilde{W}_{k}+\alpha_{k}v_{k}),\leavevmode\nobreak\ d_{k}=\nabla f_{1}(u_{k})+\mathcal{A}^{\top}\lambda_{k}$
7:     Update $v_{k+1}\leftarrow$Inner Layer Solver
8:     Update $\widetilde{W}_{k+1}=(1+\alpha_{k})^{-1}(\widetilde{W}_{k}+\alpha_{k}v_{k+1})$
9:     Update $\bar{\lambda}_{k+1}=\lambda_{k}+\alpha_{k}/\theta_{k}(\mathcal{A}v_{k+1}+\mathcal{B}w_{k})$
10:     Update $\widetilde{P}_{k+1}$ by (43)
11:     Update $\lambda_{k+1}=\lambda_{k}+\alpha_{k}/\theta_{k}(\mathcal{A}v_{k+1}+\mathcal{B}w_{k+1})$
12:     Update $\theta_{k+1},\gamma_{k+1},\beta_{k+1}$ by (18)
13:     if Stopping Criterion==absent$=$ =True then
14:         return $\widetilde{W}_{k+1},\widetilde{P}_{k+1}$
15:     end if
16:end for
Interestingly, the only difference between Algorithm 2 and Algorithm 3 is the update of $\widetilde{P}_{k+1}$. However, the former holds $\mathcal{O}\left(\frac{1}{k}\right)$ convergence rate, while the latter holds $\mathcal{O}\left(\frac{1}{k^{2}}\right)$ convergence rate. Note that Nesterov’s accelerated gradient descent is an extension of vanilla gradient descent by introducing an additional sequence and modifying the fundamental framework; Algorithm 2 and Algorithm 3 have a completely consistent framework. Specifically, the following theorem is proposed.
Theorem 6.
Provided that for all $i,j$, $w_{ij}\neq 0$, and $a_{1}>0,a_{2}>0,b_{1}<0<b_{2}$, under the initial condition
$\gamma_{0}>0,\leavevmode\nobreak\ \leavevmode\nobreak\ \text{\rm and }\beta_{0}=\mu_{g_{\mathcal{Q}}}>0$
and the condition
$\|\mathcal{B}\|^{2}\alpha_{k}^{2}=\beta_{k}\theta_{k},$ (44)
it holds $\{\widetilde{W}_{k},\widetilde{P}_{k}\}\subset\mathbb{R}^{p^{2}}\times\mathbb{R}^{mn}$ and
$\left\|\mathcal{A}\widetilde{W}_{k}+\mathcal{B}\widetilde{P}_{k}\right\|\leq\theta_{k}\mathcal{R}_{0},\quad\left|F(\widetilde{W}_{k},\widetilde{P}_{k})-F^{*}\right|\leq\theta_{k}\left(\mathcal{E}_{0}+\left\|\lambda^{*}\right\|\mathcal{R}_{0}\right),$
where $F(\widetilde{W},\widetilde{P})=f(\widetilde{W})+h(\widetilde{P})$. Above, $\mathcal{R}_{0}=\sqrt{2\mathcal{E}_{0}}+\|\lambda_{0}-\lambda^{*}\|+\|\mathcal{A}\widetilde{W}_{0}+\mathcal{B}\widetilde{P}_{0}\|$ and $\theta_{k}$ satisfies
$\theta_{k}\lesssim\min\left\{\frac{\|\mathcal{B}\|}{\sqrt{\beta_{0}}k},\frac{\|\mathcal{B}\|^{2}}{\mu_{g_{\mathcal{Q}}}k^{2}}\right\},$
where $\mu_{g_{\mathcal{Q}}}=\min_{i,j}w_{ij}\mu_{g_{2}}$, provided that $\beta_{0}\leq\left\|\mathcal{B}\right\|^{2}$.
Proof.
Noticing that $f_{1}(\widetilde{W})$ is a linear function, $\nabla f_{1}$ is Lipschitz continuous with Lipschitz parameter $L_{f_{1}}=0$. In addition, $g_{\mathcal{Q}}$ is $\mu_{g_{\mathcal{Q}}}$-strongly convex. Then, this theorem is a direct corollary of Theorem 4.1 of [31]. ∎
6 Directly Optimizing $\ell_{0}$-penalty
In the above sections, our primary objective is to minimize the $\ell_{0}$ norm (See Definition 9) of $W_{2}$; and indeed, the $\ell_{1}$-penalty and the piecewise quadratic penalty function can be considered as a convex relaxation of the $\ell_{0}$ norm (see Figure 1). In this section, the $\ell_{0}$ sparsity promoting function is considered directly, which makes the proposed optimization problem nonconvex and discontinuous. We will show that a series of coordinatewise convex (but nonconvex) problems can serve as an apporximation for the nonconvex problem at hand. Then, the BSUM framework is employed, allowing for the effective handling of the coordinatewise convex problem. Finally, some theorems are proposed to illustrate the variational properties of $\ell_{0}$-penalty induced problem. We first introduce some standard definitions below.
Definition 3.
Let $f\colon\mathcal{D}\to\mathbb{R}$ be a function where $\mathcal{D}\subseteq\mathbb{R}^{m}$ is a convex set. The directional derivative of $f$ at point $x$ in direction $d$ is defined by
$f^{\prime}(x;d)\doteq\underset{\lambda\downarrow 0}{\rm liminf}\frac{f(x+\lambda d)-f(x)}{\lambda}.$
Definition 4.
Let $f\colon\mathcal{D}\to\mathbb{R}$ be a function where $\mathcal{D}\subseteq\mathbb{R}^{m}$ is a convex set. The point $x$ is a stationary point of $f(\cdot)$ if $f^{\prime}(x;d)\geq 0$ for all $d$ such that $x+d\in\mathcal{D}$.
Definition 5.
$z\in{\rm dom}f\subseteq\mathbb{R}^{m}$ is the coordinatewise minimizer of $f$ with respective to the coordinates in $\mathbb{R}^{m_{1}},\dots,\mathbb{R}^{m_{n}}$, $m_{1}+\dots+m_{n}=m$ if
$f(z+d_{k}^{0})\geq f(z)\leavevmode\nobreak\ \forall d_{k}\in\mathbb{R}^{m_{k}}\leavevmode\nobreak\ {\rm with}\leavevmode\nobreak\ z+d_{k}^{0}\in{\rm dom}f\leavevmode\nobreak\ \forall k=1,\dots,n,$
where $d_{k}^{0}=(0,\dots,d_{k},\dots,0)$.
Definition 6.
The function $f\colon\mathbb{R}^{m}\to\mathbb{R}$ is regular at the point $z\in{\rm dom}f$ with respective to the coordinates $m_{1},m_{2},\dots,m_{n}$,$m_{1}+\dots+m_{n}=m$ if $f^{\prime}(z;d)\geq 0$ for all $d=(d_{1},d_{2},\dots,d_{n})$ with $f^{\prime}(z;d_{k}^{0})\geq 0$, where $d_{k}^{0}\doteq(0,\dots,d_{k},\dots,0)$ and $d_{k}\in\mathbb{R}^{m_{k}}$ for all $k$.
Definition 7.
The function is lower semi-continuous and is the greatest of all the lower semi-continuous functions $g$ such that $g\leq f$. It is called the lower closure of $f$, denoted by ${\rm cl}f$.
Definition 8.
For any sequence $\{f^{v}\}_{v\in\mathbb{N}}$ of functions on $\mathbb{R}^{n}$, the lower epi-limit ${\rm e\text{-}liminf}_{v}f^{v}$ is the function having as its epigraph the outer limit of the sequence of sets ${\rm epi}f^{v}$:
${\rm epi}({\rm e\text{-}liminf}_{v}f^{v})\doteq{\rm limsup}_{v}({\rm epi}f^{v}).$
The upper epi-limit ${\rm e\text{-}limsup}_{v}f^{v}$ is the function having as its epigraph the inner limit of the sets ${\rm epi}f^{v}$:
${\rm epi}({\rm e\text{-}limsup}_{v}f^{v})\doteq{\rm liminf}_{v}({\rm epi}f^{v}).$
When these two functions coincide, the epi-limits function ${\rm e\text{-}lim}f^{v}$ is said to exist:
${\rm e\text{-}lim}f^{v}\doteq{\rm e\text{-}liminf}_{v}f^{v}={\rm e\text{-}limsup}_{v}f^{v}.$
The subsequent analysis in this section is predicated upon the following assumption.
Assumption 4.
$C^{\top}C\succ 0$.
Definition 9.
For $x=(x_{1},\dots,,x_{n})^{\top}\in\mathbb{R}^{n}$, consider the $\ell_{0}$-norm function $f:\mathbb{R}^{n}\to\mathbb{R}$ given by
$f(x)=\|x\|_{0}\equiv\#\{i\colon x_{i}\neq 0\}.$
That is, $\|x\|_{0}$ is the number of nonzero elements in $x$.
Remark 10.
In reality, $\ell_{0}$-norm is not a norm, since it does not satisfy the homogeneity property. Note that $\ell_{0}$-norm is a discontinuous and nonconvex function, but can be viewed as a sum of closed functions
$\|x\|_{0}=\sum_{i=1}^{n}I(x_{i}),$
where $I:\mathbb{R}\to\{0,1\}$ is given by
$I(y)=\left\{\begin{aligned} &0,\leavevmode\nobreak\ \leavevmode\nobreak\ y=0,\\ &1,\leavevmode\nobreak\ \leavevmode\nobreak\ \text{otherwise}.\end{aligned}\right.$
Hence, $\ell_{0}$-norm is closed. Whatever, significant challenges arise when $\ell_{0}$-penalty is introduced into the optimization problems.
By the same token with (10), $\ell_{0}$-penalty induced optimization problem can be equivalently converted into the following form
$\min_{\widetilde{W}}\leavevmode\nobreak\ \leavevmode\nobreak$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})+\gamma\|(V_{2}\otimes V_{1})\widetilde{W}\|_{0}$ (45)
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $\left(V_{i2}^{\top}\otimes V_{i1}\right)\widetilde{W}=0,\leavevmode\nobreak\ \leavevmode\nobreak\ i=1,\dots,M.$
Denote
$\widehat{\mathcal{A}}=\begin{pmatrix}V_{12}^{\top}\otimes V_{11}\\ \vdots\\ V_{M2}^{\top}\otimes V_{M1}\\ \end{pmatrix},$ (46)
and let $\Omega=\{\widetilde{W}\colon\widehat{\mathcal{A}}\widetilde{W}=0\}$. Then, the optimization problem can be expressed in the unconstrained form, i.e.,
$\min_{\widetilde{W}}\leavevmode\nobreak\ \langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})+\delta_{\Omega}(\widetilde{W})+\gamma\|(V_{2}\otimes V_{1})\widetilde{W}\|_{0}.$ (P)
The discontinuous nature of optimization problem (P) makes it extremely challenging to design a suitable algorithm. The work [34] introduces a class of continuous folded concave penalty $f_{\sigma}:\mathbb{R}\to\mathbb{R}$ with $f_{\sigma}(|y|)$ converging to $I(y)$ as $\sigma\downarrow 0$. Hence, $f_{\sigma}(|y|)$ can be viewed as a class of continuous approximation of $\ell_{0}$-norm. Specifically, Property $\mathcal{D}$ is introduced to define the class of folded concave penalty $f_{\sigma}(|y|)$.
Definition 10 ([34]).
Let $f\colon\mathbb{R}\to\mathbb{R}$ and define $f_{\sigma}(y)\doteq f(y/\sigma)$ for any $\sigma>0$. The function $f$ is said to possess Property $\mathcal{D}$, if
  1. 1.
    $f$ is real analytic on $(y_{0},\infty)$ for some $y_{0}<0$,
  2. 2.
    $\forall y\geq 0$, $f^{\prime\prime}(y)\geq-\mu_{0}$, where $\mu_{0}>0$ is some constant,
  3. 3.
    $f$ is concave on $\mathbb{R}$,
  4. 4.
    $f(y)=0\Leftrightarrow y=0$,
  5. 5.
    $\lim_{y\to\infty}f(y)=1$.
It is obvious that if $f$ possesses Property $\mathcal{D}$, then
$\lim\limits_{\sigma\downarrow 0^{+}}f_{\sigma}(|y|)=I(y)=\left\{\begin{aligned} &0,\leavevmode\nobreak\ \leavevmode\nobreak\ y=0,\\ &1,\leavevmode\nobreak\ \leavevmode\nobreak\ \text{otherwise}.\end{aligned}\right.$
In fact, there are a plenty of functions that satisfy Property $\mathcal{D}$, for instance, $f(y)=1-e^{-y}$. For the record, function $f(y)=1-e^{-y}$ is utilized in this paper without special instructions, and in this case, $f_{\sigma}(y)=1-e^{\frac{y}{\sigma}}$. For $x=(x_{1},\dots,x_{n})^{\top}\in\mathbb{R}^{n}$, denote $f_{\sigma}(x)=\sum_{i=1}^{n}f_{\sigma}(x_{i})$.
Based on the fact that $\lim\limits_{\sigma\downarrow 0^{+}}f_{\sigma}(|x|)=\|x\|_{0}$, a class of optimization problems are proposed:
$\min_{\widetilde{W}}\leavevmode\nobreak\ \langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})+\delta_{\Omega}(\widetilde{W})+\gamma f_{\sigma}\left(\left|(V_{2}\otimes V_{1})\widetilde{W}\right|\right).$ ($\textbf{P}_{\sigma}$)
($\textbf{P}_{\sigma}$) is a nonconvex optimization problem since the folded concave penalty term $f_{\sigma}$ is introduced. To avoid optimizing nonconvex penalty term directly, [42] proposes a convex relaxation framework for solving $f_{\sigma}$ penalty induced optimization problem.
Lemma 4.
Let $f(x)$ satisfying Property $\mathcal{D}$, and consider the function $f_{\sigma}(x)(x\geq 0)$. It holds
$f_{\sigma}(x)=\inf\limits_{y\leq 0}\{-yx+g^{*}_{\sigma}(y)\},$
where $g_{\sigma}(x)=-f_{\sigma}(x)$ and $g^{*}_{\sigma}$ is the conjugate function of $g_{\sigma}$.
Proof.
By Property $\mathcal{D}$, $g_{\sigma}(x)$ is a proper closed and convex function. Hence, by Theorem 4.3 of [1], $g^{*}_{\sigma}(y)$ $(y<0)$ is a closed convex function. By Theorem 4.8 of [1], we have $g_{\sigma}^{**}=g_{\sigma}$. Hence, $g_{\sigma}=\sup\limits_{y\leq 0}\{yx-g_{\sigma}^{*}(y)\}$. Equivalently, $f_{\sigma}(x)=\inf\limits_{y\leq 0}\{-yx+g^{*}_{\sigma}(y)\}$. The proof is complete. ∎
Hence, $f_{\sigma}$ penalty can be expressed as
$f_{\sigma}(|x|)=\inf\limits_{y\geq 0}\left\{y^{\top}|x|+g_{\sigma}^{*}(-y)\right\},$
and optimization problem ($\textbf{P}_{\sigma}$) is equivalent to the following problem
$\min_{\widetilde{W},y\in\mathbb{R}^{mn}_{+}}\leavevmode\nobreak$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})+\delta_{\Omega}(\widetilde{W})$ ($\textbf{P}^{\prime}_{\sigma}$)
$+\gamma y^{\top}\left|(V_{2}\otimes V_{1})\widetilde{W}\right|+\gamma g_{\sigma}^{*}(-y).$
Denote
$r_{1}(\widetilde{W})=\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})+\delta_{\Omega}(\widetilde{W}),$
$r_{2,\sigma}(y)=\gamma g_{\sigma}^{*}(-y),$
$h(\widetilde{W},y)=\gamma y^{\top}\left|(V_{2}\otimes V_{1})\widetilde{W}\right|,$
then problem ($\textbf{P}^{\prime}_{\sigma}$) can be denoted in a compact form:
$\min_{\widetilde{W},y\in\mathbb{R}^{mn}_{+}}\leavevmode\nobreak\ H_{\sigma}(\widetilde{W},y)=r_{1}(\widetilde{W})+r_{2,\sigma}(y)+h(\widetilde{W},y).$ ($\textbf{P}^{\prime\prime}_{\sigma}$)
It is important to acknowledge that $r_{1},r_{2}$ are closed convex functions, but $h$ is nonconvex and nonsmooth. It seems that, compared to optimization problem ($\textbf{P}_{\sigma}$), ($\textbf{P}^{\prime\prime}_{\sigma}$) remains a nonconvex problem but introduces auxiliary variable $y$, which even makes ($\textbf{P}^{\prime\prime}_{\sigma}$) more challenging to solve. Interestingly, it is evident that $H_{\sigma}(\widetilde{W},y)$ is closed convex with regard to $\widetilde{W}$ and $y$, respectively. This observation suggests that by employing block coordinate descent, ($\textbf{P}^{\prime\prime}_{\sigma}$) can be converted to successive convex optimization problems. Specifically, given $y_{k-1}$, update $\widetilde{W}_{k}$; then given $\widetilde{W}_{k}$, update $y_{k}$, i.e.,
$\widetilde{W}_{k}\in\underset{\widetilde{W}\in\mathbb{R}^{p^{2}}}{\operatorname*{argmin}}\leavevmode\nobreak\ r_{1}(\widetilde{W})+h(\widetilde{W},y_{k-1}),$ ($\textbf{BP}\sigma-1$)
$y_{k}\in\underset{y\in\mathbb{R}_{+}^{mn}}{\operatorname*{argmin}}\leavevmode\nobreak\ r_{2,\sigma}(y)+h(\widetilde{W}_{k},y).$ ($\textbf{BP}\sigma-2$)
It is visible that both ($\textbf{BP}\sigma-1$) and ($\textbf{BP}\sigma-2$) may be classified as convex optimization problems. Our objective is then to solve ($\textbf{BP}\sigma-1$) and ($\textbf{BP}\sigma-2$) alternatively in order to approximate the global minimizers of nonconvex problem ($\textbf{P}^{\prime\prime}_{\sigma}$). Before discussing the relation between ($\textbf{P}^{\prime\prime}_{\sigma}$) and ($\textbf{BP}\sigma-1$) ($\textbf{BP}\sigma-2$), we first propose the relation between ($\textbf{P}^{\prime\prime}_{\sigma}$) and (P). Denote
$H_{\sigma}(\widetilde{W})$ $=r_{1}(\widetilde{W})+\gamma f_{\sigma}\left(\left|(V_{2}\otimes V_{1})\widetilde{W}\right|\right),$
$H(\widetilde{W})$ $=r_{1}(\widetilde{W})+\gamma\|(V_{2}\otimes V_{1})\widetilde{W}\|_{0},$
and the following theorem proves that
[…]
Theorem 7.
Let $\sigma_{i}\downarrow 0$, then the following statements hold.
  1. 1.
    $\inf\limits_{\widetilde{W}\in\mathbb{R}^{p^{2}}}H_{\sigma_{i}}(\widetilde{W})\to\inf\limits_{\widetilde{W}\in\mathbb{R}^{p^{2}}}H(\widetilde{W})$.
  2. 2.
    For $v$ in some index set $N\in\mathcal{N}_{\infty}$, the sets ${\operatorname*{argmin}}H_{\sigma_{v}}$ are nonempty and form a bounded sequence with
    $\limsup_{v}\left({\operatorname*{argmin}}H_{\sigma_{v}}\right)\subset{\operatorname*{argmin}}H.$
  3. 3.
    For any choice of $\epsilon_{i}\downarrow 0$ and $\widetilde{W}_{i}\in\epsilon_{i}\text{-}{\operatorname*{argmin}}H_{\sigma_{i}}$, the sequence $\{\widetilde{W}_{i}\}_{i\in\mathbb{N}}$ is bounded and such that all its cluster points belong to ${\operatorname*{argmin}}H$.
Proof.
Since for $x=(x_{1},\dots,x_{n})^{\top}\in\mathbb{R}^{n}$
$f_{\sigma}(x)=\sum_{i=1}^{n}f_{\sigma}(x_{i}),$
we have $f_{\sigma_{i+1}}(x)\geq f_{\sigma_{i}}(x)$ for every $x\geq 0$. Hence, for every $\widetilde{W}\in\mathbb{R}^{p^{2}}$, it follows that $H_{\sigma_{i+1}}(\widetilde{W})\geq H_{\sigma_{i}}(\widetilde{W})$, and $\{H_{\sigma_{i}}(\widetilde{W})\}_{i\in\mathbb{N}}$ is nondecreasing. By Proposition 7.4 of [41], ${\operatorname*{e-lim}}_{i}H_{\sigma_{i}}$ exists and equals $\sup_{i}[{\operatorname*{cl}}H_{\sigma_{i}}]$. Based on the fact $\lim\limits_{\sigma\downarrow 0^{+}}f_{\sigma}(|x|)=\|x\|_{0}$, it follows that $\sup_{i}[{\operatorname*{cl}}H_{\sigma_{i}}](\widetilde{W})=H(\widetilde{W})$. Obviously, for every $\sigma_{i}$, $H_{\sigma_{i}}(\widetilde{W})\geq r_{1}(\widetilde{W})$, and $r_{1}(\widetilde{W})$ is a coercive function based on the fact
$\langle{\rm vec}(R),\widetilde{W}\rangle=\langle R,W\rangle\geq\lambda_{{\rm min}}(R){\rm tr}(W)\geq\lambda_{{\rm min}}(R)\|W\|_{2}$ (47)
for $W\in\mathbb{S}_{+}^{p}$. According to Exercise 7.32 of [1], the sequence $\{H_{\sigma_{i}}\}_{i\in\mathbb{N}}$ is eventually level-bounded. By noticing that $H_{\sigma_{i}}$ and $H$ are lower semi-continuous and proper, we finish the proof by Theorem 7.33 of [1]. ∎
Hence, theoretically, one can obtain the minimizers of optimization (P) by successively solving optimization problem ($\textbf{P}_{\sigma}$) (or, equivalently, ($\textbf{P}^{\prime\prime}_{\sigma}$)) with $\sigma\downarrow 0$. As mentioned above, we hope to solve optimization problem ($\textbf{P}^{\prime\prime}_{\sigma}$) by block coordinate descent, i.e, finding the global optimal of convex optimization problems ($\textbf{BP}\sigma-1$) and ($\textbf{BP}\sigma-2$) alternately. Regretfully, since unseparable term $h(\widetilde{W},y)$ is nonconvex and nonsmooth, direct block coordinate descent method may cause stable cyclic behavior, whose gradient is bounded away from zero in the limiting path [39]. Hence, we will then modify the optimization problems ($\textbf{BP}\sigma-1$) and ($\textbf{BP}\sigma-2$). Before that, we first introduce the auxiliary functions.
Definition 11.
For a continuous function $f:\mathcal{X}\to\mathbb{R}$, where $\mathcal{X}\subseteq\mathbb{R}^{m}$ is the Cartesian product of $n$ closed convex sets: $\mathcal{X}=\mathcal{X}_{1}\times\dots\times\mathcal{X}_{n}$ with $\mathcal{X}_{i}\in\mathbb{R}^{m_{i}}$ and $\sum_{i}m_{i}=m$. We call $\{u_{i}(x_{i},x)\}_{i=1}^{n}$ are auxiliary functions of $f$, if
  1. 1.
    $u_{i}(y_{i},y)=f(y)\leavevmode\nobreak\ \leavevmode\nobreak\ \forall y\in\mathcal{X},\forall i$,
  2. 2.
    $u_{i}(x_{i},y)\geq f(y_{1},\dots,y_{i-1},x_{i},y_{i+1},\dots,y_{n})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall x_{i}\in\mathcal{X}_{i},\forall y\in\mathcal{X},\forall i,$
  3. 3.
    $u_{i}^{\prime}(x_{i},y;d_{i})\big{|}_{x_{i}=y_{i}}=f^{\prime}(y;d)\leavevmode\nobreak\ \leavevmode\nobreak\ \forall d=(0,\dots,d_{i},\dots,0)\text{s.t. }y_{i}+d_{i}\in\mathcal{X}_{i}\forall i$,
  4. 4.
    $u_{i}(x_{i},y)$ is continuous in $(x_{i},y)\forall i$.
Remark 11.
For a continuous coordinatewise convex function $f$, auxiliary functions always exist. A trivial choice of auxiliary functions is
$u_{i}(x_{i},y)=f(y_{1},\dots,y_{i-1},x_{i},y_{i+1},\dots,y_{n})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall x_{i}\in\mathcal{X}_{i},\forall y\in\mathcal{X},\forall i.$
Noticing that a convex function is differentiable in any direction at any interior point of its domain [36], the conditions 1-4 of the above definition naturally hold.
In [40], an unified algorithm framework is introduced to solve the following optimization problem
$\min f(x),\leavevmode\nobreak\ \leavevmode\nobreak\ \text{s.t. }x\in\mathcal{X},$ (48)
where $x=(x_{1},\dots,x_{n})^{\top}\in\mathbb{R}^{m}$, and $\mathcal{X}\subseteq\mathbb{R}^{m}$ is the Cartesian product of $n$ closed convex sets: $\mathcal{X}=\mathcal{X}_{1}\times\dots\times\mathcal{X}_{n}$ with $\mathcal{X}_{i}\in\mathbb{R}^{m_{i}}$ and $\sum_{i}m_{i}=m$.
  Algorithm 4 BSUM
1:Given : Feasible initial point $x^{0}\in\mathcal{X}$
2:Result: $x^{r}$
3:Set $r=0$
4:for $k=0,1,2,...$ do
5:     Let $r=r+1,i=(r\mod n)+1$
6:     Let $\mathcal{X}^{r}=\underset{x_{i}\in\mathcal{X}_{i}}{\operatorname*{argmin}}u_{i}(x_{i},x^{r-1})$
7:     Set $x^{r}_{i}$ to be an arbitrary element in $\mathcal{X}^{r}$
8:     Set $x_{k}^{r}=x_{k}^{r-1}\leavevmode\nobreak\ \forall k\neq i$
9:     if Stopping Criterion==absent$=$ =True then
10:         return $x^{r}$
11:     end if
12:end for
The above pseudocode of BSUM (Block Successive Upper-bound Minimization) framework introduced in [40] gives methodology for solving the optimization problem (48), where $\{u_{i}\}_{i=1}^{n}$ are auxiliary functions of $f$. It should be noted that the straight implementation of the BSUM framework is not viable due to the lack of knowledge on the auxiliary functions and the presence of multiple subproblems (see line 6 in Algorithm 4) that need to be addressed.
Specifically, back to our topic, optimization problem ($\textbf{BP}\sigma-2$) at hand remains unaltered. In other word, the auxiliary function $u_{2,\sigma}(y,X_{k})$ is selected as
$u_{2,\sigma}(y,X_{k})=\langle{\rm vec}(R),\widetilde{W}_{k}\rangle+r_{2,\sigma}(y)+h(\widetilde{W}_{k},y),$ (49)
where $X_{k}=(\widetilde{W}_{k},y_{k})$. ($\textbf{BP}\sigma-2$) is equivalent to
$\min\limits_{y\in\mathbb{R}^{mn}_{+}}g^{*}_{\sigma}(-y)+y^{\top}\left|(V_{2}\otimes V_{1})\widetilde{W}_{k}\right|.$ (50)
Lemma 5.
The global minimizer $y_{k,\sigma}^{*}$ of optimization problem (50) (equivalently, ($\textbf{BP}\sigma-2$)) is unique, and can be expressed as
$y_{k,\sigma}^{*}=\nabla f_{\sigma}\left(\left|(V_{2}\otimes V_{1})\widetilde{W}_{k}\right|\right).$ (51)
Proof.
Optimality condition implies that
$0\in-\partial g^{*}_{\sigma}(-y_{k,\sigma}^{*})+\left|(V_{2}\otimes V_{1})\widetilde{W}_{k}\right|.$
Since $g^{*}_{\sigma}$ is a proper, closed and convex function, by Proposition 11.3 of [41], we have
$-y_{k,\sigma}^{*}\in\partial g_{\sigma}\left(\left|(V_{2}\otimes V_{1})\widetilde{W}_{k}\right|\right).$
Noticing that $g_{\sigma}=-f_{\sigma}$ is a smooth function on +$\mathbb{R}^{mn}_{+}$, then subgradient operator degenerates to gradient operator, i.e.,
$y_{k,\sigma}^{*}=\nabla f_{\sigma}\left(\left|(V_{2}\otimes V_{1})\widetilde{W}_{k}\right|\right).$
The proof is complete. ∎
We then discuss the selection of $u_{1}(\widetilde{W},X_{k})$. The proximal term is introduced, i.e.,
$u_{1,\sigma}(\widetilde{W},X_{k};\lambda)=\langle{\rm vec}(R),\widetilde{W}\rangle+r_{2,\sigma}(y_{k})+h(\widetilde{W},y_{k})+\frac{1}{2\lambda}\left\|\widetilde{W}-\widetilde{W}_{k}\right\|^{2},$ (52)
where $\lambda\in\mathbb{R}$ is a constant. Now, we direct our attention on solving the subsequent optimization problem rather than ($\textbf{BP}\sigma-1$)
$\min\limits_{\widetilde{W}}\leavevmode\nobreak$ $u_{1,\sigma}(\widetilde{W},X_{k};\lambda)$ (53)
s.t. $\widetilde{W}\in\mathcal{X}_{1},$
where $\mathcal{X}_{1}=\{\widetilde{W}\in\mathbb{R}^{p^{2}}\colon\widetilde{W}\in\Gamma_{+}^{p},\widetilde{\Psi}_{i}\in\Gamma_{+}^{n},\widehat{A}\widetilde{W}=0\}$ which is closed convex set. First, we claim that $\{H_{\sigma}(\widetilde{W}_{k})\}$ is monotonically nonincreasing.
Lemma 6.
Given an initial feasible point $\widetilde{W}_{0}$, let $\widetilde{W}_{k}$ be the iterative sequence generated by Algorithm 4 with $u_{1},u_{2}$ given by (52), (49). Then
$H_{\sigma}(\widetilde{W}_{k+1})\leq H_{\sigma}(\widetilde{W}_{k}).$
Proof.
Let $F(\widetilde{W},X_{k})=\langle{\rm vec}(R),\widetilde{W}\rangle+r_{2,\sigma}(y_{k})+h(\widetilde{W},y_{k})$. According to the fact that
$y_{k}=\nabla f_{\sigma}\left(\left|(V_{2}\otimes V_{1})\widetilde{W}_{k}\right|\right),$
$F(\widetilde{W},X_{k})$ can be equivalently regarded as $F(\widetilde{W},\widetilde{W}_{k})$. Obviously, for all $\widetilde{W}$, we have
$u_{1,\sigma}(\widetilde{W},X_{k};\lambda)\geq F(\widetilde{W},\widetilde{W}_{k})\geq H_{\sigma}(\widetilde{W}).$
Hence, the following inequality holds
$H_{\sigma}(\widetilde{W}_{k+1})\leq u_{1,\sigma}(\widetilde{W}_{k+1},X_{k};\lambda)\leq u_{1,\sigma}(\widetilde{W}_{k},X_{k};\lambda)=F(\widetilde{W}_{k},\widetilde{W}_{k})=H_{\sigma}(\widetilde{W}_{k}).$
The proof is complete. ∎
Obviously, $\{H_{\sigma}(\widetilde{W}_{k})\}$ is bounded below by 00. Hence, by basic mathematic analysis, $H_{\sigma}(\widetilde{W}_{k})\downarrow H_{\sigma}^{\infty}$, where $H_{\sigma}^{\infty}$ is a positive constant. An inquiry that arises is whether or not $H_{\sigma}^{\infty}=H_{\sigma}^{*}\doteq\min H_{\sigma}(\widetilde{W})$. Regretfully, the above inquiry remains open, but we have the following result.
Theorem 8.
Every cluster point $z=(z_{1},z_{2})$ of the iterates generated by the Algorithm 4 with $u_{1},u_{2}$ given by (52), (49) is a coordinatewise minimum of the optimization problem ($\textbf{P}^{\prime\prime}_{\sigma}$). In addition, if $H_{\sigma}(\cdot,\cdot)$ is regular at $z$, then $z$ is a stationary point of ($\textbf{P}^{\prime\prime}_{\sigma}$).
Proof.
For a feasible $X_{0}=(\widetilde{W}_{0},y_{0})$, by (47), the sublevel set
$\mathcal{X}^{0}\doteq\left\{X=(\widetilde{W},y)\colon H_{\sigma}(X)\leq H_{\sigma}(X_{0})\right\}$
is compact. By lemma 5, optimization problem ($\textbf{BP}\sigma-2$) has a unique solution for any point $X_{r-1}\in\mathcal{X}$. Then, the proof is complete by Theorem 2 of [40]. ∎
There exists a significant disparity between coordinatewise minimizer and stationary point (local minimizer). When the objective function is regular at coordinatewise minimizer, the coordinatewise minimum becomes a local minimum by the above theorem. In Lemma 3.1 of [45], there is a few discussion about regularity property of a Gateaux-differentiable function, but $H_{\sigma}(\widetilde{W},y)$ is generally not Gateaux-differentiable in this paper. Furthermore, if we assume that $H_{\sigma}^{\infty}=H_{\sigma}^{*}$ holds, we can obtain the following theorem.
Theorem 9.
If $H_{\sigma}^{\infty}=H_{\sigma}^{*}$, then any cluster point of $\{\widetilde{W}_{k}\}$ belongs to ${\rm argmin}H_{\sigma}(\widetilde{W})$.
Proof.
Since $f_{\sigma}(|(V_{2}\otimes V_{1})\widetilde{W}|)\geq 0$, it is obvious that, for all $\alpha\in\mathbb{R}$, ${\rm lev}_{\leq\alpha}H_{\sigma}(\widetilde{W})$ is compact. Then the set
$\left\{\widetilde{W}\colon H_{\sigma}(\widetilde{W})\leq H_{\sigma}(\widetilde{W_{0}})\right\}$
is compact. Hence, by $H_{\sigma}(\widetilde{W}_{k})\downarrow H_{\sigma}^{\infty}$, $\{\widetilde{W}_{k}\}_{k\in\mathbb{N}}$ is compact. Thus, there exists a subsequence $\{\widetilde{W}_{k_{n}}\}_{n\in\mathbb{N}}$ of $\{\widetilde{W}_{k}\}$ that converges to $\widetilde{W}_{\infty}$. Based on the fact that $H_{\sigma}(\widetilde{W})$ is lower semi-continuous, we have
$H_{\sigma}^{*}=H_{\sigma}^{\infty}=\underset{k\to\infty}{\rm liminf}H_{\sigma}(\widetilde{W}_{k})\geq H_{\sigma}(\widetilde{W}_{\infty})\geq H_{\sigma}^{*}.$
The proof is complete. ∎
Our focus is now directed towards resolving subproblem (53). Obviously, (53) can be expressed as the following constrained convex optimization problem:
$\min_{\widetilde{W}}\leavevmode\nobreak\ \leavevmode\nobreak$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\gamma y_{k}^{\top}\left|(V_{2}\otimes V_{1})\widetilde{W}\right|+\frac{1}{2\lambda}\left\|\widetilde{W}-\widetilde{W}_{k}\right\|^{2}$ (54)
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $\widetilde{W}\in\Gamma^{p}_{+},$
$\widetilde{\Psi}_{i}\in\Gamma_{+}^{n},\leavevmode\nobreak\ i=1,\dots,M,$
$\widehat{\mathcal{A}}\widetilde{W}=0.$
By introducing augment variable $\widetilde{P}=(V_{2}\otimes V_{1})\widetilde{W}$, (54) becomes
$\min_{\widetilde{W},\widetilde{P}}\leavevmode\nobreak\ \leavevmode\nobreak$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\gamma y_{k}^{\top}\left|\widetilde{P}\right|+\frac{1}{2\lambda}\left\|\widetilde{W}-\widetilde{W}_{k}\right\|^{2}$ (55)
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $\widetilde{W}\in\Gamma^{p}_{+},$
$\widetilde{\Psi}_{i}\in\Gamma_{+}^{n},\leavevmode\nobreak\ i=1,\dots,M,$
$\mathcal{A}\widetilde{W}+\mathcal{B}\widetilde{P}=0.$
In this section, we use the notation
$f_{1}(\widetilde{W};\lambda)$ $=\langle{\rm vec}(R),\widetilde{W}\rangle+\frac{1}{2\lambda}\left\|\widetilde{W}-\widetilde{W}_{k}\right\|^{2},$
$f_{2}(\widetilde{W})$ $=\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M}),$
$g(\widetilde{P})$ $=\gamma y_{k}^{\top}\left|\widetilde{P}\right|,$
$f(\widetilde{W})$ $=f_{1}(\widetilde{W})+f_{2}(\widetilde{W}),$
and optimization problem (55) is equivalent to
$\min_{\widetilde{W},\widetilde{P}}\leavevmode\nobreak\ \leavevmode\nobreak$ $f_{1}(\widetilde{W};\lambda)+f_{2}(\widetilde{W})+g(\widetilde{P})$ (56)
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $\mathcal{A}\widetilde{W}+\mathcal{B}\widetilde{P}=0.$
Interestingly, $g(\widetilde{P})$ is nothing but a weighted-$\ell_{1}$ norm with weight values given by $y_{k}$. Hence, compared with optimization problem (11), the only difference is the smooth and strongly convex term $\frac{1}{2\lambda}\|\widetilde{W}-\widetilde{W}_{k}\|^{2}$ in $f_{1}(\widetilde{W};\lambda)$. Thus, we can utilize optimization scheme (13) by slightly modifying Algorithm 2 (just change $w_{ij}$ to $y_{k}^{(ij)}$, where $y_{k}^{(ij)}$ is the $(i,j)$-th element of $y_{k}$). The following theorem gives the convergence rate of optimization scheme (13) to solve (56).
Theorem 10.
Under the initial setting
$\gamma_{0}=\frac{1}{\lambda}>0,\leavevmode\nobreak\ \leavevmode\nobreak\ \text{\rm and }\beta_{0}>0$
and the condition
$\left(L_{f}\beta_{k}\theta_{k}+\gamma_{k}\|B\|^{2}\right)\alpha_{k}^{2}=\gamma_{k}\beta_{k}\theta_{k},$
we have $\{\widetilde{W}_{k},\widetilde{P}_{k}\}\subset\mathbb{R}^{p^{2}}\times\mathbb{R}^{mn}$ and
$\left\|\mathcal{A}\widetilde{W}_{k}+\mathcal{B}\widetilde{P}_{k}\right\|\leq\theta_{k}\mathcal{R}_{0},\quad\left|F(\widetilde{W}_{k},\widetilde{P}_{k})-F^{*}\right|\leq\theta_{k}\left(\mathcal{E}_{0}+\left\|\lambda^{*}\right\|\mathcal{R}_{0}\right),$
where $F(\widetilde{W},\widetilde{P})=f(\widetilde{W})+h(\widetilde{P})$. Above, $\mathcal{R}_{0}=\sqrt{2\mathcal{E}_{0}}+\|\lambda_{0}-\lambda^{*}\|+\|\mathcal{A}\widetilde{W}_{0}+\mathcal{B}\widetilde{P}_{0}\|$ and $\theta_{k}$ satisfies
$\theta_{k}\lesssim\frac{\|\mathcal{B}\|}{\sqrt{\beta_{0}}k}+\min\left\{\frac{1}{k^{2}},\exp\left(-\frac{k}{4}\right)\right\},$
provided that $\gamma_{0}\beta_{0}\leq L_{f}\beta_{0}+\gamma_{0}\left\|\mathcal{B}\right\|^{2}$.
Proof.
By noticing that $f_{1}(\widetilde{W})$ is $\frac{1}{\lambda}$-strongly convex and $\frac{1}{\lambda}$-smooth, this theorem is a direct corollary of Theorem 4.1 of [31]. ∎
  Algorithm 5 BSUM-based $\ell_{0}$-penalty Decentralized LQ
1:Given : Feasible initial point $X_{0}=(\widetilde{W}_{0},y_{0})\in\mathcal{X}_{1}\times\mathbb{R}_{+}^{mn},\lambda>0,\gamma>0,\sigma_{0}\leq 1,\alpha\in(0,1)$
2:Result: $X_{k}=(\widetilde{W}_{k},y_{k})$
3:for $i=1,2,3\dots$ do
4:     Set $\sigma_{i}=\alpha\sigma_{i-1}$
5:     Set $k=0$
6:     for $k=0,1,2,...$ do
7:         Let $k=k+1,i=(k\mod 2)+1$
8:         if $i==1$ then
9:              $\widetilde{W}_{k}\leftarrow$ALM based Decentralized LQ
10:              Set $y_{k}=y_{k-1},X_{k}=(\widetilde{W}_{k},y_{k})$
11:         end if
12:         if $i==0$ then
13:              Update $y_{k}=\nabla f_{\sigma_{i}}\left(\left|(V_{2}\otimes V_{1})\widetilde{W}_{k}\right|\right)$
14:              Set $\widetilde{W}_{k}=\widetilde{W}_{k-1},X_{k}=(\widetilde{W}_{k},y_{k})$
15:         end if
16:         if Stopping Criterion==absent$=$ =True then
17:              Break
18:         end if
19:     end for
20:     Set $X_{0}=X_{k}$
21:end for
Remark 12.
In fact, if $u_{1,\sigma}(\widetilde{W},X_{k})$ is selected as
$u_{1,\sigma}(\widetilde{W},X_{k})=\langle{\rm vec}(R),\widetilde{W}\rangle+r_{2,\sigma}(y_{k})+h(\widetilde{W},y_{k}),$
the iterative sequence of the algorithm BSUM can also converge to the coordinatewise minimum of $H_{\sigma}(\widetilde{W},y)$. However, we introduce the proximal term $\frac{1}{2\lambda}\|\widetilde{W}-\widetilde{W}_{k}\|^{2}$ into $u_{1,\sigma}$, which appears to be gilding the lily or potentially resulting in a decelerated convergence of the BSUM algorithm. Empirically, the proximal term makes the objective of (56) be strongly convex and accelerate the solving for subproblem (56). Therefore, a trade-off exists between the convergence rate of the algorithm BSUM for optimization problem ($\textbf{P}^{\prime\prime}_{\sigma}$) and the convergence rate of (13) for subproblem (56).
Remark 13.
The algorithm exhibits a significant level of complexity, and there exists no empirical evidence to support the superiority of Algorithm 5 over Algorithm 2 (or Algorithm 3). Therefore, this section mostly focuses on theoretical discourse, particularly about the variational characteristic of the optimization problem induced by the $\ell_{0}$-penalty and the relation between $\ell_{1}$-penalty problem (8) and $\ell_{0}$-penalty problem. It is recommended that, in the pursuit of a decentralized controller, Algorithm 2 (or Algorithm 3) is employed instead of Algorithm 5.
7 Numerical Examples
In this section, a few numerical examples are represented that illustrate the theoretical results of this paper, where the noise $w$ (see (57)) is characterized by a impulse disturbance vector.
Example 1: Consider $x=[x_{1},x_{2},x_{3}]^{\top}$ and a linear system
$\dot{x}$ $-Ax+B_{2}u+B_{1}w,$ (57)
$z$ $=Cx+Du,$
$u$ $=-Kx,$
where
$A=\begin{bmatrix}0.2220&0.9186&0.7659\\ 0.8707&0.4884&0.5184\\ 0.2067&0.6117&0.2968\end{bmatrix},B_{1}=I_{3},B_{2}=\begin{bmatrix}0.9315&0.7939\\ 0.9722&0.1061\\ 0.5317&0.7750\end{bmatrix},C=\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix},D=\begin{bmatrix}0&0\\ 1&0\\ 0&1\end{bmatrix}.$
By Algorithm 2, we can obtain the relation between $\gamma$ and $N$ (the number of 00 in feedback gain $K$; see Figure 2).
Figure 2: Relation between $\gamma$ and $N$
Specifically, when $\gamma=10$, we can obtain the following result (see Figure 3).
Figure 3: Feasible gain with $\gamma=10$
The optimal $W$ is given by
$W=\begin{bmatrix}2.1348&0&0&1.3084&0\\ 0&0.2698&0&0.6783&0\\ 0&0&1.0833&0&1.4155\\ 1.3084&0.6783&0&2.5604&0.0129\\ 0&0&1.4155&0.0129&1.8711\end{bmatrix},$
and the optimal decentralized feedback gain $K$ is given by
$K=\begin{bmatrix}0.6192&2.5269&0\\ 0&0&1.3068\par\end{bmatrix}.$
The responses of all the state variables are illustrated in Figure 4, and it can be seen that the stability of system is guaranteed.
Figure 4: System response
Example 2: Consider the linear system (57) with
$A=\begin{bmatrix}0.3079&0.1879&0.1797&0.2935&0.6537\\ 0.5194&0.2695&0.5388&0.9624&0.5366\\ 0.7683&0.4962&0.2828&0.9132&0.9957\\ 0.7892&0.7391&0.7609&0.5682&0.1420\\ 0.8706&0.1950&0.2697&0.4855&0.9753\end{bmatrix},B_{2}=\begin{bmatrix}0.6196&0.6414\\ 0.7205&0.9233\\ 0.2951&0.8887\\ 0.6001&0.6447\\ 0.7506&0.2956\end{bmatrix}$
$B_{1}=I_{5},C=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\end{bmatrix},D=\begin{bmatrix}0&0\\ 0&0\\ 1&0\\ 0&1\\ 0&0\end{bmatrix}.$
Letting $\gamma=10$, and by Algorithm 2, the optimal $W$ is given by
$W=\begin{bmatrix}2.5248&0&0&0&0&3.3422&2.2528\\ 0&3.0027&0&0&0&0&5.6741\\ 0&0&4.2745&0&0&0&4.4252\\ 0&0&0&0.9312&0&0&3.5367\\ 0&0&0&0&0.3125&1.7188&0\\ 3.3422&0&0&0&1.7188&14.0092&3.0413\\ 2.2528&5.6741&4.4252&3.5367&0&3.0413&30.9070\end{bmatrix},$
and the optimal decentralized feedback gain $K$ is given by
$K=\begin{bmatrix}1.3355&0&0&0&5.5233\\ 0.8946&1.8952&1.0401&3.8119&0\end{bmatrix}.$
The responses of all the state variables are illustrated in Figure 5, and the stability of system is guaranteed.
Figure 5: System response
Example 3: Consider the linear system (57) with
$A=\begin{bmatrix}0&1&0\\ 0&0&1\\ 0&0&0\end{bmatrix},B_{1}=I_{3},B_{2}=\begin{bmatrix}0.9315&0.7939\\ 0.9722&0.1061\\ 0.5317&0.7750\end{bmatrix},C=\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix},D=\begin{bmatrix}0&0\\ 1&0\\ 0&1\end{bmatrix}.$
We utilize Algorithm 2 and Algorithm 3 respectively and the results are shown below (see Figure 6 and Figure 7).
Figure 6: Convergence of Algorithm 2
Figure 7: Convergence of Algorithm 3
We can observe that the Algorithm 2 converges within 800 steps, while Algorithm 3 converges within only 80 steps.
8 Conclusion
This work investigates a decentralized LQ problem, specifically focusing on the inclusion of a sparsity constraint on the feedback gain. The formulation of approximate separable constraint optimization issues is presented, and a convex optimization framework is suggested to achieve the global optimizer for these problems. In the near future, an examination will be conducted on the decentralized LQ problem with a fixed sparse topology.