### v3

# Two-Timescale Optimization Framework for Decentralized Linear-Quadratic Optimal Control

Abstract
A $\mathcal{H}_{2}$-guaranteed decentralized linear-quadratic optimal control with convex parameterization and convex-bounded uncertainty is studied in this paper, where several sparsity promoting functions are added, respectively, into the $\mathcal{H}_{2}$ cost to penalize the number of communication links among decentralized controllers. Then, the sparse feedback gain is investigated to minimize the modified $\mathcal{H}_{2}$ cost together with the stability guarantee, and the corresponding main results are of three parts. First, the weighted-$\ell_{1}$ sparsity promoting function is of concern, and a two-timescale algorithm is developed based on the BSUM (Block Successive Upper-bound Minimization) framework and a primal-dual splitting approach. Second, the optimization problem induced by piecewise quadratic sparsity penalty is investigated, which exhibits an accelerated convergence rate. Third, the nonconvex sparse optimization problem with $\ell_{0}$-penalty is studied, which can be approximated by successive coordinatewise convex optimization problems.
{IEEEkeywords}
decentralized optimal control, linear-quadratic system, two-timescale algorithm, $\ell_{0}$-penalty
1 Introduction
During the last few decades, the concept of decentralized control has gained significant attention due to its appealing characteristics, such as the potential to lower communication costs and parallelizable processing. In particular, decentralized linear-quadratic (LQ) optimal control has been also a topic of general interest, which has been extensively investigated [3, 4]. Nevertheless, decentralized LQ problems present a significant challenge, as obtaining a closed-form solution is exceedingly difficult, with the exception of highly specific cases [5]. Optimization, an “end-to-end” method allowing us to optimize the performance criterion directly, has significantly contributed to the centralized LQ problems [6, 7, 8, 9, 10, 11, 12, 13] and decentralized LQ problems [14, 15, 16, 17], respectively. However, the question of how to design a controller with a general sparsity structure has yet to be resolved. Here, the term “sparsity structure” in this context pertains to the presence of zero elements within the stabilizing linear feedback gain matrix. In this paper, the $\mathcal{H}_{2}$-guaranteed cost decentralized control problem under convex parameterization and convex-bounded uncertainty is studied. Motivated by recent achievements in optimization-based LQ controls, the objective of this study is to establish a convex optimization framework for decentralized control incorporating a sparsity promoting penalty, allowing us to obtain a sparse feedback gain to lower the communication costs among controllers.
1.1 Related Works
1) Optimization-based LQ control. LQ optimal control is a landmark problem in the field of optimal control theory, and direct policy gradient is a hot topic of centralized LQ problems in recent years. If the full state is measured, the LQ performance criterion satisfies the Polyak-Lojasiewicz (PL) condition [12]. Hence, gradient-based methods can be ensured the convergence to the global optimal linear feedback gain exponentially in both model-based and model-free settings [6, 7, 11]. The challenge of LQ problems is significantly amplified when the state cannot be fully measured, mostly due to the absence of the PL condition and the resulting disconnection of feasible set [8, 9, 10, 13, 18].
On the other hands, a canonical decentralized LQ optimal control can be cast as a centralized problem where the stabilizing feedback gain matrix is restricted to lie in a particular subspace $\mathcal{K}$. If $\mathcal{K}$ denotes the block-diagonal matrix space, classic convex parametrization methods can be utilized to obtain an approximate convex problem [9], and ADMM (Alternating Direction Method of Multipliers) based method is shown to converge to the global optimizer of the abovementioned approximate convex problem [14, 17, 19]. If $\mathcal{K}$ possesses general sparsity structure, the SDP (Semidefinite Programming) relaxation perspective is of concern. If the SDP relaxation has a rank-$1$ solution, a globally linear optimal distributed controller can be recovered from the solution of SDP [15]. Moreover, an exact and inexact convex reformulation of sparse linear controller synthesis is introduced based on Youla parameterization and quadratically invariant assumption [41, 42], and infinite-dimensional convex optimization problems are proposed, which can be solved approximately by Laurent expansion. Additionally, zero-order policy optimization algorithm is utilized to optimize the decentralized LQ cost directly, and it can converge to a stationary point with a PAC (Probably Approximately Correct) guarantee [16].
2) Sparse optimization. Sparse optimization, a task that is intrinsically NP-hard, is a fundamental challenge in the fields of optimization, machine learning, image processing, and control theory. One popular method is the relaxed approximation method, for example, convex relaxed method [20, 21, 22], nonconvex relaxed method [23, 24] and DC (difference of convex functions) relaxed method [25, 26]. Due to an enhanced comprehension of the variational characteristics associated with the $\ell_{0}$-norm, numerous studies have addressed sparse optimization by directly optimizing the $\ell_{0}$-norm [27, 28]. Additionally, successive coordinatewise convex problem can approximate $\ell_{0}$-norm penalty problem [29].
3) Separable constrained optimization. It aims to solve the optimization problem (20) below. The most classic method to solve this problem is the ADMM method [30], which is based on the primal-dual framework. Additionally, there exists a wide range of variations of ADMM; see, for example, the work [31, 32, 33]. Currently, some dynamical system based methods are proposed to solve optimization problem (20), for instance, first-order dynamical system [34, 35, 36], second-order dynamical system [37, 38]. In particular, [39] introduces a new primal-dual splitting algorithms with provable nonergodic convergence rates based on differential equation solver approach, which is utilized in this paper.
1.2 Contributions
This paper examines decentralized LQ optimal control with general sparsity structure. If compared with existing results, the main contributions of the presented paper are as follows.
1) The optimal sparse feedback gain of the considered LQ problem can be represented by the global minimizer of a constrained optimization problem with sparsity penalty. Compared with [40, 17, 14], the feedback gain with general sparsity structure is investigated rather than the block-diagonal feedback gain, and instead of the infinite-dimensional optimization that is solved by finite-dimensional approximation method [41, 42, 43], the proposed optimization problems of this paper is finite-dimensional. Furthermore, in contrast to [15, 4, 44, 16], the algorithms introduced in this paper provide an optimality and stabilizing guarantee rather than heuristic algorithm or stationary-point-guarantee.
2) The results of this paper is of independent interest in optimization field. Note that the core algorithm used to solve the constrained convex optimization problem with weighted-$\ell_{1}$ sparsity promoting penalty is the one of [39], and whereas the methodology of [39] cannot be used directly due to the presence of a complicated subproblem, necessitating the development of an iterative algorithm based on BSUM methodology. To highlight that the BSUM framework consists of successive solving of subproblems induced by auxiliary functions, we design special auxiliary functions such that each subproblem exhibits a closed-form solution. Furthermore, we have established a direct relationship between the acceleration of solving decentralized LQ problems and the design of a strongly convex sparsity promoting function. To our best knowledge, this phenomenon has not been reported in the field of sparse optimization.
3) The variational properties of $\ell_{0}$-penalty is firstly studied for decentralized LQ problem. Motivated by the convex relaxation framework of [29], we show that the $\ell_{1}$-penalty induced problems can be seen as approximations of $\ell_{0}$-penalty induced problem in the field of decentralized LQ optimal control. Noting that [29] studies an unconstrained smooth optimization problem, the considered problem of this paper is constrained and nonsmooth. Hence, the conclusion and techniques of proof presented in our analysis do not exhibit parallel generalizations of those of [29]. Moreover, the interconnection between the proposed optimization problems is illustrated in Fig. 1 below.
The rest of this paper is organized as follows. In Section 2, the basic formulations and assumptions of decentralized LQ problem are proposed. In Section 3, we will introduce a two-timescale algorithm framework based on primal-dual splitting method to solve the weighted-$\ell_{1}$-sparsity promoting optimization problem, which encompasses a complicated subproblem. Such a subproblem is investigated in Section 4. Section 5 provides the convergence analysis of the proposed algorithm and studies quadratic sparsity promoting optimization problem, which can be regarded as an acceleration of weighted-$\ell_{1}$ promoting problem. In Section 6, we will study an $\ell_{0}$-sparsity promoting optimization problem which is nonconvex and discontinuous, and propose some theorems to characterize its variational properties.
$\ell_{0}$-sparsity promoting Section 6 weighted-$\ell_{1}$-sparsity promoting Section 3, 4, 5 quadratic sparsity promoting Section 5 coordinate convex problem Section 6 subproblem 1 $\Leftrightarrow$ weighted-$\ell_{1}$-sparsity promoting, subproblem 2 $\Leftrightarrow$ closed-form solution convex relaxationconvex relaxationaccelerationapproximationBSUM method
Figure 1: Interconnection between the proposed optimization problems
[…]
$\mathbf{prox}_{\tau g}^{\mathcal{Y}}(z):=\underset{y\in\mathcal{Y}}{\operatorname*{argmin}}\left\{g(y)+\frac{1}{2\tau}\left\|y-z\right\|^{2}\right\},\quad\forall z\in\mathbb{R}^{n}.$
For any $\alpha\in\mathbb{R}$, the $\alpha$-level set of a function $f:\mathbb{R}^{n}\to[-\infty,\infty]$ is the set ${\rm lev}_{\leq\alpha}f=\{x\in\mathbb{R}^{n}\colon f(x)\leq\alpha\}.$ We introduce the notation $M\lesssim N$, meaning $M\leq CN$ with some generic bounded constant $C>0$. For any subset $C\subseteq\mathbb{R}^{n}$, the indicator function of $C$ is defined to be the extended real-valued function given by
$\delta_{C}(x)=\left\{\begin{aligned} 0,~{}~{}&x\in C,\\ \infty,~{}~{}&x\notin C.\end{aligned}\right.$
2 Formulation
Consider a linear time-invariant system
$\dot{x}(t)$ $=Ax(t)+B_{2}u(t)+B_{1}w(t),$ (1)
$z(t)$ $=Cx(t)+Du(t)$
with state $x(t)\in\mathbb{R}^{n}$, input $u(t)\in\mathbb{R}^{m}$, exogenous disturbance input $w(t)\in\mathbb{R}^{l}$, controlled output $z(t)\in\mathbb{R}^{q}$ and matrices $A\in\mathbb{R}^{n\times n},B_{2}\in\mathbb{R}^{n\times m},B_{1}\in\mathbb{R}^{n\times l},C\in\mathbb{R}^{q\times n},D\in\mathbb{R}^{q\times m}$. The infinite-horizon LQ problem of this paper is to find a linear static state feedback gain $K\in\mathbb{R}^{m\times n}$ such that
$u(t)=-Kx(t)$
minimizes the following performance criterion
$J=\int_{0}^{\infty}z(t)^{\top}z(t){\rm d}t.$ (2)
By [45], to optimize (2), it is equivalent to minimize the $\mathcal{H}_{2}$-norm of the transfer function
$H(s)=(C-DK)(sI_{n}-A+B_{2}K)^{-1}B_{1}$
from $w$ to $z$, and the performance criterion can be reformulated as
$J(K)=\|H(s)\|_{\mathcal{H}_{2}}^{2}={\rm Tr}\left((C-DK)W_{c}(C-DK)^{\top}\right),$
where $W_{c}$ is the controllability Gramian associated with the closed-loop system. In this paper, the stabilizing controller below refers to stabilizing system (1) internally; see Section 14 of [46] for details. The following assumption is introduced in [40, 16, 47].
Assumption 1.
Assume that $C^{\top}D=0,D^{\top}D\succ 0,B_{1}B_{1}^{\top}\succ 0,(A,B_{2})$ is stabilizable and $(A,C)$ has no unobservable modes on the imaginary axis.
We introduce the following notations with $p=m+n$
$F=\begin{bmatrix}A&B_{2}\\ 0&0\end{bmatrix}\in\mathbb{R}^{p\times p},~{}G=\begin{bmatrix}0\\ I_{m}\end{bmatrix}\in\mathbb{R}^{p\times m},$
$Q=\begin{bmatrix}B_{1}B_{1}^{\top}&0\\ 0&0\end{bmatrix}\in\mathbb{S}^{p},~{}R=\begin{bmatrix}C^{\top}C&0\\ 0&D^{\top}D\end{bmatrix}\in\mathbb{S}^{p}.$
Assumption 2.
The parameter $F$ is unknown but convex-bounded, i.e., $F$ belongs to a polyhedral domain, which is expressed as a convex combination of the extreme matrices, where $F=\sum_{i=1}^{M}\xi_{i}F_{i},\xi_{i}\geq 0,\sum_{i=1}^{M}\xi_{i}=1$, and $F_{i}=\begin{bmatrix}A_{i}&B_{2,i}\\ 0&0\end{bmatrix}\in\mathbb{R}^{p\times p}$ denotes the extreme vertex of the uncertain domain.
Let block matrix
$W=\begin{bmatrix}W_{1}&W_{2}\\ W_{2}^{\top}&W_{3}\end{bmatrix}\in\mathbb{S}^{p}$
with $W_{1}\in\mathbb{S}_{++}^{n},W_{2}\in\mathbb{R}^{n\times m},W_{3}\in\mathbb{S}^{m}$. Regard $\Theta_{i}(W)=F_{i}W+WF_{i}^{\top}+Q$ as a block matrix, i.e.,
$\Theta_{i}(W)=\begin{bmatrix}\Theta_{i,1}(W)&\Theta_{i,2}(W)\\ \Theta_{i,2}^{\top}(W)&\Theta_{i,3}(W)\end{bmatrix}\in\mathbb{S}^{p}.$
The following theorem introduces a subset of stabilizing controller gains, and finds an upper bound of performance criterion (2).
Theorem 1.
([40]) One can define the set
$\mathscr{C}$ =$=\{W\in\mathbb{S}^{p}:W\succeq 0,\Theta_{i,1}(W)\preceq 0,\forall i=1,2,\ldots,M\},$
and let $\mathscr{K}=\{K=W_{2}^{\top}W_{1}^{-1}\colon W\in\mathscr{C}\}$. Then,
  1. 1.
    $K\in\mathscr{K}$ stabilizes the closed-loop system;
  2. 2.
    $K\in\mathscr{K}$ gives
    $\langle R,W\rangle\geq\|H_{i}(s)\|_{\mathcal{H}_{2}}^{2},~{}~{}i=1,\dots,M,$
    where $\|H_{i}(s)\|_{\mathcal{H}_{2}}$ represents the $\mathcal{H}_{2}$-norm with respect to the $i$-th extreme system.
Proof: The proof is familiar with that of Theorem 1 of [40], but for the sake of completeness, we will include short proof here. For statement 1, $\Theta_{i,1}\preceq 0$ implies that
$(A_{i}-B_{2,i}W_{2}^{\top}W_{1}^{-1})W_{1}$ $+W_{1}(A_{i}-B_{2,i}W_{2}^{\top}W_{1}^{-1})^{\top}$ (3)
$+B_{1}B_{1}^{\top}\preceq 0.$
It is obvious that $K=W_{2}^{\top}W_{1}^{-1}$ stabilizes the extreme system. To ensure the stability over the entire uncertain domain that is convex, it suffices to check the stability at the vertices of the convex polyhedron. Since, for all $i=1,\dots,M$, stability of extreme system $i$ is guaranteed, the stability of (1) holds.
For statement 2, we have
$W_{c,i}=\int_{0}^{+\infty}e^{(A-B_{2,i}K)t}B_{1}B_{1}^{\top}e^{(A-B_{2,i}K)^{\top}t}{\rm d}t.$
Hence, by (3), $W_{1}\succeq W_{c,i}$ holds. Then, based on Schur complement lemma, $W\succeq 0$ implies that
$W_{3}\succeq W_{2}^{\top}W_{1}^{-1}W_{2}=KW_{1}K^{\top}\succeq KW_{c,i}K^{\top}.$
Thus, combining with the assumption $C^{\top}D=0$, it follows that
$\langle R,W\rangle$ $={\rm Tr}(CW_{1}C^{\top}+DW_{3}D^{\top})$
$\geq{\rm Tr}(CW_{c,i}C^{\top}+DKW_{c,i}K^{\top}D^{\top})$
$={\rm Tr}((C-DK)W_{c,i}(C-DK)^{\top}).$
[…]
Remark 1.
If $z(t)$ exhibits the following form which is studied in [14]
$z(t)=\begin{bmatrix}Q^{\frac{1}{2}}\\ 0\end{bmatrix}x(t)+\begin{bmatrix}0\\ R^{\frac{1}{2}}\end{bmatrix}u(t)$
with $Q,R\succ 0$, then $C^{\top}C\succ 0$, $D^{\top}D\succ 0$ and $C^{\top}D=0$ naturally hold. In fact, when $F$ is known, the classic infinite-horizon LQ problem
$\min_{K\in\mathbb{R}^{m\times n}}{}{}$ $J(K)$ (4)
${\rm s.t.}{}{}$ $K\in\mathcal{S}$
with stabilizing feedback gain set $\mathcal{S}$ is equivalent to the following convex optimization problem [48]
$\min_{W\in\mathbb{S}^{p}}{}{}$ $\langle R,W\rangle$ (5)
${\rm s.t.}{}{}$ $W\in\mathscr{C}.$
However, we want to expand the range of applicability in cases where $F$ is unknown but convex-bounded; see Assumption 2. In this case, by Theorem 1, the optimization problem (5) is no longer equivalent to (4), but the objective function of (5) is the upper bound of $J(K)$.
In this study, we focus on a general decentralized control problem, assuming that the feedback gain exhibits a general sparsity structure rather than a block diagonal structure [40]. The fundamental idea in our paper is that the sparse stabilizing feedback gain, regardless of its structure, can be achieved by the design of a diagonal matrix $W_{1}$ and a sparse matrix $W_{2}$ with some sparse structure. Then, feedback gain $K=W_{2}^{\top}W_{1}^{-1}$ can preserve the sparsity structure of $W_{2}$ and stabilize the closed-loop system.
Definition 1.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be proper and lower semi-continuous function. Then $f$ is said to be a sparsity promoting function provided that:
  1. 1.
    $f(0)=0$ and f achieves its global minimum at origin;
  2. 2.
    the set $\partial f(0)$ contains at least one nonzero element.
In this paper, we will use the weighted-$\ell_{1}$-norm, the piecewise quadratic function, and the $\ell_{0}$-norm as sparsity promoting functions, which correspond to different optimization problems. Letting $N=n(n-1)/2$, $e_{i}=(\mathbf{0}_{i-1}^{\top},1,\mathbf{0}_{p-i}^{\top})^{\top}\in\mathbb{R}^{p}$,
$E_{i}^{j}=\begin{bmatrix}e_{i}^{\top}\\ \vdots\\ e_{i}^{\top}\end{bmatrix}\in\mathbb{R}^{j\times p},~{}~{}\hat{E}_{i}^{j}=\begin{bmatrix}e_{i}\\ \vdots\\ e_{j}\end{bmatrix}\in\mathbb{R}^{p(j-i+1)},$
$V_{*1}=\begin{bmatrix}V_{11}\\ \vdots\\ V_{N1}\end{bmatrix}\in\mathbb{R}^{N\times p},~{}~{}V_{*2}=\begin{bmatrix}V_{12}\\ \vdots\\ V_{N2}\end{bmatrix}\in\mathbb{R}^{Np},$
we denote
$V_{*1}=\begin{bmatrix}E_{1}^{n-1}\\ E_{2}^{n-2}\\ \vdots\\ E_{n-1}^{1}\end{bmatrix},~{}~{}V_{*2}=\begin{bmatrix}\hat{E}_{2}^{n}\\ \hat{E}_{3}^{n}\\ \vdots\\ \hat{E}_{n}^{n}\end{bmatrix}.$ (6)
Then, for the weighted-$\ell_{1}$-norm of a matrix $P$
$g(P)=\sum_{i,j}w_{ij}|P_{ij}|$ (7)
with $w_{ij}>0$ for all $i,j$, we introduce the following optimization problem
$\min_{W,P}~{}~{}$ $\langle R,W\rangle+\gamma g(P)$ (8)
${\rm s.t.}~{}~{}$ $W\in\mathbb{S}^{p}_{+},$ (9)
$\Psi_{i}\in\mathbb{S}_{+}^{n},\forall i=1,\dots,M,$ (10)
$-V_{j1}WV_{j2}=0,\forall j=1,\dots,N,$ (11)
$V_{1}WV_{2}^{\top}-P=0$ (12)
with $\gamma>0$, $V_{1}=[0,I_{m}],V_{2}=[I_{n},0]$ and
$\Psi_{i}=-V_{2}(F_{i}W+WF_{i}^{\top}+Q)V_{2}^{\top},\forall i=1,\dots,M.$
Remark 2.
The above problem is the core optimization problem of this paper, where the constraints (9) and (10) are equivalent to $W\in\mathscr{C}$. Additionally, by the definitions of $V_{*1}$ and $V_{*2}$, the constraint (11) implies that $W_{1}$ is a diagonal matrix, and the constraint (12) is equivalent to $P=W^{\top}_{2}$ so that sparsity penalty term $\gamma g(P)=\gamma g(W^{\top}_{2})$. Together with constraints (9)-(12) and Theorem 1, the matrix $K=W_{2}^{\top}W_{1}^{-1}$ is a stabilizing feedback gain with general sparsity structure.
Remark 3.
In fact, let
$D_{k}=\begin{bmatrix}0\\ I_{k}\end{bmatrix}\in\mathbb{R}^{n\times k},$
then the $N$ equality constraints
$V_{j1}WV_{j2}=0,~{}j=1,\dots,N$ (13)
can be converted into $n-1$ equality constraint:
$e_{k}V_{2}WV_{2}^{\top}D_{n-k}=0,~{}k=1,\dots,n-1$
with $e_{k}=(0,\dots,1,\dots,0)$. For the convenience and simplicity of formulas’ deducing, we still use $N$ equality constraints (i.e., (11)) to describe the diagonal property of matrix $W_{1}$.
Remark 4.
The group-sparsity promoting function can be defined likewise. For example, letting the matrix $P\in\mathbb{R}^{m\times n}$ be partitioned into submatrices $P_{ij}\in\mathbb{R}^{m_{i}\times n_{j}}$, the group weighted-$\ell_{1}$-norm can be defined as
$g_{\rm group}(P)=\sum_{i,j}w_{ij}\|P_{ij}\|_{F}.$
By replacing $\gamma g(P)$ of (8) by $\gamma g_{\rm group}(P)$ and modifying the constraint (11) such that $W_{1}$ is a block-diagonal matrix with proper size, we then obtain a group sparse feedback gains $K=W_{2}^{\top}W_{1}^{-1}$. Remarkably, all the results of this paper could be generalized into the group-sparsity-promoting setting, and for the sake of simplicity we only focus on how to obtain a sparse (instead of group sparse) feedback gain in the following sections.
Remark 5.
Generally, a sparse (or group-sparse) feedback gain $K$ can lead to a multi-agent decentralized control system. For example, given a group-sparse feedback gain
$K=\begin{bmatrix}K_{11}&\cdots&K_{1q}\\ \vdots&\ddots&\vdots\\ K_{p1}&\cdots&K_{pq}\end{bmatrix}\in\mathbb{R}^{m\times n}$
with $K_{ij}\in\mathbb{R}^{m_{i}\times n_{j}}$ and $\sum_{i=1}^{p}m_{i}=m,\sum_{j=1}^{q}n_{j}=n$, it can induce a $p$-agents decentralized system. Specifically, the global control action is composed of local control actions: $u(t)=[u_{1}(t)^{\top},\dots,u_{p}(t)^{\top}]^{\top}$, where $u_{i}(t)\in\mathbb{R}^{m_{i}}$ is the control input of agent $i$. At time $t$, agent $i$ directly observes a partial state $x_{\mathcal{I}_{i}}(t)$, and $\mathcal{I}_{i}$ is given by
$\mathcal{I}_{i}=\bigcup_{j=1,\dots,q,~{}K_{ij}\neq 0}\left\{n_{j-1}+1,\dots,n_{j}\right\}$
with $n_{0}=0$. Here, $\mathcal{I}_{i}$ is a fixed subset of $[n]\doteq\{1,\dots,n\}$ and $x_{\mathcal{I}_{i}}(t)$ denotes the subvector of $x(t)$ with indices in $\mathcal{I}_{i}$. That is why in this paper we investigate the decentralized LQ problem by studying optimization problem
$\min_{W\in\mathbb{S}^{p}}{}{}$ $\langle R,W\rangle+\gamma s(V_{1}WV_{2}^{\top})$ (14)
${\rm s.t.}{}{}$ $W\in\mathscr{C}$
with sparsity promoting function $s(\cdot)$. It is important to note that the $p$-agents decentralized systems in [40, 17, 14] satisfy
$\mathcal{I}_{i}\cap\mathcal{I}_{j}=\emptyset,~{}~{}\forall i,j\in[p],i\neq j.$
However, $\mathcal{I}_{i}$ and $\mathcal{I}_{j}$ may overlap in our setting, and hence the considered problem of this paper is a generalization of those of [40, 17, 14].
Optimization problem (8)-(12) can be equivalently expressed in a compact form
$\min_{W,P}{}{}$ $\langle R,W\rangle+\gamma g(P)$ (15)
${\rm s.t.}{}{}$ $\mathcal{G}(W)\in\mathcal{K},$
$V_{1}WV_{2}^{\top}-P=0,$
where
$\mathcal{G}(W)=(W,\Psi_{1},\dots,\Psi_{M},V_{11}WV_{12},\dots,V_{N1}WV_{N2}),$
and convex cone $\mathcal{K}$ is defined as $\mathcal{K}=\mathbb{S}_{+}^{p}\times\mathbb{S}_{+}^{n}\times\cdots\times\mathbb{S}_{+}^{n}\times 0\times\cdots\times 0.$ Since the positive semi-definite cone is self-dual, it is straightforward to express the dual cone $\mathcal{K}^{*}$ as $\mathcal{K}^{*}=\mathbb{S}_{+}^{p}\times\mathbb{S}_{+}^{n}\times\cdots\times\mathbb{S}_{+}^{n}\times\mathbb{R}\times\cdots\times\mathbb{R}.$
Assumption 3.
Problem (15) is strictly feasible, i.e., Slater’s condition is satisfied and strong duality holds.
Remark 6.
This assumption is popular in solving decentralized LQ problems via optimization landscape [40, 17, 14]. The question of when the aforementioned assumption is valid remains unresolved, and [14] offers some insights from the standpoint of block-diagonal strategy when $(A,B_{2})$ is known. Specifically, let $D$ be the feasible set of optimization problem (15), and denote
$\widetilde{D}=\left\{W_{2}\in\mathbb{R}^{n\times m}\colon W_{2}=\mathrm{blocdiag}\{W_{2,1},\dots,W_{2,p}\}\right\}$
with $W_{2,j}\in\mathbb{R}^{n_{j}\times m_{j}}$ for all $j=1,\dots,p$ and $\sum_{j=1}^{p}m_{j}=m$, $\sum_{j=1}^{p}n_{j}=n$. It is demonstrated that, if system (1) is fully actuated or weakly coupled in terms of topological connections or weakly coupled in terms of dynamical interactions (see Definition 4, Definition 5 and Definition 8 of [14]), then $D\cap\widetilde{D}$ (a subset of $D$) is nonempty (see Proposition 1-3 of [14]). Hence, if system (1) is fully actuated or weakly coupled in terms of topological connections or weakly coupled in terms of dynamical interactions, problem (15) is feasible. Combining with the fact that $B_{1}B_{1}^{\top}\succ 0$ and $W_{3}$ is a free variable to be chosen such that $W\succ 0$ (utilizing Schur complement lemma), problem (15) is strictly feasible, i.e., Assumption 3 holds. It is important to note that the feasible set of (15) is larger than the ones in [40, 17, 47, 14], since there is no structural constraints on $W_{2}$ in this paper.
Remark 7.
By modifying optimization problem (15) as
$\min_{W,P}{}{}$ $\langle R,W\rangle$ (16)
${\rm s.t.}{}{}$ $\mathcal{G}(W)\in\mathcal{K},$
$[V_{1}WV_{2}^{\top}]_{ij}=0,~{}(i,j)\in\mathcal{U}\subseteq[m]\times[n],$
with $[\cdot]_{ij}$ the $(i,j)$-th element of a matrix and $\mathcal{U}$ a given subset, the methods proposed in this paper can be used to decentralized LQ problems with fixed communication topology, which have been widely studied [15, 16, 41, 42]. However, our paper is of independent interests for problems with fixed communication topology. Specifically, this paper provides an optimal-guarantee w.r.t. (16) rather than heuristic method [15] or stationary-point-guarantee [16]. Additionally, the proposed problem (16) is finite-dimensional, in comparison to the infinite-dimensional optimization [41, 42]. Moreover, if we only hope that agent $i_{k}$ cannot observe the state $x_{j_{k}}(t)$ ($k=1,\dots,p$) directly rather than giving a fixed communication topology, we can modify problem (15) as
$\min_{W,P}{}{}$ $\langle R,W\rangle+\gamma g(P)$ (17)
${\rm s.t.}{}{}$ $\mathcal{G}(W)\in\mathcal{K},$
$V_{1}WV_{2}^{\top}-P=0,$
$[V_{1}WV_{2}^{\top}]_{i_{k}j_{k}}=0,~{}k=1,\dots,p$
with $(i_{k},j_{k})\in[m]\times[n]$ for $k=1,\dots,p$. To the best of the authors’ knowledge, such special formulation, i.e., the design of a sparse feedback gain $K$ with forcing $[K]_{i_{k}j_{k}}=0$ for $k=1,\dots,p$, has not been reported in existing literature. It is worth to underline that problem (17) can be solved analogously by the proposed algorithms in Section 3, 4.
3 Two-Timescale Optimization Algorithm Framework
Let $\widetilde{W}={\rm vec}(W),\widetilde{P}={\rm vec}(P)$, then by the definition of indicator function $\delta_{\Gamma_{+}^{n}}(\cdot)$, (15) is equivalent to the following optimization problem
$\min_{\widetilde{W},\widetilde{P}}{}{}$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})+\gamma g(\widetilde{P})$ (18)
${\rm s.t.}{}{}$ $\left(V_{i2}^{\top}\otimes V_{i1}\right)\widetilde{W}=0,~{}~{}i=1,\dots,M,$
$(V_{2}\otimes V_{1})\widetilde{W}-\widetilde{P}=0,$
where $\widetilde{\Psi}_{i}={\rm vec}(\Psi_{i})$ and $\Gamma_{+}^{p}=\left\{{\rm vec}(W)\colon W\in\mathbb{S}_{+}^{p}\right\}$. Denoting $\mathcal{A},\mathcal{B},\widehat{\mathcal{A}}$ with
$\mathcal{A}=\begin{pmatrix}V_{12}^{\top}\otimes V_{11}\\ \vdots\\ V_{M2}^{\top}\otimes V_{M1}\\ V_{2}\otimes V_{1}\end{pmatrix}\doteq\begin{pmatrix}\widehat{\mathcal{A}}\\ V_{2}\otimes V_{1}\end{pmatrix},~{}~{}\mathcal{B}=\begin{pmatrix}0\\ \vdots\\ 0\\ -I_{mn}\end{pmatrix},$
optimization problem (18) is summarized in the following form
$\min_{\widetilde{W},\widetilde{P}}{}{}$ $f(\widetilde{W})+h(\widetilde{P})$ (19)
${\rm s.t.}{}{}$ $\mathcal{A}\widetilde{W}+\mathcal{B}\widetilde{P}=0,$
where
$f(\widetilde{W})$ $=\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M}),$
$h(\widetilde{P})$ $=\gamma g(\widetilde{P}).$
To solve optimization problem (19), we briefly review the basic setup of primal-dual splitting method; see [39] for a comprehensive review. The following separable convex optimization problem is considered in [39]:
$\min_{x\in\mathcal{X},y\in\mathcal{Y}}F(x,y)=\bar{f}(x)+\bar{g}(y)~{}~{}{\rm s.t.}~{}Ax+By=b,$ (20)
where $\mathcal{X}\subseteq\mathbb{R}^{m}$ and $\mathcal{Y}\subseteq\mathbb{R}^{n}$ are two closed convex sets, and $\bar{f}$ and $\bar{g}$ are two proper closed convex functions. Assume that $\bar{g}$ is closed and $\mu_{g}$-strongly convex with $\mu_{g}>0$ and $\bar{f}=f_{1}+f_{2}$ with closed convex $f_{2}$, $\mu_{f}$-strongly convex and $L_{f}$-smooth $f_{1}$ and $0\leq\mu_{f}\leq L_{f}<\infty$. Then, the following scheme is proposed to solve problem (20)
$u_{k}=(x_{k}+\alpha_{k}v_{k})/(1+\alpha_{k}),$ (21)
$d_{k}=\nabla f_{1}(u_{k})+A^{\top}\lambda_{k},$
$v_{k+1}=\underset{v\in\mathcal{X}}{\operatorname*{argmin}}\left\{f_{2}(v)+\langle d_{k},v\rangle+\frac{\alpha_{k}}{2\theta_{k}}\left\|Av+Bw_{k}-b\right\|^{2}\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.+\frac{\widetilde{\eta}_{f,k}}{2\alpha_{k}}\left\|v-\widetilde{v}_{k}\right\|^{2}\right\},$
$x_{k+1}=(x_{k}+\alpha_{k}v_{k+1})/(1+\alpha_{k}),$
$\bar{\lambda}_{k+1}=\lambda_{k}+\alpha_{k}/\theta_{k}(Av_{k+1}+Bw_{k}-b),$
$y_{k+1}=\mathbf{prox}_{\tau_{k}\bar{g}}^{\mathcal{Y}}(\widetilde{y}_{k}-\tau_{k}B^{\top}\bar{\lambda}_{k+1}),\quad\tau_{k}=\alpha_{k}^{2}/\eta_{g,k},$
$w_{k+1}=y_{k+1}+(y_{k+1}-y_{k})/\alpha_{k},$
$\lambda_{k+1}=\lambda_{k}+\alpha_{k}/\theta_{k}(Av_{k+1}+Bw_{k+1}-b),$
where $(\tilde{v}_{k},\tilde{y}_{k},\tilde{\eta}_{f,k},\eta_{g,k})$ is given by
$\eta_{g,k}=(\alpha_{k}+1)\beta_{k}+\mu_{g}\alpha_{k},$ (22)
$\widetilde{y}_{k}=y_{k}+\frac{\alpha_{k}\beta_{k}}{\eta_{g,k}}(w_{k}-y_{k}),$ (23)
$\widetilde{\eta}_{f,k}=\gamma_{k}+\mu_{f}\alpha_{k},$ (24)
$\widetilde{v}_{k}=\frac{1}{\widetilde{\eta}_{f,k}}(\gamma_{k}v_{k}+\mu_{f}\alpha_{k}u_{k}),$ (25)
and $\{(u_{k},d_{k},v_{k},x_{k},\bar{\lambda}_{k},y_{k},w_{k},\lambda_{k})\}$ is the iteration sequence. To emphasis, $\alpha_{k}>0$ is the step size and the parameter system is given by
$\frac{\theta_{k+1}-\theta_{k}}{\alpha_{k}}=-\theta_{k+1},$ (26)
$\frac{\gamma_{k+1}-\gamma_{k}}{\alpha_{k}}=\mu_{f}-\gamma_{k+1},$
$\frac{\beta_{k+1}-\beta_{k}}{\alpha_{k}}=\mu_{g}-\beta_{k+1}$
with initial conditions $\theta_{0}=1,\gamma_{0}>0$ and $\beta_{0}>0$.
Remark 8.
In fact, the following differential inclusion
[…]
is the continuous-time version of algorithm (21), where
$\mathcal{L}(x,y,\lambda)=\bar{f}(x)+\bar{g}(y)+\left\langle\lambda,Ax+By-b\right\rangle.$ (28)
It can be observed that differential inclusion (27) encompasses two Nesterov-type differential inclusions pertaining to primal variables $x$ and $y$, as well as a gradient flow about the dual variable $\lambda$. Besides, an alternative presentation of (27) reads as
[…]
Therefore, (21) represents the semi-implicit Euler discretization of differential inclusion (29), which justifies its classification as a differential equation solver approach. We utilize this particular primal-dual splitting method instead of classic ADMM mainly because we can establish the relationship between the acceleration of solving decentralized LQ problems and the design of a strongly convex sparsity promoting function; this will be discussed in the following sections.
Remark 9.
[…]
Back to our topic (optimization problem (19)), denote
$f_{1}(\widetilde{W})$ $=\langle{\rm vec}(R),\widetilde{W}\rangle,$
$f_{2}(\widetilde{W})$ $=\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M}).$
We first consider the iteration of variable $\widetilde{P}\in\mathbb{R}^{mn\times 1}$. By (21), we have
$\widetilde{P}_{k+1}=\mathbf{prox}_{\tau_{k}h}^{\mathbb{R}^{mn\times 1}}\left(\widetilde{P}_{k}+\frac{\alpha_{k}\beta_{k}}{\eta_{g,k}}(w_{k}-\widetilde{P}_{k})-\tau_{k}\mathcal{B}^{\top}\bar{\lambda}_{k+1}\right)$
Denote
$\Omega^{k}={\operatorname*{vec}}^{-1}\left(\widetilde{P}_{k}+\frac{\alpha_{k}\beta_{k}}{\eta_{g,k}}(w_{k}-\widetilde{P}_{k})-\tau_{k}\mathcal{B}^{\top}\bar{\lambda}_{k+1}\right),$ (30)
$\rho_{k}=\frac{1}{\tau_{k}}$, $a_{k}=\frac{\gamma}{\rho_{k}}w_{ij}$, and let
$P_{ij;k+1}^{*}$ $\doteq\underset{P_{ij}\in\mathbb{R}}{\operatorname*{argmin}}\left\{\gamma w_{ij}|P_{ij}|+\frac{\rho_{k}}{2}\left(P_{ij}-\Omega^{k}_{ij}\right)^{2}\right\}$ (31)
[…]
By Example 6.19 of [49], we can deduce that $\widetilde{P}_{k+1}=(P_{11;k+1}^{*},\dots,P_{mn;k+1}^{*})^{\top}$ which exhibits naturally sparsity structure.
4 Solving Inner Layer Optimization Problem
Algorithm (21) consists of two subproblems: updating $v_{k+1}$ and $y_{k+1}$. The solution to the latter can be obtained explicitly using equation (31), while the solution to the former will be discussed in this section.
By the update rule of $v_{k+1}$ in (21) and the definitions of $f_{2}$ and $b$, when employing the symbol system of this paper, the subproblem in (21) can be transformed to
$\min_{{\rm vec}(W)}{}$ $\langle d_{k},{\rm vec}(W)\rangle+\frac{\alpha_{k}}{2\theta_{k}}\left\|\mathcal{A}{\rm vec}(W)+\mathcal{B}w_{k}\right\|^{2}$ (32)
$+\frac{\widetilde{\eta}_{f,k}}{2\alpha_{k}}\left\|{\rm vec}(W)-\widetilde{v}_{k}\right\|^{2}$
${\rm s.t.}{}$ $\mathrm{vec}(W)\in\Gamma_{+}^{p},$
$\mathrm{vec}(\Psi_{i})\in\Gamma_{+}^{n},~{}i=1,\dots,M.$
By Assumption 3, for all $k$, problem (32) satisfies Slater’s condition and strong duality holds. Since $W,\Psi_{i}$ are symmetric matrices, there exist matrices $T_{1},T_{2}$ such that
${\rm svec}(W)$ $=T_{1}{\rm vec}(W)~{}{\rm with}~{}T_{1}T_{1}^{\top}=I_{\frac{p(p+1)}{2}},$
${\rm svec}(\Psi_{i})$ $=T_{2}{\rm vec}(\Psi_{i})~{}{\rm with}~{}T_{2}T_{2}^{\top}=I_{\frac{n(n+1)}{2}},$
where $\mathrm{svec}(\cdot)$ is the vectorization operator on $\mathbb{S}^{n}$ defined by
${\rm svec}(S)=\left[s_{11},s_{12},\cdots,s_{n1},\cdots,s_{nn}\right]^{\top}\in\mathbb{R}^{\frac{n(n+1)}{2}}.$
Inversely, there exist matrix $D_{1},D_{2}$ such that
${\rm vec}(W)=D_{1}{\rm svec}(W),~{}~{}{\rm vec}(\Psi_{i})=D_{2}{\rm svec}(\Psi_{i}).$
Remark 10.
We partition the identity matrix $I_{N}$ ($N=n(n-1)/2$) as follows
$I_{N}=(u_{11},u_{21},\dots,u_{n1},u_{22},\dots,u_{n2},\dots,u_{nn}).$
Then, for every matrix $A\in\mathbb{S}^{n}$, the elimination matrix $L$ (which performs the transformation ${\rm svec}(A)=L{\rm vec}(A)$) can be explicitly expressed as
$L=\sum_{i\geq j}u_{ij}({\rm vec}E_{ij})^{\top}=\sum_{i\geq j}(u_{ij}\otimes e_{j}^{\top}\otimes e_{i}^{\top})$
with $E_{ij}=e_{i}e_{j}^{\top}$. By Lemma 3.2 of [50], $L^{\top}=L^{\dagger}$, and naturally $LL^{\top}=I_{N}$.
Inversely, there exist matrix $D_{1}\in\mathbb{R}^{p^{2}\times\frac{p(p+1)}{2}}$ and $D_{2}\in\mathbb{R}^{n^{2}\times N}$ such that
${\rm vec}(W)=D_{1}{\rm svec}(W),~{}~{}{\rm vec}(\Psi_{i})=D_{2}{\rm svec}(\Psi_{i}).$
In addition, $D_{1}$ can be explicitly expressed as
$D_{1}=\sum_{i\geq j}u_{ij}{\rm vec}(T_{ij})^{\top},$
where $T_{ij}$ is a $p\times p$ matrix with $1$ in position $(i,j),(j,i)$ and zero elsewhere.
Then, by utilizing $\mathrm{svec}(\cdot)$ operator and denoting ($\widetilde{\Gamma}_{+}^{n}$ is defined analogously)
$\widetilde{\Gamma}_{+}^{p}=\left\{{\rm svec}(W)\colon W\in\mathbb{S}_{+}^{p}\right\},$
optimization (32) can be readily converted into the subsequent problem
$\min_{{\rm svec}(W)}{}$ $\langle d_{k},D_{1}{\rm svec}(W)\rangle+\frac{\alpha_{k}}{2\theta_{k}}\left\|\mathcal{A}D_{1}{\rm svec}(W)+\mathcal{B}w_{k}\right\|^{2}$ (33)
$+\frac{\widetilde{\eta}_{f,k}}{2\alpha_{k}}\left\|D_{1}{\rm svec}(W)-\widetilde{v}_{k}\right\|^{2}$
${\rm s.t.}{}$ ${\rm svec}(W)\in\widetilde{\Gamma}_{+}^{p}$
${\rm svec}(\Psi_{i})\in\widetilde{\Gamma}_{+}^{n},~{}i=1,\dots,M,$
where ${\rm svec}(\Psi_{i})=T_{2}{\rm vec}(\Psi_{i}).$ The primal problem of (33) presents a level of difficulty that warrants exploration of its dual problem. The Lagrangian of (33) is defined as
$\mathcal{L}({\rm svec}(W);X)$ (34)
$=$ $\langle d_{k},D_{1}{\rm svec}(W)\rangle+\frac{\alpha_{k}}{2\theta_{k}}\left\|\mathcal{A}D_{1}{\rm svec}(W)+\mathcal{B}w_{k}\right\|^{2}$
$+\frac{\widetilde{\eta}_{f,k}}{2\alpha_{k}}\left\|D_{1}{\rm svec}(W)-\widetilde{v}_{k}\right\|^{2}-\left\langle X_{0},{\rm svec}(W)\right\rangle$
$+\left\langle(X_{1}+\dots+X_{M}),T_{2}(V\otimes V){\rm vec}(Q)\right\rangle$
$+\sum_{i=1}^{M}\left\langle T_{2}\left((V\otimes VF_{i})+(VF_{i}\otimes V)\right)D_{1}{\rm svec}(W),X_{i}\right\rangle,$
where $X=(X_{0},X_{1},\dots,X_{M})\in\widetilde{\Gamma}_{+}^{p}\times\widetilde{\Gamma}_{+}^{n}\times\dots\times\widetilde{\Gamma}_{+}^{n}$ is the Lagrange multiplier. Define
$\tilde{d}_{k}=D_{1}^{\top}d_{k}-X_{0}$
$~{}~{}~{}~{}~{}~{}+D_{1}^{\top}\sum_{i=1}^{M}\left[(V\otimes VF_{i})^{\top}+(VF_{i}\otimes V)^{\top}\right]T_{2}^{\top}X_{i}$
$~{}~{}~{}\doteq\mathcal{L}_{k}(X_{0},X_{1},\dots,X_{M}),$
$\sigma_{1}=\frac{\alpha_{k}}{2\theta_{k}},~{}\sigma_{2}=\frac{\widetilde{\eta}_{f,k}}{2\alpha_{k}},~{}\tilde{b}_{k}=\mathcal{B}w_{k},$
and by rearranging (34) we can derive
$\mathcal{L}({\rm svec}(W);X)$
$=$ $\langle\tilde{d}_{k},{\rm svec}(W)\rangle+\sigma_{1}\left\|\mathcal{A}D_{1}{\rm svec}(W)+\tilde{b}_{k}\right\|^{2}$
$+\sigma_{2}\left\|D_{1}{\rm svec}(W)-\widetilde{v}_{k}\right\|^{2}$
$+\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle.$
It may be inferred that the Lagrange dual function $\theta(X)$ is derived by
$\theta(X)$ $=\min_{{\rm svec}(W)}\mathcal{L}({\rm svec}(W);X)$ (35)
$=\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle$
$~{}~{}~{}+\min_{{\rm svec}(W)}\hat{\theta}({\rm svec}(W))$
with
$\hat{\theta}({\rm svec}(W))=$ $\langle\tilde{d}_{k},{\rm svec}(W)\rangle+\sigma_{1}\left\|\mathcal{A}D_{1}{\rm svec}(W)+\tilde{b}_{k}\right\|^{2}$
$+\sigma_{2}\left\|D_{1}{\rm svec}(W)-\widetilde{v}_{k}\right\|^{2}.$
Letting $\nabla\hat{\theta}({\rm svec}(W))=0$, and denoting
$M=2\sigma_{1}(\mathcal{A}D_{1})^{\top}(\mathcal{A}D_{1})+2\sigma_{2}D_{1}^{\top}D_{1},$
$q_{k}=-2\sigma_{1}(\mathcal{A}D_{1})^{\top}\tilde{b}_{k}+2\sigma_{2}D_{1}^{\top}\widetilde{v}_{k},$
we can obtain
${\rm svec}(W)=M^{-1}(q_{k}-\tilde{d}_{k}).$ (36)
Substituting (36) into (35), it holds
$~{}~{}~{}~{}\theta(X)$
$=\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle$
$~{}~{}~{}+\left\langle\tilde{d}_{k},M^{-1}(q_{k}-\tilde{d}_{k})\right\rangle$
$~{}~{}~{}+\sigma_{1}\left\|\mathcal{A}D_{1}M^{-1}(q_{k}-\tilde{d}_{k})+\tilde{b_{k}}\right\|^{2}$
$~{}~{}~{}+\sigma_{2}\left\|D_{1}M^{-1}(q_{k}-\tilde{d}_{k})-\widetilde{v}_{k}\right\|^{2}$
$=\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle$
$~{}~{}~{}+\tilde{d}_{k}^{\top}\left(\sigma_{1}(\mathcal{A}D_{1}M^{-1})^{\top}(\mathcal{A}D_{1}M^{-1})+\sigma_{2}M^{-2}\right.$
$~{}~{}~{}\left.-M^{-1}\right)\tilde{d}_{k}+\left\langle\tilde{d}_{k},M^{-1}q_{k}+2\sigma_{1}(\mathcal{A}D_{1}M^{-1})^{\top}\xi_{k}^{1}\right.$
$~{}~{}~{}\left.+2\sigma_{2}(D_{1}M^{-1})^{\top}\xi_{k}^{2}\right\rangle+\|\xi_{k}^{1}\|^{2}+\|\xi_{k}^{2}\|^{2},$
where
$\xi_{k}^{1}=-\mathcal{A}D_{1}M^{-1}q_{k}-\tilde{b}_{k},~{}\xi_{k}^{2}=\widetilde{v}_{k}-D_{1}M^{-1}q_{k}.$
Denoting
$\Omega_{k}$ $=M^{-1}q_{k}+2\sigma_{1}(\mathcal{A}D_{1}M^{-1})^{\top}\xi_{k}^{1}+2\sigma_{2}(D_{1}M^{-1})^{\top}\xi_{k}^{2},$
$\Psi$ $=\sigma_{1}(\mathcal{A}D_{1}M^{-1})^{\top}(\mathcal{A}D_{1}M^{-1})+\sigma_{2}M^{-2}-M^{-1},$
the formulation of the dual problem is presented in the subsequent manner:
$\max_{X=(X_{0},\dots,X_{M})}{}$ $\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle$ (37)
$+\left\langle\tilde{d}_{k},\Omega_{k}\right\rangle+\tilde{d}_{k}^{\top}\Psi\tilde{d}_{k}$
${\rm s.t.}{}$ $X_{0}\in\widetilde{\Gamma}_{+}^{p},$
$X_{i}\in\widetilde{\Gamma}_{+}^{n},~{}i=1,\dots,M.$
In the following text, the sGS (symmetric Gauss-Seidel) semi-proximal ALM (Augmented Lagrangian Method) is presented to solve dual problem (37). Define
$~{}~{}~{}~{}\Theta(X_{0},X_{1},\dots,X_{M})=$
$-\langle X_{1}+\dots+X_{M},T_{2}(V\otimes V){\rm vec}(Q)\rangle$
$-\left\langle\mathcal{L}_{k}(X_{0},\dots,X_{M}),\Omega_{k}\right\rangle$
$+\left\|\sqrt{-\Psi}\mathcal{L}_{k}(X_{0},\dots,X_{M})\right\|^{2}$
$+\delta_{\widetilde{\Gamma}_{+}^{p}}(X_{0})+\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{1})+\dots+\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{M}),$
and dual problem (37) can be equivalently expressed as
$\min_{X=(X_{0}\dots,X_{M})}~{}\Theta(X_{0},X_{1},\dots,X_{M}).$ (38)
For $i=1,\dots,M$, we can compute that
$~{}~{}~{}\partial_{X_{i}}\Theta(X_{0},X_{1},\dots,X_{M})=$
$-T_{2}(V\otimes V){\rm vec}(Q)-T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}\Omega_{k}$
$+2T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}(-\Psi)$
$~{}~{}~{}~{}\mathcal{L}_{k}(X_{0},\dots,X_{i-1},X_{i+1},\dots,X_{M})$
$+2T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}(-\Psi)$
$~{}~{}~{}~{}D_{1}^{\top}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]^{\top}T_{2}^{\top}X_{i}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{i}).$
Define
$~{}~{}~{}H_{i}(X_{0},\dots,X_{i-1},X_{i+1},\dots,X_{M})=$
$T_{2}(V\otimes V){\rm vec}(Q)+T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}\Omega_{k}$
$-2T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}(-\Psi)$
$~{}~{}~{}\mathcal{L}_{k}(X_{0},\dots,X_{i-1},X_{i+1},\dots,X_{M}),$
$L_{i}=2T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]D_{1}(-\Psi)D_{1}^{\top}$
$~{}~{}~{}~{}~{}~{}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]^{\top}T_{2}^{\top},$
then we have
$~{}~{}~{}\partial_{X_{i}}\Theta(X_{0},X_{1},\dots,X_{M})$
$=-H_{i}(X_{0},\dots,X_{i-1},X_{i+1},\dots,X_{M})$
$~{}~{}~{}+L_{i}X_{i}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{i}).$
The subproblem in terms of the variable $X_{i}$ in the backward sGS sweep is given by
$\bar{X}_{i}^{k+1}=\underset{X_{i}}{\operatorname*{argmin}}{}$ $\Theta(X_{0}^{k},X_{1}^{k},\dots,X_{i-1}^{k},X_{i},\bar{X}_{i+1}^{k+1},\dots,\bar{X}_{M}^{k+1})$ (39)
$+\frac{1}{2}\left\|X_{i}-X_{i}^{k}\right\|_{\mathcal{S}_{i}}^{2},$
where $\mathcal{S}_{i}$ is a positive linear operator given by
$\mathcal{S}_{i}(X)=\rho_{i}\mathcal{I}(X)-L_{i}X,~{}\rho_{i}=\max\{{\rm eig}(L_{i})\}$
with an identity operator $\mathcal{I}$.
The optimality condition implies that
00 0 $\in-H_{i}(X_{0}^{k},X_{1}^{k},\dots,X_{i-1}^{k},\bar{X}_{i+1}^{k+1},\dots,\bar{X}_{M}^{k+1})$
$~{}~{}~{}+L_{i}X_{i}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{i})+\mathcal{S}_{i}(X_{i}-X_{i}^{k})$
$=\rho_{i}X_{i}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{i})$
$~{}~{}~{}-H_{i}(X_{0}^{k},X_{1}^{k},\dots,X_{i-1}^{k},\bar{X}_{i+1}^{k+1},\dots,\bar{X}_{M}^{k+1})$
$~{}~{}~{}-\rho_{i}X_{i}^{k}+L_{i}X_{i}^{k}.$
Denoting
$\Delta_{i}^{k}=$ $\rho_{i}^{-1}(H_{i}(X_{0}^{k},X_{1}^{k},\dots,X_{i-1}^{k},\bar{X}_{i+1}^{k+1},\dots,\bar{X}_{M}^{k+1})$
$+\rho_{i}X_{i}^{k}-L_{i}X_{i}^{k}),$
upon careful observation, it may be noted that
$\bar{X}_{i}^{k+1}=(I+\rho_{i}^{-1}\partial\delta_{\widetilde{\Gamma}_{+}^{n}})^{-1}(\Delta_{i}^{k})=\Pi_{\widetilde{\Gamma}_{+}^{n}}(\Delta_{i}^{k}),$ (40)
where the second equality is justified by the lemma provided.
Lemma 1 ([40]).
The projection operator $\Pi_{\mathcal{C}}(\cdot)$ with respect to the convex cone $\mathcal{C}$ can be expressed as
$\Pi_{\mathcal{C}}=(I+\alpha\partial\delta_{\mathcal{C}})^{-1},$
where ($\partial$($\cdot$)∂ ( ⋅ ) denotes the subdifferential operator, and $\alpha\in\mathbb{R}$ can be an arbitrary number.
Remark 11.
Due to the proximal term $\left\|X_{i}-X_{i}^{k}\right\|_{\mathcal{S}_{i}}^{2}$, optimization problem (39) admits a closed-form solution (40). In fact, if the proximal term is removed, i.e., the update of $\bar{X}_{i}^{k+1}$ is given by
$\bar{X}_{i}^{k+1}=\underset{X_{i}}{\operatorname*{argmin}}~{}\Theta(X_{0}^{k},\dots,X_{i},\dots,\bar{X}_{M}^{k+1}),$ (41)
then Algorithm 2 given below still converges to the global minimizer of (38). However, optimization problem (41) has no closed-form solution, necessitating the development of an iterative algorithm for solving problem (41), e.g., gradient descent (GD) or Nesterov GD. It is worth to highlight that the extra iterative algorithm for problem (41) will significantly improve the complexity of Algorithm 2, and hence the introduction of special proximal term $\left\|X_{i}-X_{i}^{k}\right\|_{\mathcal{S}_{i}}^{2}$ is necessary.
For $i=0$, we can compute that
$\partial_{X_{0}}\Theta(X_{0},\dots,X_{M})=L_{0}X_{0}+\partial\delta_{\widetilde{\Gamma}_{+}^{p}}(X_{0})-H_{0}(X_{1},\dots,X_{M}),$
with
$L_{0}=-2\Psi,~{}H_{0}(X_{1},\dots,X_{M})=-2\Psi\mathcal{L}_{k}(X_{1},\dots,X_{M})-\Omega_{k}.$
Similarly, the subproblem in terms of the variable $X_{0}$ in the backward sGS sweep is given by
$X_{0}^{k+1}=\underset{X_{0}}{\operatorname*{argmin}}~{}\Theta(X_{0},\bar{X}_{1}^{k+1},\dots,\bar{X}_{M}^{k+1})+\frac{1}{2}\left\|X_{0}-X_{0}^{k}\right\|_{\mathcal{S}_{0}}^{2},$
where the positive linear operator $\mathcal{S}_{0}$ is given by
$\mathcal{S}_{0}(X)=\rho_{0}\mathcal{I}(X)-L_{0}X,~{}\rho_{0}=\max\{{\rm eig}(L_{0})\}.$
Then, by the same token, we have
$X_{0}^{k+1}=(I+\rho_{0}^{-1}\partial\delta_{\widetilde{\Gamma}_{+}^{p}})^{-1}(\Delta_{0}^{k})=\Pi_{\widetilde{\Gamma}_{+}^{p}}(\Delta_{0}^{k})$ (42)
with
$\Delta_{0}^{k}=\rho_{0}^{-1}(H_{0}(\bar{X}_{1}^{k+1},\dots,\bar{X}_{M}^{k+1})+\rho_{0}X_{0}^{k}-L_{0}X_{0}^{k}).$
For forward sGS sweep, for all $i=1,\dots,M$, $X_{i}^{k+1}$ is given by
$X_{i}^{k+1}=\Pi_{\widetilde{\Gamma}_{+}^{n}}(\tilde{\Delta}_{i}^{k}),$ (43)
where $\tilde{\Delta}_{i}^{k}$ is defined as
$\tilde{\Delta}_{i}^{k}=$ $\alpha_{i}^{-1}(H_{i}(X_{0}^{k+1},X_{1}^{k+1},\dots,X_{i-1}^{k+1},\bar{X}_{i+1}^{k+1},\dots,\bar{X}_{M}^{k+1})$
$+\rho_{i}\bar{X}_{i}^{k+1}-L_{i}\bar{X}_{i}^{k+1}).$
Remark 12.
The update of $X_{i}^{k+1}$ requires the projection onto the convex cone $\widetilde{\Gamma}_{+}^{n}$ for $i=1,\dots,M$. Obviously,
$\bar{X}_{i}^{k+1}=\Pi_{\widetilde{\Gamma}_{+}^{n}}(\Delta_{i}^{k})=\Pi_{\mathbb{S}_{+}^{n}}({\rm svec}^{-1}(\Delta_{i}^{k})),$
and obtaining an explicit solution appears to be challenging. However, the subsequent lemma provides a computationally efficient method for calculating the aforementioned projection.
Lemma 2.
Let $X=\sum_{i=1}^{n}\lambda_{i}v_{i}v_{i}^{\top}\in\mathbb{S}^{n}$ be the eigenvalue decomposition of matrix $X$ with the eigenvalues satisfying $\lambda_{1}\geq\dots\geq\lambda_{n}$, where $v_{i}$ denotes the eigenvector corresponding to the $i$-th eigenvalue. Then, the projection onto the positive semi-definite cone of matrix X can be expressed by
$\Pi_{\mathbb{S}_{+}^{n}}(X)=\sum_{i=1}^{n}\max\{\lambda_{i},0\}v_{i}v_{i}^{\top}.$
Introducing the proximal term for solving separable optimization problem is first proposed in [51], and the convergence analysis can be considered by the analogous methodology of [52], which originally studies the proximal terms in variational inequalities. Remarkably, references [51, 52] only show the fact that introducing the proximal terms has no influence on the algorithm convergence, while how to design the suitable proximal terms remains unstudied, necessitating a case-by-case analysis of individual optimization problem. Hence, the proximal term $\|X_{i}-X_{i}^{k}\|_{\mathcal{S}_{i}}^{2}$ in this paper does not exhibit parallel generalizations of those of [51, 52], but requires additional efforts with consideration of the special structure of dual function $\Theta$; see Remark 11 for details. Additionally, a full understanding of the proximal term can be achieved by examining it from the perspective of auxiliary functions and BSUM framework [53].
Definition 2.
For a continuous function $f:\mathcal{X}\to\mathbb{R}$, where $\mathcal{X}\subseteq\mathbb{R}^{m}$ is the Cartesian product of $n$ closed convex sets: $\mathcal{X}=\mathcal{X}_{1}\times\dots\times\mathcal{X}_{n}$ with $\mathcal{X}_{i}\in\mathbb{R}^{m_{i}}$ and $\sum_{i}m_{i}=m$. We call $\{u_{i}(x_{i},x)\}_{i=1}^{n}$ are auxiliary functions of $f$, if
  1. 1.
    $u_{i}(y_{i},y)=f(y)~{}~{}\forall y\in\mathcal{X},\forall i$,
  2. 2.
    $u_{i}(x_{i},y)\geq f(y_{1},\dots,y_{i-1},x_{i},y_{i+1},\dots,y_{n})$ $\forall x_{i}\in\mathcal{X}_{i}$, $\forall y\in\mathcal{X},\forall i,$
  3. 3.
    $u_{i}^{\prime}(x_{i},y;d_{i})|_{x_{i}=y_{i}}=f^{\prime}(y;d)$ $\forall d=(0,\dots,d_{i},\dots,0)$, s.t. $y_{i}+d_{i}\in\mathcal{X}_{i}$ $\forall i$,
  4. 4.
    $u_{i}(x_{i},y)$ is continuous in $(x_{i},y)$ $\forall i$.
Remark 13.
The following separable (possibly nonconvex) problem is considered in [53]:
$\min~{}f(x),~{}~{}{\rm s.t.~{}}x\in\mathcal{X},$ (44)
where $x=(x_{1},\dots,x_{n})^{\top}\in\mathbb{R}^{m}$, and $\mathcal{X}\subseteq\mathbb{R}^{m}$ is the Cartesian product of $n$ closed convex sets: $\mathcal{X}=\mathcal{X}_{1}\times\dots\times\mathcal{X}_{n}$ with $\mathcal{X}_{i}\in\mathbb{R}^{m_{i}}$ and $\sum_{i}m_{i}=m$. The following pseudocode of BSUM framework proposed in [53] gives methodology for solving separable optimization problem (44), where $\{u_{i}\}_{i=1}^{n}$ is a sequence of auxiliary functions of $f$.
Algorithm 1 BSUM
  Given : Feasible initial point $x^{0}\in\mathcal{X}$
  Result: $x^{r}$
  Set $r=0$
  for $k=0,1,2,...$ do
     Let $r=r+1,i=(r\mod n)+1$
     Let $\mathcal{X}^{r}=\underset{x_{i}\in\mathcal{X}_{i}}{\operatorname*{argmin}}~{}u_{i}(x_{i},x^{r-1})$
     Set $x^{r}_{i}$ to be an arbitrary element in $\mathcal{X}^{r}$
     Set $x_{k}^{r}=x_{k}^{r-1}~{}\forall k\neq i$
     if Stopping Criterion==absent$=$ =True then
        return  $x^{r}$
     end if
  end for
We use the notations
$g(X_{0},\dots,X_{M})=\left\|\sqrt{-\Psi}\mathcal{L}_{k}(X_{0},\dots,X_{M})\right\|^{2},$
$h_{0}(X_{0})=\langle X_{0},\Omega_{k}\rangle,$
$h_{i}(X_{i})=-\left\langle X_{i},T_{2}(V^{\top}\otimes V){\rm vec}(Q)\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.+T_{2}\left[(V\otimes VF_{i})+(VF_{i}\otimes V)\right]T_{1}^{\top}\Omega_{k}\right\rangle,$
for $i=1,\dots,M$. Then, optimization problem (38) is equivalent to the following standard form:
$\min\limits_{X_{0},\dots,X_{M}}$ $\widehat{\Theta}(X_{0},\dots,X_{M})=g(X_{0},\dots,X_{M})+\sum_{i=0}^{M}h_{i}(X_{i})$
${\rm s.t.}{}{}{}$ $X_{0}\times X_{1}\dots\times X_{M}\in\widetilde{\Gamma}_{+}^{p}\times\widetilde{\Gamma}_{+}^{n}\times\dots\times\widetilde{\Gamma}_{+}^{n}.$
Hence, ignoring the forward sGS sweep (i.e., regarding Algorithm 2 as BSUM algorithm with Gauss-Seidel (G-S) rule), let
$u_{i}(X_{i},X^{k})=g(X_{0}^{k},\dots,X_{i},\dots,X_{M}^{k})+\frac{1}{2}\|X_{i}-X_{i}^{k}\|_{\mathcal{S}_{i}}^{2},$
with $X^{k}=(X_{0}^{k},\dots,X_{M}^{k})$, for $i=0,\dots,M$, which leads us to derive the following corollary.
Definition 3.
Consider the following optimization problem
$\min{}{}$ $f(x)=g(x_{1},\dots,x_{K})+\sum_{i=1}^{K}h_{i}(x_{i})$
${\rm s.t.}{}{}$ $x_{i}\in\mathcal{X}_{i},~{}i=1,\dots,K,$
where $g$is a smooth convex function and $h_{i}$ is a closed convex function. We call that $\{u_{i}\}$ is a sequence of strict auxiliary functions, if
  1. 1.
    $u_{i}(y_{i},y)=g(y)~{}~{}\forall y\in\mathcal{X}=\mathcal{X}_{1}\times\dots\times\mathcal{X}_{K},\forall i$,
  2. 2.
    $u_{i}(x_{i},y)\geq g(y_{1},\dots,y_{i-1},x_{i},y_{i+1},\dots,y_{n})~{}~{}\forall x_{i}\in\mathcal{X}_{i},\forall y\in\mathcal{X},\forall i,$
  3. 3.
    $\nabla u_{i}(y_{i},y)=\nabla_{i}g(y),\forall y\in\mathcal{X},\forall i$,
  4. 4.
    $u_{i}(x_{i},y)$ is continuous in $(x_{i},y),\forall i$. Further, for any given $y\in\mathcal{X}$, it is proper, closed and strongly convex function of $x_{i}$, satisfying
    $u_{i}(x_{i},y)\geq$ $u_{i}(\hat{x}_{i},y)+\left\langle\nabla u_{i}(\hat{x}_{i},y),x_{i}-\hat{x}_{i}\right\rangle$
    $+\frac{\gamma_{i}}{2}\|x_{i}-\hat{x}_{i}\|^{2},\forall x_{i},\hat{x}_{i}\in\mathcal{X}_{i},$
    where $\gamma_{i}>0$ is independent of the choice of $y$.
  5. 5.
    For any given $y\in\mathcal{X}$, $u_{k}(x_{i},y)$ has Lipschitz continuous gradient, i.e., $\forall\hat{x}_{i},x_{i}\in\mathcal{X}_{i},\forall i,$
    $\|\nabla u_{i}(x_{i},y)-\nabla u_{i}(\hat{x}_{i},y)\|\leq L_{i}\|x_{k}-\hat{x}_{i}\|,$
    where $L_{i}>0$ is some constant. Further, we have $\forall i,\forall y,z\in\mathcal{X}$
    $\|\nabla u_{i}(x_{i},y)-\nabla u_{i}(x_{i},z)\|\leq G_{i}\|y-z\|,~{}\forall x_{i}\in\mathcal{X}_{i}.$
    Define $L_{{\rm max}}=\max_{i}L_{i},G_{{\rm max}}=\max_{i}G_{i}$.
Corollary 1.
Let $\{X^{k}\}$ be the sequence generated by the BSUM algorithm with G-S rule for solving optimization problem (38). Then it holds
$\widehat{\Theta}(X^{k})-\widehat{\Theta}^{*}\leq\frac{c_{1}}{\sigma_{1}}\frac{1}{k},~{}\forall k\geq 1,$
where $\gamma=\frac{1}{2}\min_{k}\gamma_{k}$,
$R=\max\limits_{X\in\mathcal{X}}\max\limits_{X^{*}\in\mathcal{X}^{*}}\left\{\|X-X^{*}\|\colon\widehat{\Theta}(X)\leq\widehat{\Theta}(X^{0})\right\},$
$c_{1}=\max\{4\sigma_{1}-2,\widehat{\Theta}(X^{0})-\widehat{\Theta}^{*},2\},$
$\sigma_{1}=\frac{\gamma}{(M+1)G_{\rm max}^{2}R^{2}}.$
[…]
In Algorithm 2 below, the sGS rule is employed instead of the G-S rule in order to enhance the algorithm’s performance using empirical means. Review the optimality condition of (38):
00 0 $\in\partial_{X_{0}}\Theta(X_{0},\dots,X_{M})=-H_{0}+L_{0}X_{0}+\partial\delta_{\widetilde{\Gamma}_{+}^{p}}(X_{0}),$
00 0 $\in\partial_{X_{1}}\Theta(X_{0},\dots,X_{M})=-H_{1}+L_{1}X_{1}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{1}),$
$\vdots$
00 0 $\in\partial_{X_{M}}\Theta(X_{0},\dots,X_{M})=-H_{M}+L_{M}X_{M}+\partial\delta_{\widetilde{\Gamma}_{+}^{n}}(X_{M}),$
which are equivalent to
$X_{0}-\Pi_{\widetilde{\Gamma}_{+}^{p}}\left(X_{0}-L_{0}X_{0}+H_{0}\right)=0,$
$X_{1}-\Pi_{\widetilde{\Gamma}_{+}^{n}}\left(X_{1}-L_{1}X_{1}+H_{1}\right)=0,$
[…]
$X_{M}-\Pi_{\widetilde{\Gamma}_{+}^{n}}\left(X_{M}-L_{M}X_{M}+H_{M}\right)=0.$
Then, for $i=1,\dots,M$, the relative residual error can be defined as
${\rm err}_{X_{0}}^{k}$ $=\frac{\left\|X_{0}^{k}-\Pi_{\widetilde{\Gamma}_{+}^{p}}\left(X_{0}^{k}-L_{0}X_{0}^{k}+H_{0}\right)\right\|}{1+\|X_{0}^{k}\|+\|L_{0}X_{0}^{k}-H_{0}\|},$ (45)
${\rm err}_{X_{i}}^{k}$ $=\frac{\left\|X_{i}^{k}-\Pi_{\widetilde{\Gamma}_{+}^{n}}\left(X_{i}^{k}-L_{i}X_{i}^{k}+H_{i}\right)\right\|}{1+\|X_{i}^{k}\|+\|L_{i}X_{i}^{k}-H_{i}\|}.$
Let
${\rm err}^{k}=\max\{{\rm err}_{X_{0}}^{k},{\rm err}_{X_{1}}^{k},\dots,{\rm err}_{X_{M}}^{k}\},$ (46)
and the inner layer optimization process will terminate if ${\rm err}^{k}<\epsilon$. By (36),
${\rm svec}(W)=M^{-1}(q_{k}-\mathcal{L}_{k}(X_{0}^{k},X_{1}^{k},\dots,X_{M}^{k}))$ (47)
is the optimal solution to (32).
It is noteworthy that, for $i=0,\dots,M$, the auxiliary function $u_{i}$ is nothing but the nonseparable term $g$ of dual function $\Theta$ plus the special proximal term $\|X_{i}-X_{i}^{k}\|_{\mathcal{S}_{i}}^{2}$ designed in this paper. Hence, in fact, the design of proximal terms is equivalent to the design of the auxiliary functions of BSUM framework. Significantly, we have so far designed the special sequence of auxiliary functions $\{u_{i}\}_{i=0}^{M}$ such that problem (39) admits a closed-form solution, and thus decreased the algorithmic complexity of Inner Layer Solver; see Remark 11 and Algorithm 2 below.
Algorithm 2 Coordinate Descent with Proximal Term (Inner Layer Solver)
  Given : Initial point $(X_{0}^{0},X_{1}^{0},\dots,X_{M}^{0})$, stopping criterion parameter $\epsilon$
  Result : ${\rm vec}(W)$
  for $k=0,1,2,...$ do
     Update $\bar{X}_{M}^{k+1},\dots,\bar{X}_{1}^{k+1},X_{0}^{k+1}$ by (40), (42)
     Update $X_{1}^{k+1},\dots,X_{M}^{k+1}$ by (43)
     Compute ${\rm err}_{X_{0}}^{k+1},{\rm err}_{X_{1}}^{k+1},\dots,{\rm err}_{X_{M}}^{k+1}$ by (45)
     Determine ${\rm err}^{k+1}$ by (46)
     if ${\rm err}^{k+1}<\epsilon$ then
        Compute ${\rm svec}(W)$ by (47)
        return  ${\rm vec}(W)=D_{1}{\rm svec}(W)$
     end if
  end for
5 Discussion of Convergence Rate and Acceleration
In this section, the convergence of Algorithm 3 (using iteration scheme (21) for solving problem (15)) is studied. Additionally, the optimization problem induced by quadratic sparsity promoting function is of concern, which is shown to exhibit an accelerated convergence rate.
Algorithm 3 ALM based Decentralized LQ (Outer Layer Solver)
  Given : Initial point $\widetilde{W}_{0}\in\Gamma^{p},v_{0}\in\Gamma^{p},\widetilde{P}_{0}\in\mathbb{R}^{mn},w_{0}\in\mathbb{R}^{mn},\lambda_{0}\in\mathbb{R}^{mn},\theta_{0}=1,\beta_{0}>0,\gamma_{0}>0$, smoothness parameter $L_{f}$
  Result : $\widetilde{W}_{k}$, $\widetilde{P}_{k}$
  for $k=0,1,2,...$ do
     Compute $\alpha_{k}$ by (49)
     Compute $\eta_{g,k},\widetilde{y}_{k},\widetilde{\eta}_{f,k},\widetilde{v}_{k}$ by (22), (23), (24), (25)
     Determine $u_{k}=(1+\alpha_{k})^{-1}(\widetilde{W}_{k}+\alpha_{k}v_{k}),~{}d_{k}=\nabla f_{1}(u_{k})+\mathcal{A}^{\top}\lambda_{k}$
     Update $v_{k+1}\leftarrow$Inner Layer Solver
     Update $\widetilde{W}_{k+1}=(1+\alpha_{k})^{-1}(\widetilde{W}_{k}+\alpha_{k}v_{k+1})$
     Update $\bar{\lambda}_{k+1}=\lambda_{k}+\alpha_{k}/\theta_{k}(\mathcal{A}v_{k+1}+\mathcal{B}w_{k})$
     Update $\widetilde{P}_{k+1}$ by (31)
     Update $w_{k+1}=\widetilde{P}_{k+1}+\alpha_{k}^{-1}(\widetilde{P}_{k+1}-\widetilde{P}_{k})$
     Update $\lambda_{k+1}=\lambda_{k}+\alpha_{k}/\theta_{k}(\mathcal{A}v_{k+1}+\mathcal{B}w_{k+1})$
     Update $\theta_{k+1},\gamma_{k+1},\beta_{k+1}$ by (26)
     if Stopping Criterion==absent$=$ =True then
        return  $\widetilde{W}_{k+1},\widetilde{P}_{k+1}$
     end if
  end for
Remark 14.
Denoting
$r_{k+1}=\mathcal{A}\widetilde{W}_{k+1}+\mathcal{B}\widetilde{P}_{k+1},$
$s_{k+1}=\rho\mathcal{A}^{\top}\mathcal{B}(\widetilde{P}_{k+1}-\widetilde{P}_{k})$
with constant $\rho>0$, it is recommended that the stopping criterion of Algorithm 3 can be selected as
$\|r_{k}\|\leq\epsilon^{(k)}_{\rm pri},~{}\|s_{k}\|\leq\epsilon^{(k)}_{\rm dua},$
where
$\epsilon^{(k)}_{\rm pri}=\sqrt{M+mn}\epsilon_{1}+\epsilon_{2}\max\{\|\mathcal{A}\widetilde{W}_{k}\|,\|\mathcal{B}\widetilde{P}_{k}\|\},$
$\epsilon^{(k)}_{\rm dua}=p\epsilon_{1}+\epsilon_{2}\|\mathcal{A}^{\top}\lambda_{k}\|$
with some constants $\epsilon_{1},\epsilon_{2}$ [54]. Moreover, the stopping criterion can be selected in many other ways; see [55] for some examples.
Algorithm 3 is the outer layer solver of optimization problem (19). Let the Lagrange function associated to problem (19) be
$\mathcal{L}(\widetilde{W},\widetilde{P},\lambda)=f(\widetilde{W})+h(\widetilde{P})+\left\langle\lambda,\mathcal{A}\widetilde{W}+\mathcal{B}\widetilde{P}\right\rangle,$ (48)
and let the discrete Lyapunov function be
$\mathcal{E}_{k}=$ $\mathcal{L}(\widetilde{W}_{k},\widetilde{P}_{k},\lambda^{*})-\mathcal{L}(\widetilde{W}^{*},\widetilde{P}^{*},\lambda_{k})+\frac{\gamma_{k}}{2}\|v_{k}-\widetilde{W}^{*}\|^{2}$
$+\frac{\beta_{k}}{2}\|w_{k}-\widetilde{P}^{*}\|^{2}+\frac{\theta_{k}}{2}\|\lambda_{k}-\lambda^{*}\|^{2},$
where $(\widetilde{W}^{*},\widetilde{P}^{*},\lambda^{*})$ is a saddle point of $\mathcal{L}$.
$\mathcal{L}(\widetilde{W}^{*},\widetilde{P}^{*},\lambda)\leq\mathcal{L}(\widetilde{W}^{*},\widetilde{P}^{*},\lambda^{*})\leq\mathcal{L}(\widetilde{W},\widetilde{P},\lambda^{*}),~{}\forall(\widetilde{W},\widetilde{P},\lambda).$
The convergence analysis of Algorithm 3 can be derived, as stated in the following corollary.
Corollary 2.
Under the initial condition $\gamma_{0}>0$, $\beta_{0}>0$ and the condition
$\|\mathcal{B}\|^{2}\alpha_{k}^{2}=\beta_{k}\theta_{k},$ (49)
it holds that $\{\widetilde{W}_{k},\widetilde{P}_{k}\}\subset\mathbb{R}^{p^{2}}\times\mathbb{R}^{mn}$ and
$\left\|\mathcal{A}\widetilde{W}_{k}+\mathcal{B}\widetilde{P}_{k}\right\|\leq\theta_{k}\mathcal{R}_{0},$
$\left|F(\widetilde{W}_{k},\widetilde{P}_{k})-F^{*}\right|\leq\theta_{k}\left(\mathcal{E}_{0}+\left\|\lambda^{*}\right\|\mathcal{R}_{0}\right),$
with $F(\widetilde{W},\widetilde{P})=f(\widetilde{W})+h(\widetilde{P})$. Above, $\mathcal{R}_{0}=\sqrt{2\mathcal{E}_{0}}+\|\lambda_{0}-\lambda^{*}\|+\|\mathcal{A}\widetilde{W}_{0}+\mathcal{B}\widetilde{P}_{0}\|$ and $\theta_{k}$ satisfies
$\theta_{k}\lesssim\frac{\|\mathcal{B}\|}{\sqrt{\beta_{0}}k},$
provided that $\beta_{0}\leq\left\|\mathcal{B}\right\|^{2}$.
[…]
The following proposition elucidates the asymptotic behavior of optimization problem (15) when the parameter $\gamma_{n}\to 0$. Therefore, it is possible to achieve a near-optimal centralized controller by selecting a sufficiently small value for $\gamma$ and utilizing Algorithm 3.
Proposition 2.
Let $\gamma_{n}\to 0$. Assume that $C^{\top}C\succ 0$, and denote
$F^{\gamma_{n}}(\widetilde{W})$ $=f(\widetilde{W})+\gamma_{n}g((V_{2}\otimes V_{1})\widetilde{W})+\delta_{\widehat{\mathcal{A}}\widetilde{W}=0}(\widetilde{W}),$
$F(\widetilde{W})$ $=f(\widetilde{W})+\delta_{\widehat{\mathcal{A}}\widetilde{W}=0}(\widetilde{W}).$
Then it holds $\inf F^{\gamma_{n}}\to\inf F.$
Proof: Obviously, $F^{\gamma_{n}}(\widetilde{W})\stackrel{{\scriptstyle\rm e}}{{\longrightarrow}}F(\widetilde{W})$ (the epigraph convergence; see Definition 7.1 of [56]). Based on the fact
$\langle{\rm vec}(R),\widetilde{W}\rangle=\langle R,W\rangle\geq\lambda_{{\rm min}}(R){\rm tr}(W)\geq\lambda_{{\rm min}}(R)\|W\|_{2},$
[…]
Above, our choice of sparsity promoting function is the weighted-$\ell_{1}$-norm, i.e., $g(P)=\sum_{i,j}w_{ij}|P_{ij}|$, which is convex but not strongly convex. Hence, regretfully, algorithm (21) exhibits $\mathcal{O}\left(\frac{1}{k}\right)$ convergence rate. Alternatively, designing a strongly convex sparsity promoting function is one approach to solving the issue described above. For instance, let $g_{2}:\mathbb{R}\to\mathbb{R}$ be
$g_{2}(x)=\left\{\begin{aligned} &\frac{1}{2}a_{1}x^{2}+b_{1}x,~{}{\rm if}~{}x\leq 0,\\ &\frac{1}{2}a_{2}x^{2}+b_{2}x,~{}{\rm if}~{}x>0,\end{aligned}\right.$ (50)
and $g_{\mathcal{Q}}(P)=\sum_{ij}w_{ij}g_{2}(p_{ij})$ called a weighted piecewise quadratic function. The proof of the following lemma is straightforward.
Lemma 3.
$g_{\mathcal{Q}}(P)$ is sparsity promoting if and only if
$a_{1}\geq 0,~{}a_{2}\geq 0,~{}b_{1}\leq 0\leq b_{2},~{}{\rm and}~{}b_{2}-b_{1}>0.$
Obviously, by choosing $a_{1},a_{2}>0,b_{1}<0<b_{2}$, $g_{\mathcal{Q}}(P)$ becomes a strongly convex sparsity promoting function with modulus
$\mu_{g_{\mathcal{Q}}}=\min\limits_{i,j}\left\{w_{ij}\min\{a_{1},a_{2}\}\right\}.$
Hence, there comes a surprising result that designing an algorithm with an accelerated convergence rate is equivalent to designing a strongly convex sparsity promoting function. If we consider the piecewise quadratic function $g_{\mathcal{Q}}(P)$, the primal optimization problem becomes
$\min_{\widetilde{W},\widetilde{P}}{}{}$ $f(\widetilde{W})+\gamma g_{\mathcal{Q}}(\widetilde{P})$
${\rm s.t.}{}{}$ $\mathcal{A}\widetilde{W}+\mathcal{B}\widetilde{P}=0.$
Similarly, by (21), we have
$\widetilde{P}_{k+1}=\mathbf{prox}_{\tau_{k}\gamma g_{\mathcal{Q}}}^{\mathbb{R}^{mn\times 1}}(\widetilde{P}_{k}+\frac{\alpha_{k}\beta_{k}}{\eta_{g,k}}(w_{k}-\widetilde{P}_{k})-\tau_{k}\mathcal{B}^{\top}\bar{\lambda}_{k+1}).$
Remarkably, $\widetilde{P}_{k+1}$ can be obtained by the following theorem.
Theorem 3.
Provided that $a_{1}>0,a_{2}>0,b_{1}<0<b_{2}$, letting
$\hat{P}_{ij;k+1}^{*}=\left\{\begin{aligned} &0,~{}{\rm if}~{}\Omega_{ij}^{k}<\frac{\gamma w_{ij}b_{2}}{\rho_{k}},\Omega_{ij}^{k}\geq 0,\\ &0,~{}{\rm if}~{}\Omega_{ij}^{k}>\frac{\gamma w_{ij}b_{1}}{\rho_{k}},\Omega_{ij}^{k}<0,\\ &\frac{\rho_{k}\Omega_{ij}^{k}-\gamma w_{ij}b_{2}}{\gamma w_{ij}a_{2}+\rho_{k}},~{}{\rm if}~{}\Omega_{ij}^{k}\geq\frac{\gamma w_{ij}b_{2}}{\rho_{k}},\Omega_{ij}^{k}\geq 0,\\ &\frac{\rho_{k}\Omega_{ij}^{k}-\gamma w_{ij}b_{1}}{\gamma w_{ij}a_{1}+\rho_{k}},~{}{\rm if}~{}\Omega_{ij}^{k}\leq\frac{\gamma w_{ij}b_{1}}{\rho_{k}},\Omega_{ij}^{k}<0,\end{aligned}\right.$
where $\Omega^{k}$ is given in (30), we have
$\widetilde{P}_{k+1}=(\hat{P}_{11;k+1}^{*},\hat{P}_{21;k+1}^{*},\dots,\hat{P}_{mn;k+1}^{*})^{\top}.$ (51)
Proof: By the definition of $\widetilde{P}_{k+1}$, we have
$\widetilde{P}_{k+1}$ $=\mathbf{prox}_{\tau_{k}\gamma g_{\mathcal{Q}}}^{\mathbb{R}^{mn\times 1}}(\widetilde{P}_{k}-\tau_{k}\mathcal{B}^{\top}\bar{\lambda}_{k+1})$
$=\left(\underset{P_{ij}\in\mathbb{R}}{\operatorname*{argmin}}\left\{\gamma w_{ij}g_{2}(P_{ij})+\frac{\rho_{k}}{2}\left(P_{ij}-\Omega^{k}_{ij}\right)^{2}\right\}\right)_{mn},$
Noticing the fact that for all $x,y\in\mathbb{R}$ and $\alpha\in[0,1]$
$|\alpha x+(1-\alpha)y|^{2}=\alpha|x|^{2}+(1-\alpha)|y|^{2}-\alpha(1-\alpha)|x-y|^{2},$
it is obvious that optimization problem
$\hat{P}_{ij;k+1}^{*}\in\min_{P_{ij}\in\mathbb{R}}\gamma w_{ij}g_{2}(P_{ij})+\frac{\rho_{k}}{2}\left(x-\Omega^{k}_{ij}\right)^{2}$ (52)
is strongly convex. Thus, $\hat{P}_{ij;k+1}^{*}$ exists and is unique. Denote $\varphi(P_{ij})=\gamma w_{ij}g_{2}(P_{ij})+\frac{\rho_{k}}{2}\left(P_{ij}-\Omega^{k}_{ij}\right)^{2}$. When $\Omega_{ij}^{k}\geq 0$, we can obtain
$\varphi(P_{ij})=\left\{\begin{aligned} &\left(\frac{\gamma w_{ij}a_{1}+\rho_{k}}{2}\right)P_{ij}^{2}+\left(\gamma w_{ij}b_{1}-\rho_{k}\Omega_{ij}^{k}\right)P_{ij}\\ &\qquad\qquad\qquad+\frac{\rho_{k}}{2}{\Omega_{ij}^{k}}^{2},~{}{\rm if}~{}P_{ij}\leq 0,\\ &\left(\frac{\gamma w_{ij}a_{2}+\rho_{k}}{2}\right)P_{ij}^{2}+\left(\gamma w_{ij}b_{2}-\rho_{k}\Omega_{ij}^{k}\right)P_{ij}\\ &\qquad\qquad\qquad+\frac{\rho_{k}}{2}{\Omega_{ij}^{k}}^{2},~{}{\rm if}~{}P_{ij}>0.\end{aligned}\right.$
Hence, $\varphi(P_{ij})$ monotonously decreases on $(-\infty,0]$. Moreover, if $\Omega_{ij}^{k}<\frac{\gamma w_{ij}b_{2}}{\rho_{k}}$, $\varphi(P_{ij})$ will monotonously increase on $[0,+\infty)$, and if $\Omega_{ij}^{k}\geq\frac{\gamma w_{ij}b_{2}}{\rho_{k}}$, $\varphi(P_{ij})$ will monotonously decrease on $\left[0,\frac{\rho_{k}\Omega_{ij}^{k}-\gamma w_{ij}b_{2}}{\gamma w_{ij}a_{2}+\rho_{k}}\right]$, while increase on $\left[\frac{\rho_{k}\Omega_{ij}^{k}-\gamma w_{ij}b_{2}}{\gamma w_{ij}a_{2}+\rho_{k}},+\infty\right)$. Consequently, when $\Omega_{ij}^{k}\geq 0$, we have
$\hat{P}_{ij;k+1}^{*}=\left\{\begin{aligned} &0,~{}~{}~{}~{}{\rm if}~{}\Omega_{ij}^{k}<\frac{\gamma w_{ij}b_{2}}{\rho_{k}},\\ &\frac{\rho_{k}\Omega_{ij}^{k}-\gamma w_{ij}b_{2}}{\gamma w_{ij}a_{2}+\rho_{k}},~{}~{}~{}~{}{\rm if}~{}\Omega_{ij}^{k}\geq\frac{\gamma w_{ij}b_{2}}{\rho_{k}}.\end{aligned}\right.$
When $\Omega_{ij}^{k}<0$, by the same token, we have
$\hat{P}_{ij;k+1}^{*}=\left\{\begin{aligned} &0,~{}~{}~{}~{}{\rm if}~{}\Omega_{ij}^{k}>\frac{\gamma w_{ij}b_{1}}{\rho_{k}},\\ &\frac{\rho_{k}\Omega_{ij}^{k}-\gamma w_{ij}b_{1}}{\gamma w_{ij}a_{1}+\rho_{k}},~{}~{}~{}~{}{\rm if}~{}\Omega_{ij}^{k}\leq\frac{\gamma w_{ij}b_{1}}{\rho_{k}}.\end{aligned}\right.$
[…]
Algorithm 4 Accelerated ALM based Decentralized LQ (Outer Layer Solver)
  Given : Initial point $\widetilde{W}_{0}\in\Gamma^{p},v_{0}\in\Gamma^{p},y_{0}\in\mathbb{R}^{mn},w_{0}\in\mathbb{R}^{mn},\lambda_{k}\in\mathbb{R}^{mn},\theta_{0}=1,\beta_{0}=\mu_{g_{\mathcal{Q}}},\gamma_{0}>0$, smoothness parameter $L_{f}$
  Result : $\widetilde{W}_{k}$, $\widetilde{P}_{k}$
  for $k=0,1,2,...$ do
     Compute $\alpha_{k}$ by (53)
     Compute $\eta_{g,k},\widetilde{y}_{k},\widetilde{\eta}_{f,k},\widetilde{v}_{k}$ by (22), (23), (24), (25)
     Determine $u_{k}=(1+\alpha_{k})^{-1}(\widetilde{W}_{k}+\alpha_{k}v_{k}),~{}d_{k}=\nabla f_{1}(u_{k})+\mathcal{A}^{\top}\lambda_{k}$
     Update $v_{k+1}\leftarrow$Inner Layer Solver
     Update $\widetilde{W}_{k+1}=(1+\alpha_{k})^{-1}(\widetilde{W}_{k}+\alpha_{k}v_{k+1})$
     Update $\bar{\lambda}_{k+1}=\lambda_{k}+\alpha_{k}/\theta_{k}(\mathcal{A}v_{k+1}+\mathcal{B}w_{k})$
     Update $\widetilde{P}_{k+1}$ by (51)
     Update $\lambda_{k+1}=\lambda_{k}+\alpha_{k}/\theta_{k}(\mathcal{A}v_{k+1}+\mathcal{B}w_{k+1})$
     Update $\theta_{k+1},\gamma_{k+1},\beta_{k+1}$ by (26)
     if Stopping Criterion==absent$=$ =True then
        return  $\widetilde{W}_{k+1},\widetilde{P}_{k+1}$
     end if
  end for
Interestingly, the only difference between Algorithm 3 and Algorithm 4 is the update of $\widetilde{P}_{k+1}$. However, the former holds $\mathcal{O}\left(\frac{1}{k}\right)$ convergence rate, while the latter holds $\mathcal{O}\left(\frac{1}{k^{2}}\right)$ convergence rate. Note that Nesterov GD is an extension of vanilla GD by introducing an additional sequence and modifying the fundamental framework; Algorithm 3 and Algorithm 4 have a completely consistent framework. We have established a direct relationship between the acceleration of solving decentralized LQ problems and the design of a strongly convex sparsity promoting function, and to our best knowledge, this phenomenon has not been reported in the field of sparse optimization.
Corollary 3.
Provided that for all $i,j$, $w_{ij}\neq 0$, and $a_{1}>0,a_{2}>0,b_{1}<0<b_{2}$, under the initial condition $\gamma_{0}>0$, $\beta_{0}=\mu_{g_{\mathcal{Q}}}>0$ and the condition
$\|\mathcal{B}\|^{2}\alpha_{k}^{2}=\beta_{k}\theta_{k},$ (53)
it holds $\{\widetilde{W}_{k},\widetilde{P}_{k}\}\subset\mathbb{R}^{p^{2}}\times\mathbb{R}^{mn}$ and
$\left\|\mathcal{A}\widetilde{W}_{k}+\mathcal{B}\widetilde{P}_{k}\right\|\leq\theta_{k}\mathcal{R}_{0},$
$\left|F(\widetilde{W}_{k},\widetilde{P}_{k})-F^{*}\right|\leq\theta_{k}\left(\mathcal{E}_{0}+\left\|\lambda^{*}\right\|\mathcal{R}_{0}\right),$
where $F(\widetilde{W},\widetilde{P})=f(\widetilde{W})+h(\widetilde{P})$. Above, $\mathcal{R}_{0}=\sqrt{2\mathcal{E}_{0}}+\|\lambda_{0}-\lambda^{*}\|+\|\mathcal{A}\widetilde{W}_{0}+\mathcal{B}\widetilde{P}_{0}\|$ and $\theta_{k}$ satisfies
$\theta_{k}\lesssim\min\left\{\frac{\|\mathcal{B}\|}{\sqrt{\beta_{0}}k},\frac{\|\mathcal{B}\|^{2}}{\mu_{g_{\mathcal{Q}}}k^{2}}\right\},$
provided that $\beta_{0}\leq\left\|\mathcal{B}\right\|^{2}$.
[…]
6 Directly Optimizing $\ell_{0}$-penalty
Reviewing the fundamental topic of this paper, we want to penalize the number of communication links among decentralized controllers by introducing the sparsity promoting function into the optimal control problem. In this section, the optimization problem
$\min_{W\in\mathbb{S}^{p}}{}{}$ $\langle R,W\rangle+\gamma\|V_{1}WV_{2}^{\top}\|_{0}$ (54)
${\rm s.t.}{}{}$ $W\in\mathscr{C}$
is directly studied, which is nonconvex and discontinuous. It is worth to highlight that penalizing the number of communication links among decentralized controllers is equivalent to the penalty term $\gamma\|W_{2}\|_{0}$ of problem (54), and the penalty promoting terms $\gamma g(W_{2})$ and $\gamma g_{\mathcal{Q}}(W_{2})$ introduced in the previous sections can be regarded as the convex relaxations of $\gamma\|W_{2}\|_{0}$. By the following discussions, we show that a series of coordinatewise convex problems can serve as an approximation for problem (54). Additionally, the variational properties of problem (54) are of concern. We first introduce some standard definitions below.
Definition 4.
Let $f\colon\mathcal{D}\to\mathbb{R}$ be a function where $\mathcal{D}\subseteq\mathbb{R}^{m}$ is a convex set. The directional derivative of $f$ at point $x$ in direction $d$ is defined by
$f^{\prime}(x;d)\doteq\underset{\lambda\downarrow 0}{\rm liminf}\frac{f(x+\lambda d)-f(x)}{\lambda}.$
Definition 5.
Let $f\colon\mathcal{D}\to\mathbb{R}$ be a function where $\mathcal{D}\subseteq\mathbb{R}^{m}$ is a convex set. The point $x$ is a stationary point of $f(\cdot)$ if $f^{\prime}(x;d)\geq 0$ for all $d$ such that $x+d\in\mathcal{D}$.
Definition 6.
$z\in{\rm dom}f\subseteq\mathbb{R}^{m}$ is the coordinatewise minimizer of $f$ with respective to the coordinates in $\mathbb{R}^{m_{1}},\dots,\mathbb{R}^{m_{n}}$, $m_{1}+\dots+m_{n}=m$ if $\forall k=1,\dots,n$
$f(z+d_{k}^{0})\geq f(z)~{}\forall d_{k}\in\mathbb{R}^{m_{k}}~{}{\rm with}~{}z+d_{k}^{0}\in{\rm dom}f,$
where $d_{k}^{0}=(0,\dots,d_{k},\dots,0)$.
Definition 7.
The function $f\colon\mathbb{R}^{m}\to\mathbb{R}$ is regular at the point $z\in{\rm dom}f$ with respective to the coordinates $m_{1},m_{2},\dots,m_{n}$,$m_{1}+\dots+m_{n}=m$ if $f^{\prime}(z;d)\geq 0$ for all $d=(d_{1},d_{2},\dots,d_{n})$ with $f^{\prime}(z;d_{k}^{0})\geq 0$, where $d_{k}^{0}\doteq(0,\dots,d_{k},\dots,0)$ and $d_{k}\in\mathbb{R}^{m_{k}}$ for all $k$.
Definition 8.
The function is lower semi-continuous and is the greatest of all the lower semi-continuous functions $g$ such that $g\leq f$. It is called the lower closure of $f$, denoted by ${\rm cl}f$.
Definition 9.
For any sequence $\{f^{v}\}_{v\in\mathbb{N}}$ of functions on $\mathbb{R}^{n}$, the lower epi-limit ${\rm e\text{-}liminf}_{v}f^{v}$ is the function having as its epigraph the outer limit of the sequence of sets ${\rm epi}f^{v}$:
${\rm epi}({\rm e\text{-}liminf}_{v}f^{v})\doteq{\rm limsup}_{v}({\rm epi}f^{v}).$
The upper epi-limit ${\rm e\text{-}limsup}_{v}f^{v}$ is the function having as its epigraph the inner limit of the sets ${\rm epi}f^{v}$:
${\rm epi}({\rm e\text{-}limsup}_{v}f^{v})\doteq{\rm liminf}_{v}({\rm epi}f^{v}).$
When these two functions coincide, the epi-limits function ${\rm e\text{-}lim}f^{v}$ is said to exist:
${\rm e\text{-}lim}f^{v}\doteq{\rm e\text{-}liminf}_{v}f^{v}={\rm e\text{-}limsup}_{v}f^{v}.$
The subsequent analysis in this section is based on the following assumption.
Assumption 4.
$C^{\top}C\succ 0$.
By the same token with (18), $\ell_{0}$-penalty induced optimization problem can be equivalently converted into the following form
$\min_{\widetilde{W}}{}{}$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})$ (55)
$+\gamma\|(V_{2}\otimes V_{1})\widetilde{W}\|_{0}$
${\rm s.t.}{}{}$ $\left(V_{i2}^{\top}\otimes V_{i1}\right)\widetilde{W}=0,~{}~{}i=1,\dots,M.$
Recall
$\widehat{\mathcal{A}}=\begin{pmatrix}V_{12}^{\top}\otimes V_{11}\\ \vdots\\ V_{M2}^{\top}\otimes V_{M1}\\ \end{pmatrix},$ (56)
and let $\Omega=\{\widetilde{W}\colon\widehat{\mathcal{A}}\widetilde{W}=0\}$. Then, the optimization problem can be expressed in the unconstrained form, i.e.,
$\min_{\widetilde{W}}{}$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})$ (P)
$+\delta_{\Omega}(\widetilde{W})+\gamma\|(V_{2}\otimes V_{1})\widetilde{W}\|_{0}.$
The discontinuous nature of optimization problem (P) makes it extremely challenging to design a suitable algorithm. The work [57] introduces a class of continuous folded concave penalty $f_{\sigma}:\mathbb{R}\to\mathbb{R}$ with $f_{\sigma}(|y|)$ converging to $I(y)$ as $\sigma\downarrow 0$. Hence, $f_{\sigma}(|y|)$ can be viewed as a class of continuous approximation of $\ell_{0}$-norm. Specifically, Property $\mathcal{D}$ is introduced to define the class of folded concave penalty $f_{\sigma}(|y|)$.
Definition 10 ([57]).
Let $f\colon\mathbb{R}\to\mathbb{R}$ and define $f_{\sigma}(y)\doteq f(y/\sigma)$ for any $\sigma>0$. The function $f$ is said to possess Property $\mathcal{D}$, if
  1. 1.
    $f$ is real analytic on $(y_{0},\infty)$ for some $y_{0}<0$,
  2. 2.
    $\forall y\geq 0$, $f^{\prime\prime}(y)\geq-\mu_{0}$, where $\mu_{0}>0$ is some constant,
  3. 3.
    $f$ is concave on $\mathbb{R}$,
  4. 4.
    $f(y)=0\Leftrightarrow y=0$,
  5. 5.
    $\lim_{y\to\infty}f(y)=1$.
It is obvious that if $f$ possesses Property $\mathcal{D}$, then
$\lim\limits_{\sigma\downarrow 0^{+}}f_{\sigma}(|y|)=I(y)=\left\{\begin{aligned} &0,~{}~{}y=0,\\ &1,~{}~{}\text{otherwise}.\end{aligned}\right.$
In fact, there are a plenty of functions that satisfy Property $\mathcal{D}$, for instance, $f(y)=1-e^{-y}$. For the record, function $f(y)=1-e^{-y}$ is utilized in this paper without special instructions, and in this case, $f_{\sigma}(y)=1-e^{\frac{y}{\sigma}}$. For $x=(x_{1},\dots,x_{n})^{\top}\in\mathbb{R}^{n}$, denote $f_{\sigma}(x)=\sum_{i=1}^{n}f_{\sigma}(x_{i})$.
Based on the fact that $\lim\limits_{\sigma\downarrow 0^{+}}f_{\sigma}(|x|)=\|x\|_{0}$, a class of optimization problems are proposed:
$\min_{\widetilde{W}}{}$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})$ ($\textbf{P}_{\sigma}$)
$+\delta_{\Omega}(\widetilde{W})+\gamma f_{\sigma}\left(\left|(V_{2}\otimes V_{1})\widetilde{W}\right|\right).$
($\textbf{P}_{\sigma}$) is a nonconvex optimization problem since the folded concave penalty term $f_{\sigma}$ is introduced. To avoid optimizing nonconvex penalty term directly, [29] proposes a convex relaxation framework for solving $f_{\sigma}$ penalty induced optimization problem.
Lemma 4.
Let $f(x)$ satisfying Property $\mathcal{D}$, and consider the function $f_{\sigma}(x)(x\geq 0)$. It holds
$f_{\sigma}(x)=\inf\limits_{y\leq 0}\{-yx+g^{*}_{\sigma}(y)\},$
where $g_{\sigma}(x)=-f_{\sigma}(x)$ and $g^{*}_{\sigma}$ is the conjugate function of $g_{\sigma}$.
[…]
Hence, $f_{\sigma}$ penalty can be expressed as
$f_{\sigma}(|x|)=\inf\limits_{y\geq 0}\left\{y^{\top}|x|+g_{\sigma}^{*}(-y)\right\},$
and optimization problem ($\textbf{P}_{\sigma}$) is equivalent to the following problem
$\min_{\widetilde{W},y\in\mathbb{R}^{mn}_{+}}~{}\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\cdots$ ($\textbf{P}^{\prime}_{\sigma}$)
$+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})+\delta_{\Omega}(\widetilde{W})+\gamma y^{\top}\left|(V_{2}\otimes V_{1})\widetilde{W}\right|+\gamma g_{\sigma}^{*}(-y).$
Denote
$r_{1}(\widetilde{W})=\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})+\delta_{\Omega}(\widetilde{W}),$
$r_{2,\sigma}(y)=\gamma g_{\sigma}^{*}(-y),$
$h(\widetilde{W},y)=\gamma y^{\top}\left|(V_{2}\otimes V_{1})\widetilde{W}\right|,$
then problem ($\textbf{P}^{\prime}_{\sigma}$) can be denoted in a compact form:
$\min_{\widetilde{W},y\in\mathbb{R}^{mn}_{+}}~{}H_{\sigma}(\widetilde{W},y)=r_{1}(\widetilde{W})+r_{2,\sigma}(y)+h(\widetilde{W},y).$ ($\textbf{P}^{\prime\prime}_{\sigma}$)
It is important to highlight that $r_{1},r_{2}$ are closed and convex, but $h$ is nonconvex and nonsmooth. It seems that, compared to optimization problem ($\textbf{P}_{\sigma}$), ($\textbf{P}^{\prime\prime}_{\sigma}$) remains a nonconvex problem but introduces auxiliary variable $y$, which even makes optimization problem ($\textbf{P}^{\prime\prime}_{\sigma}$) more challenging. Interestingly, it is evident that $H_{\sigma}(\widetilde{W},y)$ is closed convex w.r.t. $\widetilde{W}$ and $y$, respectively. This observation suggests that by employing block coordinate descent (BCD), ($\textbf{P}^{\prime\prime}_{\sigma}$) can be converted to successive convex optimization problems. Specifically, given $y_{k-1}$, update $\widetilde{W}_{k}$; then given $\widetilde{W}_{k}$, update $y_{k}$, i.e.,
$\widetilde{W}_{k}\in\underset{\widetilde{W}\in\mathbb{R}^{p^{2}}}{\operatorname*{argmin}}~{}r_{1}(\widetilde{W})+h(\widetilde{W},y_{k-1}),$ ($\textbf{BP}\sigma-1$)
$y_{k}\in\underset{y\in\mathbb{R}_{+}^{mn}}{\operatorname*{argmin}}~{}r_{2,\sigma}(y)+h(\widetilde{W}_{k},y).$ ($\textbf{BP}\sigma-2$)
It is important to underline that both ($\textbf{BP}\sigma-1$) and ($\textbf{BP}\sigma-2$) may be classified as convex optimization problems. Our objective is then to solve ($\textbf{BP}\sigma-1$) and ($\textbf{BP}\sigma-2$) alternatively in order to approximate the global minimizers of nonconvex problem ($\textbf{P}^{\prime\prime}_{\sigma}$). Before discussing the relation between ($\textbf{P}^{\prime\prime}_{\sigma}$) and ($\textbf{BP}\sigma-1$) ($\textbf{BP}\sigma-2$), we first propose the relation between ($\textbf{P}^{\prime\prime}_{\sigma}$) and (P). Denote
$H_{\sigma}(\widetilde{W})$ $=r_{1}(\widetilde{W})+\gamma f_{\sigma}\left(\left|(V_{2}\otimes V_{1})\widetilde{W}\right|\right),$
$H(\widetilde{W})$ $=r_{1}(\widetilde{W})+\gamma\|(V_{2}\otimes V_{1})\widetilde{W}\|_{0},$
and the following theorem proves that
[…]
Theorem 4.
Let $\sigma_{i}\downarrow 0$, then the following statements hold.
  1. 1.
    $\inf\limits_{\widetilde{W}\in\mathbb{R}^{p^{2}}}H_{\sigma_{i}}(\widetilde{W})\to\inf\limits_{\widetilde{W}\in\mathbb{R}^{p^{2}}}H(\widetilde{W})$.
  2. 2.
    For $v$ in some index set $N\in\mathcal{N}_{\infty}$, the sets ${\operatorname*{argmin}}H_{\sigma_{v}}$ are nonempty and form a bounded sequence with
    $\limsup_{v}\left({\operatorname*{argmin}}H_{\sigma_{v}}\right)\subset{\operatorname*{argmin}}H.$
  3. 3.
    For any choice of $\epsilon_{i}\downarrow 0$ and $\widetilde{W}_{i}\in\epsilon_{i}\text{-}{\operatorname*{argmin}}H_{\sigma_{i}}$, the sequence $\{\widetilde{W}_{i}\}_{i\in\mathbb{N}}$ is bounded and such that all its cluster points belong to ${\operatorname*{argmin}}H$.
Proof: Since for $x=(x_{1},\dots,x_{n})^{\top}\in\mathbb{R}^{n}$
$f_{\sigma}(x)=\sum_{i=1}^{n}f_{\sigma}(x_{i}),$
we have $f_{\sigma_{i+1}}(x)\geq f_{\sigma_{i}}(x)$ for every $x\geq 0$. Hence, for every $\widetilde{W}\in\mathbb{R}^{p^{2}}$, it follows that $H_{\sigma_{i+1}}(\widetilde{W})\geq H_{\sigma_{i}}(\widetilde{W})$, and $\{H_{\sigma_{i}}(\widetilde{W})\}_{i\in\mathbb{N}}$ is nondecreasing. By Proposition 7.4 of [56], ${\operatorname*{e-lim}}_{i}H_{\sigma_{i}}$ exists and equals $\sup_{i}[{\operatorname*{cl}}H_{\sigma_{i}}]$. Based on the fact $\lim\limits_{\sigma\downarrow 0^{+}}f_{\sigma}(|x|)=\|x\|_{0}$, it follows that $\sup_{i}[{\operatorname*{cl}}H_{\sigma_{i}}](\widetilde{W})=H(\widetilde{W})$. Obviously, for every $\sigma_{i}$, $H_{\sigma_{i}}(\widetilde{W})\geq r_{1}(\widetilde{W})$, and $r_{1}(\widetilde{W})$ is a coercive function based on the fact
$\langle{\rm vec}(R),\widetilde{W}\rangle=\langle R,W\rangle\geq\lambda_{{\rm min}}(R){\rm tr}(W)\geq\lambda_{{\rm min}}(R)\|W\|_{2}$ (57)
[…]
Hence, theoretically, one can obtain the minimizers of optimization ($\textbf{P}^{\prime\prime}_{\sigma}$) by successively solving optimization problem ($\textbf{P}_{\sigma}$) with $\sigma\downarrow 0$. As mentioned above, we hope to solve problem ($\textbf{P}^{\prime\prime}_{\sigma}$) by solving ($\textbf{BP}\sigma-1$) and ($\textbf{BP}\sigma-2$) alternately, known as BCD method. However, due to the nonconvex and nonsmooth term $h(\widetilde{W},y)$, direct BCD may fail to converge and cause stable cyclic behavior [58]. It is worth to note that the BCD method can be regarded as a special case of BSUM framework (Algorithm 1); and hence, instead of BCD method, we will design suitable algorithm for solving ($\textbf{P}^{\prime\prime}_{\sigma}$) under a more general framework, i.e., BSUM framework.
The auxiliary function (see Definition 2) $u_{2,\sigma}(y,X_{k})$ can be selected as
$u_{2,\sigma}(y,X_{k})=\langle{\rm vec}(R),\widetilde{W}_{k}\rangle+r_{2,\sigma}(y)+h(\widetilde{W}_{k},y)$ (58)
with $X_{k}=(\widetilde{W}_{k},y_{k})$. Then, the subproblem
$\underset{y\in\mathbb{R}^{mn}_{+}}{\operatorname*{argmin}}~{}u_{2,\sigma}(y,X_{k})$
within BSUM framework is equivalent to problem ($\textbf{BP}\sigma-2$), which is further equivalent to the following problem
$\min\limits_{y\in\mathbb{R}^{mn}_{+}}g^{*}_{\sigma}(-y)+y^{\top}\left|(V_{2}\otimes V_{1})\widetilde{W}_{k}\right|.$ (59)
Lemma 5.
The global minimizer $y_{k,\sigma}^{*}$ of optimization problem (59) (equivalently, ($\textbf{BP}\sigma-2$)) is unique, and can be expressed as
$y_{k,\sigma}^{*}=\nabla f_{\sigma}\left(\left|(V_{2}\otimes V_{1})\widetilde{W}_{k}\right|\right).$ (60)
Proof: Optimality condition implies that
$0\in-\partial g^{*}_{\sigma}(-y_{k,\sigma}^{*})+\left|(V_{2}\otimes V_{1})\widetilde{W}_{k}\right|.$
Since $g^{*}_{\sigma}$ is a proper, closed and convex function, by Proposition 11.3 of [56], we have
$-y_{k,\sigma}^{*}\in\partial g_{\sigma}\left(\left|(V_{2}\otimes V_{1})\widetilde{W}_{k}\right|\right).$
Noticing that $g_{\sigma}=-f_{\sigma}$ is a smooth function on +$\mathbb{R}^{mn}_{+}$, then subgradient operator degenerates to gradient operator, i.e.,
$y_{k,\sigma}^{*}=\nabla f_{\sigma}\left(\left|(V_{2}\otimes V_{1})\widetilde{W}_{k}\right|\right).$
[…]
We then discuss the selection of $u_{1,\sigma}(\widetilde{W},X_{k})$. The proximal term is introduced, i.e.,
$u_{1,\sigma}(\widetilde{W},X_{k};\lambda)=$ $\langle{\rm vec}(R),\widetilde{W}\rangle+r_{2,\sigma}(y_{k})+h(\widetilde{W},y_{k})$ (61)
$+\frac{1}{2\lambda}\left\|\widetilde{W}-\widetilde{W}_{k}\right\|^{2},$
where $\lambda\in\mathbb{R}$ is a constant. Now, we direct our attention on solving the subsequent optimization problem rather than ($\textbf{BP}\sigma-1$)
$\min\limits_{\widetilde{W}}{}$ $u_{1,\sigma}(\widetilde{W},X_{k};\lambda)$ (62)
s.t. $\widetilde{W}\in\mathcal{X}_{1},$
with
$\mathcal{X}_{1}=\{\widetilde{W}\colon\widetilde{W}\in\Gamma_{+}^{p},\widetilde{\Psi}_{i}\in\Gamma_{+}^{n},i=1,\dots,M,\widehat{A}\widetilde{W}=0\}$
which is closed convex set. First, we claim that $\{H_{\sigma}(\widetilde{W}_{k})\}$ is monotonically nonincreasing.
Lemma 6.
Given an initial feasible point $\widetilde{W}_{0}$, let $\widetilde{W}_{k}$ be the iterative sequence generated by Algorithm 1 with $u_{1},u_{2}$ given by (61), (58). Then $H_{\sigma}(\widetilde{W}_{k+1})\leq H_{\sigma}(\widetilde{W}_{k}).$
Proof: Let $F(\widetilde{W},X_{k})=\langle{\rm vec}(R),\widetilde{W}\rangle+r_{2,\sigma}(y_{k})+h(\widetilde{W},y_{k})$. According to the fact that
$y_{k}=\nabla f_{\sigma}\left(\left|(V_{2}\otimes V_{1})\widetilde{W}_{k}\right|\right),$
$F(\widetilde{W},X_{k})$ can be equivalently regarded as $F(\widetilde{W},\widetilde{W}_{k})$. Obviously, for all $\widetilde{W}$, we have
$u_{1,\sigma}(\widetilde{W},X_{k};\lambda)\geq F(\widetilde{W},\widetilde{W}_{k})\geq H_{\sigma}(\widetilde{W}).$
Hence, the following inequality holds
$H_{\sigma}(\widetilde{W}_{k+1})\leq u_{1,\sigma}(\widetilde{W}_{k+1},X_{k};\lambda)\leq u_{1,\sigma}(\widetilde{W}_{k},X_{k};\lambda)$
$=F(\widetilde{W}_{k},\widetilde{W}_{k})=H_{\sigma}(\widetilde{W}_{k}).$
[…]
Obviously, $\{H_{\sigma}(\widetilde{W}_{k})\}$ is bounded below by 00. Hence, by basic mathematic analysis, $H_{\sigma}(\widetilde{W}_{k})\downarrow H_{\sigma}^{\infty}$, where $H_{\sigma}^{\infty}$ is a positive constant. An inquiry that arises is whether or not $H_{\sigma}^{\infty}=H_{\sigma}^{*}\doteq\min H_{\sigma}(\widetilde{W})$. Regretfully, such question remains open, but we have the following result.
Theorem 5.
Every cluster point $z=(z_{1},z_{2})$ of the iteration generated by Algorithm 1 with $u_{1},u_{2}$ given by (61), (58) is a coordinatewise minimizer of the optimization problem ($\textbf{P}^{\prime\prime}_{\sigma}$). In addition, if $H_{\sigma}(\cdot,\cdot)$ is regular at $z$, then $z$ is a stationary point of ($\textbf{P}^{\prime\prime}_{\sigma}$).
Proof: For a feasible $X_{0}=(\widetilde{W}_{0},y_{0})$, by (57), the sublevel set
$\mathcal{X}^{0}\doteq\left\{X=(\widetilde{W},y)\colon H_{\sigma}(X)\leq H_{\sigma}(X_{0})\right\}$
[…]
There exists a significant gap between coordinatewise minimizer and stationary point (or local minimizer). When the objective function is regular at coordinatewise minimizer, the coordinatewise minimum becomes a local minimum by the above theorem. In Lemma 3.1 of [59], there is a few discussion about regularity property of a Gateaux-differentiable function, but $H_{\sigma}(\widetilde{W},y)$ is generally not Gateaux-differentiable in this paper. Furthermore, if we assume that $H_{\sigma}^{\infty}=H_{\sigma}^{*}$ holds, we can obtain the following theorem.
Theorem 6.
If $H_{\sigma}^{\infty}=H_{\sigma}^{*}$, then any cluster point of $\{\widetilde{W}_{k}\}$ belongs to ${\rm argmin}H_{\sigma}(\widetilde{W})$.
Proof: Since $f_{\sigma}(|(V_{2}\otimes V_{1})\widetilde{W}|)\geq 0$, it is obvious that, for all $\alpha\in\mathbb{R}$, ${\rm lev}_{\leq\alpha}H_{\sigma}(\widetilde{W})$ is compact. Then the set
$\left\{\widetilde{W}\colon H_{\sigma}(\widetilde{W})\leq H_{\sigma}(\widetilde{W_{0}})\right\}$
is compact. Hence, by $H_{\sigma}(\widetilde{W}_{k})\downarrow H_{\sigma}^{\infty}$, $\{\widetilde{W}_{k}\}_{k\in\mathbb{N}}$ is compact. Thus, there exists a subsequence $\{\widetilde{W}_{k_{n}}\}_{n\in\mathbb{N}}$ of $\{\widetilde{W}_{k}\}$ that converges to $\widetilde{W}_{\infty}$. Based on the fact that $H_{\sigma}(\widetilde{W})$ is lower semi-continuous, we have
$H_{\sigma}^{*}=H_{\sigma}^{\infty}=\underset{k\to\infty}{\rm liminf}H_{\sigma}(\widetilde{W}_{k})\geq H_{\sigma}(\widetilde{W}_{\infty})\geq H_{\sigma}^{*}.$
[…]
Our focus is now directed towards resolving subproblem (62). Obviously, (62) can be expressed as the following constrained convex optimization problem:
$\min_{\widetilde{W}}{}{}$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\gamma y_{k}^{\top}\left|(V_{2}\otimes V_{1})\widetilde{W}\right|+\frac{1}{2\lambda}\left\|\widetilde{W}-\widetilde{W}_{k}\right\|^{2}$ (63)
${\rm s.t.}{}{}$ $\widetilde{W}\in\Gamma^{p}_{+},$
$\widetilde{\Psi}_{i}\in\Gamma_{+}^{n},~{}i=1,\dots,M,$
$\widehat{\mathcal{A}}\widetilde{W}=0.$
By introducing augment variable $\widetilde{P}=(V_{2}\otimes V_{1})\widetilde{W}$, (63) becomes
$\min_{\widetilde{W},\widetilde{P}}{}{}$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\gamma y_{k}^{\top}\left|\widetilde{P}\right|+\frac{1}{2\lambda}\left\|\widetilde{W}-\widetilde{W}_{k}\right\|^{2}$ (64)
${\rm s.t.}{}{}$ $\widetilde{W}\in\Gamma^{p}_{+},$
$\widetilde{\Psi}_{i}\in\Gamma_{+}^{n},~{}i=1,\dots,M,$
$\mathcal{A}\widetilde{W}+\mathcal{B}\widetilde{P}=0.$
In this section, we use the notation
$f_{1}(\widetilde{W};\lambda)$ $=\langle{\rm vec}(R),\widetilde{W}\rangle+\frac{1}{2\lambda}\left\|\widetilde{W}-\widetilde{W}_{k}\right\|^{2},$
$f_{2}(\widetilde{W})$ $=\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M}),$
$g(\widetilde{P})$ $=\gamma y_{k}^{\top}\left|\widetilde{P}\right|,$
$f(\widetilde{W})$ $=f_{1}(\widetilde{W})+f_{2}(\widetilde{W}),$
and optimization problem (64) is equivalent to
$\min_{\widetilde{W},\widetilde{P}}{}{}$ $f_{1}(\widetilde{W};\lambda)+f_{2}(\widetilde{W})+g(\widetilde{P})$ (65)
${\rm s.t.}{}{}$ $\mathcal{A}\widetilde{W}+\mathcal{B}\widetilde{P}=0.$
Interestingly, $g(\widetilde{P})$ is nothing but a weighted-$\ell_{1}$-norm with weight values given by $y_{k}$. Hence, compared to optimization problem (19), the only difference is the smooth and strongly convex term $\frac{1}{2\lambda}\|\widetilde{W}-\widetilde{W}_{k}\|^{2}$ in $f_{1}(\widetilde{W};\lambda)$. Thus, we can utilize optimization scheme (21) by slightly modifying Algorithm 3 (just change $w_{ij}$ to $y_{k}^{(ij)}$, where $y_{k}^{(ij)}$ is the $(i,j)$-th element of $y_{k}$).
Algorithm 5 BSUM-based $\ell_{0}$-penalty Decentralized LQ
  Given : Feasible initial point $X_{0}=(\widetilde{W}_{0},y_{0})\in\mathcal{X}_{1}\times\mathbb{R}_{+}^{mn},\lambda>0,\gamma>0,\sigma_{0}\leq 1,\alpha\in(0,1)$
  Result: $X_{k}=(\widetilde{W}_{k},y_{k})$
  for $i=1,2,3\dots$ do
     Set $\sigma_{i}=\alpha\sigma_{i-1}$
     Set $k=0$
     for $k=0,1,2,...$ do
        Let $k=k+1,i=(k\mod 2)+1$
        if $i==1$ then
           $\widetilde{W}_{k}\leftarrow$ALM based Decentralized LQ
           Set $y_{k}=y_{k-1},X_{k}=(\widetilde{W}_{k},y_{k})$
        end if
        if $i==0$ then
           Update $y_{k}=\nabla f_{\sigma_{i}}\left(\left|(V_{2}\otimes V_{1})\widetilde{W}_{k}\right|\right)$
           Set $\widetilde{W}_{k}=\widetilde{W}_{k-1},X_{k}=(\widetilde{W}_{k},y_{k})$
        end if
        if Stopping Criterion==absent$=$ =True then
           Break
        end if
     end for
     Set $X_{0}=X_{k}$
  end for
Remark 15.
In fact, if $u_{1,\sigma}(\widetilde{W},X_{k})$ is selected as
$u_{1,\sigma}(\widetilde{W},X_{k})=\langle{\rm vec}(R),\widetilde{W}\rangle+r_{2,\sigma}(y_{k})+h(\widetilde{W},y_{k}),$
the iterative sequence of the algorithm BSUM can also converge to the coordinatewise minimizers of $H_{\sigma}(\widetilde{W},y)$. However, we introduce the proximal term $\frac{1}{2\lambda}\|\widetilde{W}-\widetilde{W}_{k}\|^{2}$ into $u_{1,\sigma}$, which appears to be meaningless or potentially resulting in a decelerated convergence of the BSUM algorithm. Empirically, the proximal term makes the objective of (65) be strongly convex and accelerate the solving for subproblem (65). Therefore, a trade-off exists between the convergence rate of the algorithm BSUM for optimization problem ($\textbf{P}^{\prime\prime}_{\sigma}$) and the convergence rate of (21) for subproblem (65).
Remark 16.
The method proposed in this section exhibits a significant level of complexity, and there exists no theoretical evidence to support the superiority of the method introduced in this section over Algorithm 3 (or Algorithm 4). Therefore, this section mostly focuses on theoretical discussions, particularly about the variational characteristics of the optimization problem induced by the $\ell_{0}$-penalty and the relation between $\ell_{1}$-penalty problem (15) and $\ell_{0}$-penalty problem. It is recommended that, for finding decentralized controllers, Algorithm 3 (or Algorithm 4) is employed.
7 Numerical Examples
In this section, a few numerical examples are represented that illustrate the theoretical results of this paper, where the noise $w$ (see (66)) is characterized by a impulse disturbance vector.
Example 1: Consider $x=[x_{1},x_{2},x_{3}]^{\top}$ and a linear system
$\dot{x}$ $-Ax+B_{2}u+B_{1}w,$ (66)
$z$ $=Cx+Du,$
$u$ $=-Kx,$
where
$A=\begin{bmatrix}0.2220&0.9186&0.7659\\ 0.8707&0.4884&0.5184\\ 0.2067&0.6117&0.2968\end{bmatrix},B_{1}=I_{3},C=\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix},$
$B_{2}=\begin{bmatrix}0.9315&0.7939\\ 0.9722&0.1061\\ 0.5317&0.7750\end{bmatrix},D=\begin{bmatrix}0&0\\ 1&0\\ 0&1\end{bmatrix}.$
By Algorithm 3, we can obtain the relation between $\gamma$ and $N$ (the number of 00 in feedback gain $K$; see Fig. 2).
$\gamma=0$ $\gamma=1$ $\gamma=5$ $\gamma=10$ $\gamma=20$ $\gamma=50$
$J(K)$ 1.92 3.51 5.78 6.56 7.20 7.91
$N$ 0 1 2 3 3 3
Table 1: Relation between $\gamma$ and $J(K)$
Figure 2: Relation between $\gamma$ and $N$
Remarkably, if $(A,B_{2})$ is known, we have $\langle R,W\rangle=J(K)$ rather than an upper bound of $J(K)$; see Remark 1 above. The TABLE 1 illustrates the relation between $\gamma$ and LQ cost $J(K)$, and we can observe that $J(K)$ increases with the increase of $\gamma$. In particular, when $\gamma=0$, the problem studied in this example degenerates to classic centralized LQ problem, which can be solved by letting $\gamma\to 0$ and utilizing Algorithm 3 (or Algorithm 4; see Proposition 2) or solving Algebraic Riccati Equation. As illustrated in Fig. 2 and TABLE 1, the parameter $\gamma$ serves as the balance between LQ cost $J(K)$ and the level of sparsity of feedback gain $K$. Additionally, the LQ cost changes very little and the level of sparsity $N$ remains unchange when $\gamma$ is large enough; which implies if we only need a sufficiently sparse feedback gain $K$, we can simply let $\gamma$ sufficiently large.
Specifically, when $\gamma=10$, we can obtain the following result (see Fig. 3). Denoting
$W_{k}=\mathrm{vec}^{-1}(\widetilde{W}_{k})=\begin{bmatrix}W_{1,k}&W_{2,k}\\ W_{2,k}^{\top}&W_{3,k},\end{bmatrix},$
where $\{\widetilde{W}_{k}\}$ is the iteration sequence of Algorithm 3, we have that
$K_{k}=W_{2,k}^{\top}W_{1,k}^{-1}=\begin{bmatrix}K_{11,k}&K_{12,k}&K_{13,k}\\ K_{21,k}&K_{22,k}&K_{23,k}\end{bmatrix}$
converges to the stabilizing feedback gain by Theorem 1 and Corollary 2, and indeed the colored graphs in Fig. 3 indicate the convergence of $K_{ij,k}$ for $i=1,2$, $j=1,2,3$.
Figure 3: Feasible gain with $\gamma=10$
The optimal decentralized feedback gain $K$ is given by
$K=\begin{bmatrix}0.6192&2.5269&0\\ 0&0&1.3068\par\end{bmatrix}.$
The responses of all the state variables are illustrated in Fig. 4, and it can be seen that the stability of system is guaranteed.
Figure 4: System response
Example 2: Consider the linear system (66) with
$A=\begin{bmatrix}0.3079&0.1879&0.1797&0.2935&0.6537\\ 0.5194&0.2695&0.5388&0.9624&0.5366\\ 0.7683&0.4962&0.2828&0.9132&0.9957\\ 0.7892&0.7391&0.7609&0.5682&0.1420\\ 0.8706&0.1950&0.2697&0.4855&0.9753\end{bmatrix},$
$B_{2}=\begin{bmatrix}0.6196&0.6414\\ 0.7205&0.9233\\ 0.2951&0.8887\\ 0.6001&0.6447\\ 0.7506&0.2956\end{bmatrix},$
$B_{1}=I_{5},C=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\end{bmatrix},D=\begin{bmatrix}0&0\\ 0&0\\ 1&0\\ 0&1\\ 0&0\end{bmatrix}.$
Letting $\gamma=10$, the optimal decentralized feedback gain $K$ is given by
$K=\begin{bmatrix}1.3355&0&0&0&5.5233\\ 0.8946&1.8952&1.0401&3.8119&0\end{bmatrix}.$
The responses of all the state variables are illustrated in Fig. 5, and the stability of system is guaranteed.
Figure 5: System response
Example 3: Consider the linear system (66) with
$A=\begin{bmatrix}0&1&0\\ 0&0&1\\ 0&0&0\end{bmatrix},B_{1}=I_{3},B_{2}=\begin{bmatrix}0.9315&0.7939\\ 0.9722&0.1061\\ 0.5317&0.7750\end{bmatrix},$
$C=\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix},D=\begin{bmatrix}0&0\\ 1&0\\ 0&1\end{bmatrix}.$
We utilize Algorithm 3 and Algorithm 4 with $\gamma=10$, respectively, and the results are shown below (see Fig. 6 and Fig. 7).
Figure 6: Convergence of Algorithm 3
Figure 7: Convergence of Algorithm 4
By using Algorithms 3 and 4, denote the obtained feedback gain as $K_{1},K_{2}$, respectively; we have $J(K_{1})=8.81$ and $J(K_{2})=9.39$. Moreover, it holds $\|K_{1}\|_{0}=3$ and $\|K_{2}\|_{0}=4$. Generally, the obtained LQ cost $J(K_{1})$ is slightly smaller than the one by Algorithm 4; this is mainly caused by $g_{\mathcal{Q}}(x)\gg g(x)$ when $\|x\|$ is large. Additionally, the obtained feedback gains exhibit nearly the same level of sparsity by using Algorithms 3 and 4, however, the required iteration numbers could be really different. As shown in Fig. 6 and 7, we can observe that Algorithm 3 converges within 800 steps, while Algorithm 4 converges within only 80 steps. That is because the Algorithm 3 exhibits $\mathcal{O}(1/k)$ convergence rate, while Algorithm 4 exhibits $\mathcal{O}(1/k^{2})$ convergence rate.
8 Conclusion
This work investigates the $\mathcal{H}_{2}$-guaranteed cost decentralized control problem under convex parameterization and convex-bounded uncertainty, specifically focusing on the inclusion of a sparsity constraint on the feedback gain. The formulation of approximate separable constraint optimization problems is presented, and a convex optimization framework is suggested to achieve the global optimizer for these problems. In the near future, an examination will be conducted on decentralized LQ problems with nonconvex optimization perspective.